### v1

# Remote Work and Women’s Labor Supply: The New Gender Division at Home

Abstract
We study how increases in remote work opportunities for men affect their spouses’ labor supply. Exploiting variation in the change in work-from-home (WFH) exposure across occupations before and after the COVID-19 pandemic, we find that women whose husbands experienced larger WFH increases are over 2 percentage points more likely to be employed, equivalent to a 4% rise relative to pre-pandemic levels. Evidence from time-use diaries and childcare questionnaires suggests these effects are driven by intra-household reallocation of child-caring time: women are less likely to engage in primary childcare activities, while men working at home partially compensate by covering more for their spouse. These results highlight the role of intra-household spillovers and bargaining in shaping the labor market consequences of remote work.
JEL Codes: D13, J13, J22, J16, J31
Keywords: work-from-home, workplace flexibility, labor supply, intra-household allocation, gender gaps, time use, childcare
1 Introduction
Over the past century, gender gaps in education, human capital, and labor market outcomes have narrowed substantially in developed countries (Olivetti and Petrongolo, 2016). Nonetheless, a persistent divergence remains in labor market trajectories following parenthood (Kleven et al., 2024). After childbirth, women frequently reduce their labor force participation or exit employment altogether (Kleven, 2022), and those who remain employed often face large earnings penalties relative to men (Bertrand et al., 2010). These gender disparities are largely driven by (i) the unequal division of childcare responsibilities within households, and (ii) the fact that many jobs - including high-paying white collar positions - have historically lacked flexibility options to allow women to balance both their careers and motherhood (Goldin, 2014).
This paper examines whether the expansion of one aspect of workplace flexibility - namely, remote work opportunities - can shift these intra-household dynamics and encourage greater labor force participation among women with children. We focus on U.S. households with young children and combine occupation-level data on remote work vacancies with rich household-level survey data from the CPS and other national datasets. Leveraging quasi-experimental variation in remote work exposure across occupations before and after the pandemic, we estimate whether large-scale changes in remote work availability altered household labor supply decisions, particularly by enabling mothers to remain in or return to the labor force.
Our motivation builds on the idea, emphasized by Goldin (2014), that the last chapter of gender convergence may depend on restructuring work to allow for greater flexibility. Workplace flexibility has been shown to help parents, especially mothers, accommodate both predictable and unexpected caregiving demands (Angelici and Profeta, 2024). Evidence also shows that women and parents of young children place a high value on flexible work arrangements (Mas and Pallais, 2017; Bloom et al., 2022; Aksoy et al., 2022), and are more averse to commuting (Le Barbanchon et al., 2021). While not synonymous with flexibility, remote work is strongly associated with it, particularly in allowing job tasks to be performed from home, thus potentially facilitating a more equitable division of household production.
From the perspective of a collective household model (Chiappori, 1992), increased parental presence at home may ease caregiving constraints and shift bargaining dynamics within couples. Despite this potential, there is limited causal evidence on how remote work affects women’s labor supply, particularly via intra-household channels. Our paper fills this gap by documenting how the expansion of remote work, especially for men, affects their spouses’ employment and time use.
The backdrop for our analysis is the COVID-19 pandemic, which triggered a rapid and persistent shift in remote work norms. In the U.S., only 7% of full paid workdays were conducted from home in 2019. By June 2023, this share had quadrupled to 28% (Barrero et al., 2023). However, the magnitude of this shift varied widely across occupations: while some sectors (e.g., computer science and professional services) adopted remote work at scale, others (e.g., transport, healthcare, education) largely returned to in-person arrangements. In Figure 1, we illustrate this heterogeneity using data from Hansen et al. (2023) on remote work job postings between 2019 and 2024. We use this variation to study whether and how women’s labor market outcomes respond to changes in household exposure to remote work opportunities, shaped by the husband’s occupation.
Our empirical strategy compares the labor market outcomes of women with young children whose husbands experienced a large increase in the availability of remote work, measured as the change in the share of job vacancies offering WFH in their occupation, to those whose husbands experienced little or no change. Our main hypothesis is that an increase in the husbands’ available WFH options leads to an increase in the wife’s engagement in the labor market. This design offers two main advantages. First, the shock plausibly expands the household’s set of feasible work–care arrangements, creating opportunities for women to enter or re-enter in the labor market. Second, defining treatment through the husband’s occupation avoids a major selection problem: in most cross-sectional surveys, a woman’s occupation is only observed if she is employed. Using changes in WFH exposure in the wife’s occupation as treatment would condition on labor force participation, precluding identification of extensive margin effects – our primary outcome of interest.
Figure 1: Evolution of WFH by occupation (2 digits).
Notes: Author’s elaboration using WFH job postings data from Hansen et al. (2023). Classification based on SOC 2018 2-digits.
Our first-stage results confirm that the large, occupation-level shifts in remote work feasibility post-COVID led to substantial increases in men’s actual WFH take-up, which almost tripled relative to pre-pandemic levels, with an average rise of about 20 percentage points. In response, we find that wives’ labor market outcomes improved meaningfully: their probability of employment and labor force participation rose by approximately 2.5 percentage points, and their wages increased by around 8%. We also document a 10 percentage point increase in women’s own WFH rates, a rise of roughly 75% relative to their pre-treatment mean. We find some evidence that these positive employment effects for women are mainly driven by households with toddlers but this heterogeneity analysis is not fully robust.
To validate the credibility of our identification strategy, we conduct a broad set of robustness checks that address potential threats to our findings. We examine the possibility of changes in household composition or selection into treatment following the pandemic, including tests for differential fertility and marriage patterns and reweighting strategies that align pre- and post-COVID covariate distributions. We also assess whether our results are driven by specific occupation groups through leave-one-out analyses, and evaluate the sensitivity of our estimates to alternative fixed effect structures and demographic controls. Additionally, we explore the robustness of our findings to varying definitions of the post-treatment period and sample restrictions. Across all these exercises, our main results remain consistent, reinforcing the interpretation that increased WFH options for men played a causal role in raising women’s labor market engagement.
We then proceed to study how shifts in the household’s allocation of labor supply, childcare responsibilities, and household production activities could be driving our employment effects. To do this, we rely on two data sources: the time-use diary data from the American Time Use Survey (ATUS) and the topical modules on childcare from the Survey of Income and Program Participation (SIPP).
Our findings reveal that treated women increase their time spent working, with no significant change in overall household production. However, we observe that women reduce time spent on childcare as a primary activity and husbands timidly seem to compensate for that as both ATUS-based evidence and SIPP topical modules highlight. Households are also less likely to pay for daycare services. We also find evidence that these adjustments in women’s employment and the allocation of childcare are not coming at the expense of the children’s well-being.
Overall, our findings suggest that increases in men’s access to remote work can influence their spouses’ labor supply decisions, particularly in households with young children. Our analysis of time-use and childcare patterns offers tentative evidence that the effects may stem from shifts in household responsibilities: women are reducing their primary childcare time and households spend less in daycare services, while husbands modestly seem to provide more primary childcaring and take care of children while women work. Taken together, the results point to the value of analyzing remote work not only as an individual labor market arrangement, but also as a household-level shock, offering new perspective on how WFH access may affect gender gaps in employment.
Related Literature
Our main contribution is to the extensive literature on gender inequality in the labor market (Waldfogel, 1998; Anderson et al., 2002; Ejrnæs and Kunze, 2013; Kleven et al., 2024). Goldin (2014) emphasizes that closing gender gaps requires structural changes in the labor market that enhance flexibility. More recently, Goldin (2022) has argued that the rise of WFH should eventually benefit employment of college-educated mothers. It has been difficult to disentangle net employment effects in the immediate aftermath of the Covid pandemic due to a shortage in the supply of childcare services (Heggeness et al., 2021). By exploiting the long-run effect of Covid on the availability of remote work options, we are among the first papers to find that WFH can lead to closing gender gaps in employment, even in developed country. Our result resembles most closely those by Jalota and Ho (2024), which find in an RCT in India that increased remote work options doubles female labor force participation among data-entry workers.
Second, we contribute to the growing literature on how working arrangements affect labor market dynamics. While existing research has primarily focused on the direct effects of flexible work arrangements on individual outcomes—highlighting their growing value for both workers and firms (Bloom et al., 2015, 2022; Angelici and Profeta, 2024; Choudhury et al., 2024; Mas and Pallais, 2017)—relatively few studies have examined how changes in work arrangements reshape decision-making at the household level. Prior work shows that women and parents place especially high value on flexibility (Mas and Pallais, 2020; Aksoy et al., 2022), yet evidence on intra-household responses remains limited. We contribute to this emerging strand by leveraging a large-scale, quasi-experimental shift in remote work feasibility to study how increased WFH access for men affects household labor supply and home production. [^1] Closest to our work is Harrington and Kahn (2023), who examine how increases in remote work opportunities, as proxied by women’s college attainment, affect mothers’ employment and earnings trajectories In contrast, we estimate the spillover effects of men’s remote work access on their spouses’ labor market behavior and on the intra-household time allocation.
Finally, our paper contributes to the literature on the labor supply effects of parental leave and temporal flexibility policies for parents. Prior work generally finds that such policies have limited, and sometimes even negative, effects on women’s long-term labor market outcomes (Kleven et al., 2024; Ginja et al., 2023, 2020; Canaan, 2022). At best, fathers’ access to temporal flexibility has been shown to improve maternal postpartum health, with little impact on female labor supply (Persson and Rossin-Slater, 2024). How, then, do we reconcile these findings with our own evidence of sizable and persistent effects from increased remote work? We argue that the scale and perceived permanence of the WFH shock fundamentally differ from traditional parental leave policies. Rather than offering short-term accommodations around childbirth, the pandemic-driven shift toward remote work restructured household production possibilities and gave families a broader set of tools to jointly optimize labor supply and caregiving responsibilities over the longer run.
Outline
The remainder of the paper is structured as follows. Section 2 describes our main data sources and outlines the sample restrictions used to construct comparable treatment and control groups. Section 3 presents our empirical strategy and discusses the identification assumptions underlying our analysis. In Section 4, we report the main results on women’s labor market outcomes. Section 5 provides a series of robustness checks and placebo tests to validate our identification strategy. In Section 6, we explore time-use and childcare data to shed light on the mechanisms behind the observed effects. Section 7 concludes.
2 Data
2.1 Data Sources
We draw on several public-use U.S. household surveys for our analysis. Our primary results are based on the Current Population Survey (CPS), with complementary evidence from the American Community Survey (ACS) and the Survey of Income and Program Participation (SIPP). To investigate underlying mechanisms, we use detailed time-use information from the American Time Use Survey (ATUS), with additional evidence from SIPP. To measure changes in WFH exposure at the occupation level, we rely on job postings data on remote and hybrid work arrangements (Hansen et al., 2023), as well as established measures of WFH feasibility by occupation from Dingel and Neiman (2020). We describe each data source in greater detail below.
Current Population Survey (CPS)
We use public-use microdata from the CPS (Flood et al., 2024) spanning the years 2016–2024. Our analysis integrates data from both the basic monthly files and the Annual Social and Economic Supplement (ASEC), also known as the March files. The CPS provides detailed information on all household members, including the presence and age of children, as well as occupational codes for both husbands and wives. Critically, it includes several measures of labor supply at the extensive margin, such as employment status and labor force participation, available at both weekly and annual frequencies [^2] See Section 3.3 and Appendix B of Kleven (2024) for a detailed discussion of differences between the CPS basic monthly and March CPS files, and between weekly and annual variables. Weekly variables refer to outcomes in the previous week, while annual variables refer to the previous calendar year., as well as intensive margin measures such part-time status, usual weekly hours worked, and wages [^3] Throughout the paper, we define intensive margin outcomes conditional on employment.. While weekly measures are available in both the basic monthly and March files, annual employment/labor force participation and income variables are available exclusively in the March files. Occupation codes at the 4-digit level are based on the Census Occupation Classification and are already harmonized to the 2010 coding scheme.
WFH Job Postings Data (Hansen et al., 2023).
Our preferred measure of WFH exposure is derived from job postings data on remote and hybrid work arrangements from 2019 to 2024. These data were constructed by a team of researchers in collaboration with Lightcast, using a large language model (LLM) trained to classify job advertisements based on whether the position allows remote work at least one day per week. The resulting measures are available at the month-year–occupation level [^4] Occupations are classified using the Standard Occupational Classification 2018 (SOC) at 3 digits. and at the month-year- industry level. [^5] Industry codes are classified using North American Industry Classification System (NAICS) 2022 at 3 digits. We use this data to build WFH probability changes by occupation, which will be key for our definition of treatment, as we discuss in Subsection 2.3.
American Time Use Survey (ATUS)
We use the ATUS (Flood et al., 2023) samples from 2017 to 2024. The ATUS draws its sample from households completing their final CPS interview, ensuring consistency in demographic and labor market variables across surveys. Each ATUS respondent reports detailed time use for a designated 24-hour period, allowing us to examine outcomes such as time spent on childcare, labor and household production. We can also assess the time doing secondary activities, such as the time spent childcaring while doing another activity. These measures are central to investigating the mechanisms underlying the intra-household spillovers we study. Importantly, the ATUS interviews only one respondent per household. By merging the ATUS with the CPS using unique person identifiers, we obtain time-use data for either the wife or the husband, but never both spouses within the same household.
American Community Survey (ACS)
We replicate our main findings using ACS (Ruggles et al., 2025) samples from 2017 to 2023. The ACS provides information on current weekly employment status and labor force participation, as well as retrospective measures covering the previous 12 months. [^6] This contrasts with the CPS, where annual outcomes refer to the previous calendar year. These include indicators for employment, months worked, usual hours worked per week, and annual wage and salary income.
Survey of Income and Program Participation (SIPP).
We also replicate our main findings using the 2018-2024 SIPP panel datasets. The SIPP is a nationally representative longitudinal survey that provides comprehensive information for dynamics on household’s employment status and composition, income and government program participation. Starting on 2018, the SIPP has started implementing a 4-year rotational panel survey. [^7] In 2018 the SIPP redesigned the survey that make it difficult to compare the 2014-2017 SIPP panels and post 2018 panels. The 2018-2024 SIPP panels cover information for reference years 2017-2023. Households are interviewed once a year and asked to recall their detailed monthly employment status and income during the previous year. After the year they join the SIPP sample, households are followed and interviewed for the next three years, after which they are dropped from the sample. Each year new households join the sample, so up to 4 different household cohorts populate each year’s SIPP sample. [^8] The 2019 SIPP sample experienced significant logistical difficulties, which led the US Census to prioritize resources on following through with the 2018 sample. As a result, the sample of new households that joined in 2019 was not adequately surveyed and the decision was made to not follow up on them in subsequent surveys. In addition to using the SIPP to validate our cross-sectional findings from the CPS and ACS, we leverage its panel structure to estimate alternative specifications of our main empirical strategy that are only feasible with panel data. [^9] Specifically, we incorporate household fixed effects into our baseline specification to account for time-invariant unobserved heterogeneity. Also, we use SIPP child care topical modules, which are available for all years in our 2018-2024 SIPP panels, to complement mechanism analysis.
Teleworkability Index (Dingel and Neiman, 2020)
As an alternative measure of remote work potential, we incorporate the teleworkability index developed by Dingel and Neiman (2020), which classifies occupations based on their suitability for being performed from home. [^10] The index is constructed using responses from two ONET surveys: the Work Context Questionnaire and the Generalized Work Activities Questionnaire. These surveys assess the importance of physical presence in task performance, allowing for an evaluation of whether specific tasks within each occupation could feasibly be performed remotely. The final occupation-level indicator is calculated as the average teleworkability score across tasks within each occupation. In contrast to our primary measure based on changes in job postings, this index captures the inherent feasibility of telework by occupation, rather than shifts in employer demand for remote work arrangements.
2.2 Sample Restrictions
We restrict our primary analysis, based on the CPS, to dual-parent households in which both the husband and wife reside together and have at least one child of child-rearing age. Specifically, we focus on households with children between the ages of 0 and 8. We exclude households with more than four children. Additionally, we restrict the sample to women aged 20 to 50 and men aged 18 to 64.
In terms of time coverage, we use both the CPS basic monthly and March files spanning 2017 to 2024. The March files contain both weekly labor force variables—comparable to those in the basic monthly files—and additional variables such as annual employment, weeks worked, and wage income, which refer to the previous calendar year. As a result, the effective data coverage for regressions using these annual March files variables is 2016 to 2023.111111Note that occupation in the March files is measured at the time of the survey, while some labor market outcomes—such as annual employment and earnings—refer to the previous year. Therefore, in regressions using these outcomes, we implicitly assume that individuals’ occupation at the time of the survey is the same as in the previous year, which may introduce some measurement error.
To avoid conflating temporary COVID-related labor market disruptions with the long-run shift in remote work, we exclude data from March 2020 to March 2023 in analyses using the CPS basic monthly and March files outcomes referring to the prior year. For regressions using only contemporaneous March CPS files variables—such as weekly labor force status—we exclude data from March 2020 through December 2022, allowing us to include the 2023 March CPS sample and thereby add more post-COVID observations. [^12] The federal COVID-19 Public Health Emergency officially ended on May 11, 2023. For more details, see this link. In the robustness section, we show that our results are not sensitive to alternative definitions of the post-COVID period.
2.3 Treated and Control Group Definitions
We use job postings data from Hansen et al. (2023) to construct occupation-level measures of the change in WFH probability before and after the COVID-19 pandemic. Specifically, we compute the difference in the share of hybrid and remote-work job postings between the post-pandemic period (2023–2024) and the pre-pandemic period (2019). [^13] As an alternative, we also construct a measure based on the change in WFH share between two specific months: July 2023 and July 2019.
We convert occupation codes from the original SOC 3-digit classification used in the job postings data to the occupation codes used in the CPS. Occupations with a WFH probability change at or above the median are classified as treated, and those below the median as control. [^14] As a robustness check, we alternatively define treatment based on the median split within the original SOC 2018 3-digit classification. These occupation-level measures capture the pandemic-induced shift in WFH feasibility and are merged to our household survey data based on the husband’s occupation at the time of the survey. [^15] Ideally, we would match individuals to their pre-pandemic occupation to avoid concerns about endogenous selection into more WFH-intensive occupations on the basis of unobservables.
Appendix Table B.1 lists the 15 most common occupations in the treated and control groups. As expected, occupations that experienced the largest increases in WFH probability after COVID-19 include those in the tech sector, law, and management,such as computer and software engineers, lawyers, and various types of managers. In contrast, occupations with minimal changes in WFH feasibility are concentrated in sales, construction, and retail, as well as in sectors like healthcare (e.g., physicians and surgeons) and education (e.g., teachers), where in-person presence remains a key job requirement.
Appendix Table B.2 and Table B.3 present descriptive statistics for baseline labor market outcomes and household-level covariates, respectively. On average, treated households are more likely to have both spouses employed, [^16] For spouse employment and labor force participation measured in the previous week, this pattern is mechanical: the variable equals 1 for all observations because occupation—and hence treatment status—is only recorded for individuals who are currently employed or in the labor force. with both husbands and wives working at higher rates than in control households. As expected, treated households are more likely to have both spouses highly educated and less likely to have attained only a high school degree or less. Treated wives are also slightly older, on average, than their counterparts in the control group. In terms of family composition, treated households tend to have fewer children and younger children on average. Additionally, labor market outcomes for women in the treated group exhibit larger average changes post-COVID compared to women in the control group, though this pattern is purely descriptive. In all regressions, we control for the education levels of both spouses, the wife’s age, and the number of children in the household. In robustness checks, we additionally include controls for the age of the youngest and oldest child (see Table B.16).
As a robustness exercise, we additionally define treatment using the teleworkability measure developed by Dingel and Neiman (2020), which classifies occupations based on their feasibility to be performed remotely. Occupations at or above the median in this measure are classified as treated, and those below as control. Unlike our baseline, which relies on realized job postings and thus reflects employer demand, this measure is constructed from occupational characteristics and reflects potential—rather than actual—telework feasibility. [^17] Importantly, this classification does not account for demand-side frictions or social norms that may affect the practical implementation of remote work. For instance, while Dingel and Neiman (2020) classify teachers as capable of WFH, post-pandemic trends suggest that in-person instruction remains the preferred mode, largely due to parental demand (Barrero et al., 2023). Crucially, both this and our baseline measure are based on occupational characteristics and do not rely on actual WFH take-up by husbands, which helps avoid concerns about endogenous sorting or selection based on realized behavior (Mas and Pallais, 2017).
3 Empirical Strategy
3.1 Baseline Difference-in-Differences
We exploit variation in the change in work-from-home (WFH) probability across occupations induced by the COVID-19 pandemic. Our identification strategy compares labor market outcomes of women whose husbands were employed in occupations that experienced larger shifts in WFH probability, before and after the pandemic. As explained in the previous section, change in WFH exposure is collapsed at the occupation level and reflects the difference in demand-driven WFH probabilities pre- and post-COVID.
Our baseline specification is:
$\begin{split}y_{i(H),s,t}=\;&f_{s,t}+f_{o(H)}\;+\\ &\beta\times 1\left(o(H)\in\{\text{High $\Delta$WFH Occupations}\}\right)\times 1(t>\text{March 2023})\;+\\ &X_{i(H),t}\delta+\epsilon_{i(H),s,t}\end{split}$ (1)
where $y_{i(H),s,t}$ denotes a labor market outcome for woman $i$, married to husband $H$, observed in state $s$ and time tt. [^18] Time is defined at the month-year level in CPS, ATUS, and SIPP. In ACS, where month of interview is unavailable, time is defined at the year level. The interaction term captures the treatment effect of interest: it equals one for observations after March 2023 in which the husband’s occupation $o(H)$ belongs to the set of occupations that experienced an above-median increase in the probability of WFH between the pre- and post-pandemic periods. [^19] In the ACS sample, where data are available only at the annual level from 2017 to 2023, we define the post-period as years 2022 and 2023, due to the absence of month-level timing and in order to have a sufficient sample size to run our regressions.202020 In the ATUS sample, where we have data from 2017 to 2024, and it is a subsample from CPS, we define March 2022 as the starting point of the post-treatment period to have sufficient sample size.
The control vector $X_{i(H),t}$ includes the woman’s age, education, and number of children, as well as the husband’s education. All regressions include state-by-time fixed effects $f_{s,t}$ to account for local labor market shocks and policy variation, and occupation fixed effects $f_{o(H)}$ to absorb time-invariant differences across husband’s occupations.
When leveraging the panel structure of one of our datasets, the SIPP, we estimate a panel version of Equation (1) by including household fixed effects. As a result, we remove any control or fixed effect that is time-invariant, differencing out all unobserved and observed factors that do not vary over time within households and/or individuals.
When conducting heterogeneity analysis by age of children or other relevant categories, we interact the treatment term with the categorical variables of interest, omitting the base category, and similarly interact controls and fixed effects to allow for differential trends across groups.
We also conduct industry-level specifications, where we construct a treatment indicator based on whether the husband’s Census 2012 industry ($n(H)$) experienced an above-median increase in WFH probability. In these regressions, we replace the occupation-level treatment term in Equation (1) with the industry-level indicator interacted with the post-COVID period. Additionally, we estimate a combined occupation–industry specification that includes three interaction terms: (i) the occupation-based treatment indicator interacted with the post-COVID period, (ii) the industry-based treatment indicator interacted with the post-COVID period, and (iii) their interaction, also interacted with the post-COVID dummy. This approach allows us to separately identify the effects of occupational and industry WFH exposure, as well as any potential complementarities between the two. All specifications include state-by-time and spouse occupation and/or industry fixed effects, along with the standard set of demographic controls.
Identification Assumption
Under the standard parallel trends assumption—namely, that in the absence of the pandemic, the labor market outcomes of women whose husbands worked in occupations that experienced large increases in WFH feasibility would have followed similar trends to those whose husbands worked in occupations with minimal changes—the coefficient $\beta$ identifies the causal effect of interest. Specifically, $\beta$ captures the intra-household spillover effect of the husband’s exogenous shift in WFH probability, induced by the pandemic, on the wife’s labor market outcomes.
We conduct a range of robustness checks to support the validity of our identification strategy. These include tests for selection and group composition, the exclusion of influential occupation groups to ensure that results are not driven by specific job categories, alternative specifications of fixed effects and controls, and several placebo exercises. Across all these tests, our results remain consistent, as discussed in detail in Section 5.
It is important to note that, due to data limitations, we are unable to fully isolate spillover effects from potential direct effects. In most surveys, occupational information is only recorded for respondents who are currently employed. As a result, we cannot systematically observe or control for the wife’s own WFH exposure or estimate the interaction between the husband’s and wife’s occupational flexibility. [^21] See Appendix C for a detailed discussion of the identification challenges posed by the lack of occupational data for non-employed individuals and the implications for isolating intra-household spillover effects from potential direct effects of the wife’s own WFH exposure.
In the next section, we present estimates of the coefficient of interest ($\beta$) from this difference-in-differences specification, focusing on a range of labor market outcomes at both the weekly and annual level.
3.2 Dynamic Difference-in-Differences
To assess pre-trends and the dynamic evolution of treatment effects, we estimate a dynamic version of Equation (1). Specifically, we estimate the following specification:
$y_{i(H),s,t}=$ $f_{s,t}+f_{o(H)}\;+$ (2)
$\sum_{\begin{subarray}{c}y=2017\\ y\neq 2019\end{subarray}}^{2024}\beta^{y}\cdot 1\left(o(H)\in\{\text{High $\Delta$WFH Occupations}\}\right)\cdot 1\left(\text{year}(t)=y\right)$
$+X_{i(H),t}\delta+\epsilon_{i(H),s,t}$
where 2019 serves as the omitted (base) year. The coefficients $\beta^{y}$ capture the differential evolution of outcomes for women whose husbands were in occupations with large WFH shifts, relative to those in occupations with smaller or no changes, for each year from 2017 to 2024 (excluding 2019). All notation follows as in Equation (1).
We plot the estimated $\beta^{y}$ coefficients to visually assess the plausibility of the parallel trends assumption and to explore the dynamic pattern of the estimated effects.
4 Main Results
4.1 Baseline Difference-in-Differences Estimates
First-Stage: WFH Shock Exposure and Husbands’ WFH uptake
Our main identification strategy exploits how occupation-level shocks in women’s spouses increased women’s labor market participation by plausibly increasing spouse’s time availability for household production. Consequently, our first stage analysis consists of verifying that large shocks on occupation-level remote work options led to sizable increase in uptake of remote work for men.
Because the CPS started asking monthly about remote work on 2023 after the COVID-19 pandemic, we cannot use our pre-period sample to estimate the baseline specification. However, the CPS questionnaire asks respondents whether they used to work from home in their occupation in the months preceding the start of the pandemic. With this recall data, we are able to build a pseudo-panel dataset with a single post-Covid and pre-Covid observation per household-month. Both observations have the same covariates and vary only in their timing status and whether or not they worked from home using the actual post-Covid and recalled pre-covid responses.
In Table 1, we present our first stage results using this pseudo-panel CPS data. We implement our baseline specification from Equation (1) and an alternative panel specification that uses household- and time- fixed effects. Both specifications show a very large increase in actual uptake of remote work of around 17 to 20 percentage points in our sample of male spouses. With a pre-pandemic average WFH rate of 10%, rates have tripled in our sample households. This increase roughly matches the large increases in share of remote work job vacancies for high exposed occupations that we documented in our motivating Figure 1.
Further, we report estimates separately for the subsample of men that are employees and the self-employed. As expected, WFH take-up increased significantly more among employees, who experienced a threefold rise relative to pre-COVID levels, compared to the self-employed, whose WFH rates increased by approximately 50 percent from a higher baseline of 16 percent. It is worth remarking that even for the self-employed the increase is large and highly significant. This motivates our choice to keep these households in our sample for the main analysis.
Table 1: First-Stage Results: Effect of Occupation-Level WFH Shocks on Husbands’ Remote Work Uptake
Cross-section Panel
Spouse Restriction: All Employee Self-employed All Employee Self-employed
$Exposed_{o(i)}\times Post_{t}$ 0.203*** 0.217*** 0.087*** 0.175*** 0.190*** 0.079***
(0.020) (0.021) (0.020) (0.020) (0.022) (0.018)
Mean pre-treatment 0.106 0.099 0.161 0.106 0.099 0.161
Obs 126,240 111,135 15,049 124,549 109,577 14,702
$R^{2}$ 0.30 0.32 0.32 0.66 0.67 0.66
  • Notes: We regress men’s self-reported telework status on their treatment status, using both cross-sectional and pseudo-panel specifications. The data comes from our household sample in the 2023–2024 monthly CPS. Respondents are asked whether they currently work from home (post-COVID) and whether they did so in February 2020 (pre-COVID), allowing us to construct a two-period pseudo-panel under the assumption that respondents remained in the same occupation. In the first three columns, we estimate cross-sectional regressions that include month-year, state, and spouse-occupation fixed effects. In the final three columns, we exploit the panel structure of the recalled data by including household fixed effects instead of occupation fixed effects. For each specification, we report results for the full sample, the subset of households where the male spouse is an employee, and the subset where he is self-employed.
Effects on Labor Market Outcomes
Our hypothesis is that the expansion of remote work opportunities for men relaxes constraints on household production, both in terms of time and physical presence, thereby enabling their spouses to increase labor market engagement while raising children. Table 2 presents our main estimates from Equation (1), quantifying the effect of increased WFH exposure for husbands on their wives’ labor market outcomes. The results indicate a sizable and statistically significant increase of approximately 2.5 percentage points in women’s probability of employment and labor force participation, relative to a pre-treatment mean of 67% — equivalent to a 4% increase.
While the effects on usual weekly hours, months worked, and part-time status are less precisely estimated, we find a meaningful decline in the probability of part-time work using the March files, consistent with an increase in labor supply intensity. In addition, we estimate an 8% increase in women’s wages, suggesting potential productivity gains or access to better job matches facilitated by the reallocation of household responsibilities. To explore treatment heterogeneity, Table B.4 reports estimates where we discretize the treatment variable into quartiles of the husband’s occupation-level WFH exposure. The first quartile corresponds to occupations with the smallest changes in WFH probability, while the fourth captures those with the largest. As expected, the effects are concentrated among households where the husband’s occupation falls into the top two quartiles of the WFH exposure distribution.
Table 2: Effect of Husbands’ WFH Shock on Wives’ Labor Market Outcomes
Monthly CPS sample March CPS sample
Employed LFP Usual work Part-time Employed LFP Part-time Months Log wages
hours status Status Worked
$Exposed_{o(H)}\times Post_{t}$ 0.024** 0.027*** 0.217 0.004 0.035** 0.027* -0.030* 0.120 0.078**
(0.011) (0.010) (0.197) (0.008) (0.015) (0.014) (0.016) (0.095) (0.036)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.665 0.708 0.247 10.897 9.992
Obs 252,309 252,309 162,882 161,095 38,637 38,637 25,920 25,920 25,920
$R^{2}$ 0.12 0.12 0.06 0.06 0.12 0.13 0.07 0.07 0.17
  • Notes: This table reports estimates from Equation (1) using the 2017–2024 CPS sample of dual-parent households with children of child-rearing age. The dependent variables include weekly labor market outcomes—employment, labor force participation, usual hours worked, and part-time status—drawn from the basic monthly CPS files, as well as annual outcomes—employment, labor force participation, months worked, part-time status, and log wages—drawn from the March CPS (ASEC) supplement. All intensive margin outcomes are defined only for employed individuals and are coded as missing for those not working. The treatment indicator is defined at the occupation level based on the change in the probability of WFH between the pre-pandemic period (2019) and the post-pandemic period (2023–2024), as described in Subsection 2.3. All regressions control for the wife’s age, education, and number of children, as well as the husband’s education. Specifications include state-by-time and husband-occupation fixed effects. Standard errors are clustered at the occupation level.
Table B.5 presents analogous results using the husband’s industry-level WFH exposure as the treatment variable. [^22] See Figure A.1 for industry-level WFH trends before and after COVID-19, analogous to the occupation-level trends shown in Figure 1. Specifically, we construct an industry-level WFH measure using job postings data aggregated at the month-year–industry level and define treatment as being above the median change in WFH probability between the pre- and post-pandemic period. The resulting estimates closely mirror those based on occupational exposure: wives of men in high-WFH industries experience an increase of approximately 2 percentage points in both employment and labor force participation.
In Table B.6, we include both the occupation- and industry-level WFH shocks, as well as their interaction, to disentangle the relative importance of each source of variation. The results indicate that the occupation-based measure primarily drives the effect. The coefficient on industry exposure is positive but not statistically significant, while the interaction term is near zero, suggesting no meaningful complementarity between the two dimensions of exposure.
Effects on Wife’s WFH Probability
The increase in women’s labor supply at the extensive margin may be accompanied by changes in their working arrangements. In particular, if husbands become more likely to work from home, as shown in Table 1, this could influence wives’ likelihood of taking jobs that offer remote work options. Such a response may occur either directly, if spouses value spending more time together (Blundell et al., 2018), or indirectly, as women entering (or re-entering) the labor market may prefer remote over in-person jobs due to documented preferences for shorter commuting time (Le Barbanchon et al., 2021).
We explore this possibility using our CPS pre- and post-Covid pseudo-panel in Table B.7 in the Appendix. In a similar fashion to the main employment outcomes, we estimate the effect of husbands’ WFH shock on their wives’ probability of working from home. As with the first stage analysis, the pseudo-panel allows us to estimate both the cross-section and panel specifications. The full sample estimates show in both specification show that women’s WFH rates have increased around 10 pp, or a 75% increase with respect to the pre-Covid average WFH rate. The estimates are smaller for households where husbands are self-employed - around 6 to 7 pp or 40% increase - but they are still highly significant.
In the last three columns of Table B.7, exploiting our pseudo-panel structure, we try to break apart the correlation between both spouses’ treatment status when estimating our treatment effects. To do this, we include interactions between post-covid dummy and (i) women’s treatment status and (ii) women-and-men combined treatment status to the panel specification. This augmented specification reveals that 2/3 of the effect on women’s probability of remote working is driven by households where both spouses were working in occupations that experienced large increases in WFH vacancy rates. Moreover, women’s own treatment status also drives a large increase in WFH rates of nearly 16 pp, or over double the pre-Covid averages.
Heterogeneity by Age of Children
It is sensible to believe that results would differ depending on how young children are within our sample, and whether parents have had previous parenting experience. In Table B.8, we examine this possibility by estimating heterogeneous effects by age of the youngest child in the household. We interact our treatment coefficient with a dummy for whether the youngest in the household is between 5 and 8 years old. In panel A, we estimate this specification on all parents in the household. In panels B and C, we restrict the sample to households with two or more children and households with only one child, respectively.
Overall, we find limited evidence of systematic differences in labor market responses between households with and without toddlers. One exception is that mothers in households with toddlers exhibit a significant increase in employment and labor force participation. In the Monthly CPS sample, these effects are concentrated among households with two or more children, where the estimates are both larger in magnitude and statistically significant; by contrast, the corresponding estimates for households with only one child are smaller and not significant. The March CPS sample shows the opposite pattern. There is also some indication that women with toddlers experience higher wage gains, while women without toddlers and with one child work significantly more hours. Despite these differences, we do not find robust or consistent evidence of heterogeneity in women’s labor supply responses by the age of children.
Wives’ Response to Own WFH Exposure
To benchmark the intra-household spillover effects, we compare the effect of husbands’ WFH shocks on their wives’ labor market outcomes to the direct effect of wives’ own WFH shocks—that is, we compare cross-elasticities to own-elasticities. [^23] Due to survey design, we are unable to estimate extensive-margin own-elasticities. Occupation is only reported for individuals who are currently employed or in the labor force, which makes occupation-based WFH exposure mechanically collinear with participation and employment status. As a result, it is not possible to separately identify the causal effect of own WFH exposure on the decision to enter the labor force or become employed. Table B.9 reports estimates from specifications where treatment is defined using the wife’s own occupation-level WFH shock, and outcomes are her intensive measures of labor supply. The results indicate no statistically significant response: wives do not appear to increase or adjust intensive margins of labor supply in response to their own WFH exposure.
4.2 Dynamic Effects and Event Study Analysis.
We implement the dynamic difference-in-differences specification described in Equation (2) to test for pre-trends and provide year-specific estimates of our main effects. Figure 2 displays the annual coefficients for weekly employment and labor force participation. Reassuringly, we find no evidence of differential trends prior to the pandemic, lending credibility to the parallel trends assumption and reducing concerns about selection into treatment based on unobservables. Beginning in 2022, the estimated effects stabilize at around a 2 percentage point increase in women’s extensive-margin labor supply, consistent with a persistent intra-household response to the husband’s increased WFH exposure and with the results found in the baseline DiD.
To assess whether these effects are driven by improvements in treated households or declines in control households, Appendix Figure A.2 plots adjusted labor market trends separately for treated and control groups. The figure shows that both groups experienced declines in employment and labor force participation during the early pandemic period. While women in the control group gradually returned to their pre-pandemic levels, those in the treated group experienced a sustained increase in labor market outcomes that exceeded pre-COVID levels starting in 2022.
Appendix Figure A.3 presents dynamic estimates and adjusted trends for additional outcomes: weekly hours worked, probability of part-time work, and log wages. While the estimates for hours and part-time status are noisier, we observe a modest increase in hours worked in 2021 and 2022, and a decline in the likelihood of part-time work for treated women. Wage effects appear more robust: consistent with our baseline results, we find that women in treated households experience an average increase in log wages of approximately 10% with respect to 2019.
Figure 2: DiD Estimates of the Impact of Husbands’ WFH Shocks on Wives’ Labor Market Outcomes
(a) Weekly Employment
(b) Weekly Labor Force Participation (LFP)
Notes: Each panel plots coefficients from a dynamic difference-in-differences specification comparing women whose husbands were in occupations with large WFH shocks to those in occupations with smaller changes. The omitted year is 2019. All regressions include state-by-time and husband-occupation fixed effects, and control for wife’s age, number of children, and education levels of both spouses. Confidence intervals are shown at the 95% level and are based on standard errors clustered at the husband-occupation level.
5 Robustness and Validation
Group Composition and Selection.
We conduct a series of robustness checks to assess whether our main results are driven by changes in group composition between treated and control households before and after the pandemic. Given the cross-sectional nature of the CPS, one potential concern is that our findings may not reflect intra-household adjustments—for example, a woman increasing her labor supply in response to her husband’s WFH shift—but instead reflect changes in assortative matching. Specifically, it is possible that, post-COVID, men in high–WFH-shock occupations are more likely to be partnered with women who are already more attached to the labor force.
We address this concern by examining changes in assortative matching over time. Figure A.4 in the Appendix provides suggestive evidence that changes in spousal composition are unlikely to be driving our main results. While there is clear evidence of assortative matching, i.e., households in which the husband is in the highest quartile of WFH exposure are more likely to have wives in similarly high-exposure occupations, this pattern appears stable over time. Specifically, the distribution of wives’ exposure conditional on their husbands’ treatment quartile does not shift meaningfully between the pre- and post-COVID periods. As discussed in Subsection 2.3, occupational WFH exposure is strongly correlated with wives’ income and labor force attachment. The stability of the matching pattern across periods is therefore reassuring and suggests that selection into treatment via assortative matching is not biasing our estimates. In addition, we leverage the ACS to examine potential changes in key demographic outcomes - namely, marriage, divorce, and fertility. Table B.10 reports estimates from specification (1), where the dependent variable is an indicator for whether the woman got married, divorced, or had a child within the 12 months preceding the interview date. Across all outcomes, we find no statistically significant effects of the husband’s WFH shock, suggesting no differential selection into or out of marriage, nor into fertility, between treated and control groups following the pandemic.
Another plausible concern is that the composition of households with young children may have changed after the pandemic. For instance, if households that chose to have children post-COVID differ systematically in characteristics such as income, education, or labor market attachment, our estimates may capture selection effects rather than causal impacts. To assess this possibility, Figure A.5 plots changes in the probability of having a child and in the number of children across quartiles of husbands’ WFH exposure, before and after COVID. The results reveal no systematic differences in fertility trends by treatment status, alleviating concerns that selection into parenthood is driving our main findings. Consistently, Table B.11 presents regression estimates based on Equation (1), showing that the husband’s WFH shock has no statistically significant effect on fertility outcomes.
To address concerns about endogenous changes in group composition over time, we implement reweighting strategies that adjust the post-COVID sample to match the distribution of pre-COVID covariates. We estimate a single logit model for the probability of an observation belonging to the pre-COVID period, using spouse occupation, wife’s age and education, state, and race as predictors. [^24] The predicted probabilities are obtained from the following specification: $\Pr(\text{Pre-COVID})=f_{s}+f_{o(H)}+f_{e}+f_{a}+\gamma\cdot\text{Non-White}+\varepsilon,$ where $f_{s}$, $f_{o(H)}$, $f_{e}$, and $f_{a}$ denote state, husband’s occupation, education, and age fixed effects, respectively. We then plot the predicted probabilities for pre- and post-COVID observations to verify common support (see Figure A.5). We construct DiNardo et al. (1995) (DFL) weights as p/(1p)p/(1-p) and inverse-propensity score (IPW) weights as $1/p$, where $p$ is the predicted probability of being in the pre-COVID period. [^25] Both sets of weights are normalized to sum to one and scaled by CPS sampling weights. As shown in Table B.12, the reweighted estimates remain consistent with our baseline results, alleviating concerns that shifting sample composition drives our findings.
Treatment Definition Robustness
A potential concern is that our baseline treatment definition — based on changes in the average WFH vacancy rate between the post-covid period in 2023 and 2024 and the pre-covid period in 2019 at the Census 2010 4-digit occupation level — may be driving the results. To address this, we examine the robustness of our findings to a range of alternative treatment definitions, as Table B.13 presents. We consider different approaches for computing changes in WFH vacancy rates before and after the pandemic, employ alternative occupational classifications, and construct the treatment variable using a different WFH index.
The estimates reported in Panel B confirm that smoothing the share of remote work job vacancies using a three-month moving average does not affect the treatment assignment, and henceforth, the resulting coefficients. Similarly, the results in Panels C, E, F, and G, which use alternative classification schemes and different measures of WFH change, yield estimates that are closely aligned with those obtained under the baseline treatment. Panel D, which defines treatment using the Teleworkability Index à la Dingel and Neiman (2020) as outlined in Subsection 2.3, yields slightly larger effect sizes but preserves the direction and statistical significance of the key estimates. Although the earnings coefficient loses significance in this specification, the overall pattern remains consistent. Panel H does the same but using the SOC 3-d classification. The resulting estimates are nearly indistinguishable from those based on the benchmark treatment.
Taken together, these findings indicate that our results are not sensitive to how treatment is defined, and that the estimated effects are robust across a range of reasonable alternatives.
Specification and Sample Robustness
We test in several ways the robustness of our main women employment results to minor changes in the specification and sampling choices. First, we test the robustness to the choice of fixed effects. By design, the correct specification should include at least an after Covid-19 dummy and a treatment dummy. In our baseline specification, we generalize these to month-year-by-state fixed effects and spouse occupation fixed effects to gain more precision. In Table B.14, we make minor alterations to the fixed effect structure in our main estimations for the CPS sample. We change the occupation group definitions and the level of interactions between state fixed effects and others. By comparing all other results to Panel E - our baseline specification - a clear pattern emerges in that coefficients do not change significantly with the fixed effects interactions or with the level at which we categorize occupation groups.
Second, we test how the CPS results on women’s extensive margin labor supply hold up when we change the specification of demographic controls, while keeping the fixed effects unchanged. In Table B.15, we estimate the baseline cross-sectional DiD specification for employment and labor force participation, starting with a model that includes no controls (Column 1) and then sequentially adding controls until reaching the full control set in Column 5. The point estimates remain stable throughout, suggesting that our results are not sensitive to the inclusion of any specific control. We further test robustness in Table B.16 by altering how we account for the age composition of children in the household. Specifically, we include the age of the youngest and oldest child, both as continuous variables and grouped into five-year bins. Once again, results remain consistent, indicating that our findings are robust to alternative specifications of child age controls.
Third, we test the sensitivity of our results to changes in how we define the post-Covid window. In our baseline CPS specification, we deliberately and conservatively omit the sample between March 2020 and March 2023 to partial out any temporary confounding effect of the Covid-19 pandemic in our estimates. The argument is that by March 2023 the pandemic was officially over and important events for career outcomes such as school and daycare closures had been phased out. However, the relevant shocks to our estimates might have phased out much earlier and dropping as much sample may not be strictly necessary. In Table B.17, we look at how our estimated effects of women’s labor market outcomes vis-à-vis their spouses’ WFH occupation shocks change when progressively push-back the post-Covid timeframe. Panel B shows our baseline choice, and from there we push the end-window back by three months at a time. Results are robust all the way until the last panel, where we only drop observations between March 2020 and September 2021.
Fourth, we examine the robustness of our results to additional sample restrictions. In Table B.18, we re-estimate the main occupation-level effects on women’s outcomes using only households where the male spouse is an employee (i.e., excluding the self-employed). The results remain largely unchanged, supporting the inclusion of self-employed men in our main specification. In Table B.19, we test sensitivity to household income by trimming the top 1% and 5% of income earners in Panels B and C, respectively. Results remain stable: estimates for extensive margin outcomes are unaffected by either trimming, and the log wage coefficient is robust to the exclusion of the top 1%.
Fifth, we assess the sensitivity of our results to influential occupation groups by conducting a leave-one-out analysis, systematically dropping each major occupation category and re-estimating the effects. Table B.21 reports the bootstrapped average treatment effects, standard errors, and p-values across these iterations. This exercise allows us to evaluate the persistence of our findings and whether they are driven by any particular broad occupation group. For the monthly CPS sample, the results indicate that no single occupation group statistically drive the observed effects on women’s employment or labor force participation. In the March CPS sample, the estimated effects on employment status and earnings remain robust to this exercise. On the other hand, the significant effects on labor force participation and part-time status observed in the baseline results are no longer statistically significant once we consider the bootstrapped p-values. Despite this loss of precision, the bootstrapped means and standard errors for these outcomes mimic the estimates Table 2 yields, suggesting that the direction and magnitude of the effects remain stable.
Finally, we consider the possibility that the occupations classified as treated or control only under the treatment definition from the job postings data might be driving the effects. To ensure that our findings are not driven by occupations that other teleworkability indicators would classify differently, we restrict the sample to occupations with a consistent treatment status to the treatment definition based on the Teleworkability Index à la Dingel and Neiman (2020). Table B.20 presents the results. The estimated effects generally increase in magnitude across most outcomes, with the exception of wages. Notably, the coefficient on usual work hours becomes statistically significant. These patterns hold when the treatment is redefined using the 2018 SOC 3-digit classification. Overall, these findings indicate that our main results are not sensitive to discrepancies in the treatment status of occupations between the two of the most reliable sources of WFH data.
Placebo tests
We conduct several placebo tests to bolster the validity of our identification strategy and our theoretical framework. First, we argued that wives affected by their husbands’ WFH shocks are those with children of child-rearing age. Indeed, employment gaps are virtually non-existent between spouses that have no children, or whose children are old enough to be in school and/or take care of themselves. We can verify this hypothesis using the full CPS data. In Table B.22, we run a placebo test by estimating our baseline cross-section specification in the sample of households outside our restriction on the age of youngest child. Across the board, we verify that women’s monthly labor market outcomes are unaffected by their spouses’ WFH occupation-level shocks when their children are as old as teenagers (12 to 18 years old), are young adults (over 18 years old), or when they have no children at all.
Second, our preferred specification uses husbands’ changes in their occupation WFH rates following Covid-19 as an exogenous shock that impacts their wives’ labor market decisions. We expect that the nationwide WFH shocks only affect female spouses in households with young children, since they are usually the ones who bear the majority of child-rearing responsibilities. Men’s labor supply is close to one on average, as shown in the descriptive statistics from previous sections. In order to test this hypothesis, in Table B.23 we invert the estimation procedure and estimate the effect of female spouses’ WFH shocks on their husbands’ labor market outcomes using the monthly CPS data. By and large, we find no effect on men’s labor market outcomes, whether we look at women’s occupation and job industry shocks.
Third, we conduct a standard timing falsification test by coding the timing of the Covid shock as starting on March 2018 - three years before the pandemic began - and estimating the baseline specification on the 2017-2019 CPS subsample. This data has no influence from the post-Covid era, and since the Covid shock is largely unexpected, we would expect no effects to appear on these estimations. We present the results from this placebo test in Table B.24, which confirm that there are no significant employment, labor force participation or wage effects occurring around this artificial time of treatment.
Evidence from the ACS
We replicate our main results on wives’ labor market outcomes using the ACS sample (2017–2023), described in Section 2. While the ACS does not include all CPS outcomes, [^26] For example, it lacks a measure of annual labor force participation, and annual outcomes refer to the 12 months preceding the interview rather than the calendar year. We also observe only the interview year (not month), so we define the post-treatment period as 2022–2023 and include year fixed effects. it offers the advantage of a substantially larger sample size. Table B.25 confirms our main findings: treated wives experience a 2 percentage point increase in employment and labor force participation, work more hours per week, are 3 percentage points less likely to work part-time, and earn roughly 5% higher wages. When we examine heterogeneity by the age of the youngest child in Table B.26, we find that, much like the CPS results, the employment effects in the women ACS sample also appear to be driven by households with toddlers aged 0–4.
Evidence from the SIPP
Finally, we replicate several of our main employment results and analysis using the SIPP data. Panel A of Table B.27 shows that baseline estimates [^27] Since the SIPP only has data for reference years up until 2023, we moved the post-Covid period to after March 2022, and remove data between March 2020 and March 2022 for estimations. for the effect of husbands’ WFH occupation shocks on their wives’ labor market outcomes are positive in the SIPP cross-section. Moreover, point estimates using the SIPP data are larger than our CPS and ACS results. We are unable to test whether this difference is larger, since all estimates come from entirely different sources. Nevertheless, the size of the standard errors would suggest that the differences may not be statistically significant. In Panel B, point estimates suggest that employment effects are larger for households with younger children - in the range of 0 to 4 years old - but the differences are not statistically significant.
As with the CPS and ACS, we perform a heterogeneity analysis based on the age of the youngest child, separately for households with multiple children and with only one children. These estimations are shown in Table B.29, and overall these are closely aligned with the CPS heterogeneity results: We find some evidence that households with toddlers experienced significant effects, whereas non-toddler households experienced non-significant effects - as seen by the test on the sum of the coefficients. This is particularly true in households with more than one child. However, because the interaction coefficients are not significant, we cannot interpret this as conclusive evidence of heterogeneous results.
In addition to our cross-section specification, the panel component of the SIPP data allows us to estimate a panel regression. Specifically, we replace our baseline fixed effects with month-year and household-level fixed effects. We present results for several samples in Table B.28.282828In addition to sampling households of dual parent households with children aged 0 to 8 years old, we restrict the sample to households that have not switched treatment status across different panel interviews. In Panel A, which uses all data from the 2017 to 2023 reference years, we find positive and large estimates for employment, labor force participation, and wage effects for women; however, these estimates are not statistically significant except for wages. The results similarly hold when we use data from the 2020 SIPP panel, which follows households from reference years 2019 to 2022, and thus is the only panel to cover same household before March 2020 and after March 2021.292929Due to logistical issues, the SIPP did not follow through the sample of new households that joined the sample in SIPP 2019. In panels C and D, we estimate our sample on balanced 2019-2022 and 2019-2021 panels, respectively. [^30] For panel D we change the post-Covid window to data after March 2021. Estimates are non-significant, which we argue is due to loss in power as we only observe a very limited number of different households in the balanced panels.
In Table B.30, we replicate and consolidate our main placebo and falsification tests developed for the CPS sample in our SIPP sample. These effects include estimating the effects of women’s WFH occupation shocks on their male spouses’ labor market outcomes, estimating women’s labor market effects on a sample without children or whose youngest child are teenagers, and moving the Covid-19 shock to March 2018 for a falsification test. The falsification test for households without children and the artificial timing of Covid-19 (Panels C and D) show the results we expected. The latter is specially reassuring as it provides evidence against pre-trends in the SIPP data. The labor employment effects for husbands (Panel A) are weakly significant and less than a half of the treatment effect for wives in the SIPP sample. However, women whose youngest are teenagers-to-adults see large effects, which goes against our expectations.
Summary
Taken together, these robustness and falsification checks support the validity of our identification strategy. The estimated effects are not driven by differential selection, sample restrictions, or alternative treatment definitions. Main estimates across different nationally representative datasets are similar in magnitude and sign.
6 Mechanisms
Understanding the household mechanisms that enabled women to increase their labor supply when their husbands became more likely to work from home is central to interpreting the overall effects of the WFH shock. We posit that our large employment effects for women with young children, to some extent, can be explained if we also observe changes in the household production function, and more specifically, in the burden of child-rearing. This could occur through the following channels: (i) a substitution channel, where greater time spent at home of the husband leads to an increase of his involvement in childcare, potentially substituting not only for the wife’s time but also for alternatives such as paid daycare; or (ii) a multitasking channel, where women provide childcare while simultaneously engaging in other activities, such as working from home or performing household tasks, effectively changing their childcare “production function”.
To explore how the intra-household allocation of time adjusts in response to the WFH shock, we estimate Equation (1) using variables from the ATUS and SIPP that capture both the intensive and extensive margins of labor supply and household production. [^31] Since the ATUS and SIPP have fewer observations, we moved the post-Covid period to after March 2022, and remove data between March 2020 and March 2022 for estimations. We still lose a lot of precision due to sample size. Our goal is to assess how time spent on domestic tasks and childcare is redistributed within households, enabling the observed increase in labor market participation among women in treated households. [^32] For the ATUS, we restrict the sample to respondents who completed the time-use diary on a weekday and non-holiday day, in order to capture household dynamics on a typical workday.
Time-Use in ATUS
First, the increase in time allocated to market work is consistent across spouses and aligns with the labor supply effects documented in Section 4. As shown in Table 3, treated women increase the share of their day spent working by 7.5 percentage points, i.e., equivalent to nearly two additional hours. [^33] A full day includes 24 hours This rise is driven entirely by an increase in main job hours, with significant gains only from working from home. Although the estimate for time spent working outside the home is not statistically significant, both its direction and the pattern shown in Figure A.7 suggest an increase during the first two years of the post-treatment period. We find no significant changes in time spent commuting or in hours worked at a secondary job.
In contrast to their wives, husbands experience a substantial reallocation in their work location, consistent with being the directly treated group. They increase time working from home by 4.7 percentage points and appear to reduce time working outside the home. Although the latter estimate is noisy due to limited sample size, Figure A.7 shows a negative and relatively stable effect across all post-treatment years. These shifts in work location are persistent over time, unlike the more transitory patterns observed for women. Notably, men do not exhibit any change in total hours worked. However, treated men consistently report reduced commuting time, a pattern strongly aligned with the nature of the WFH shock, and an increase in time spent on secondary jobs.
Second, treated households exhibit meaningful shifts in the allocation of household and childcare responsibilities. On the household production side, women spend significantly less time shopping, while time spent on general household tasks such as cooking and cleaning shows no statistically significant change. Although these estimates are not precise, they consistently point in a negative direction, as shown in Figure A.8. Regarding childcare, although total time spent caring for children does not change significantly, there is a clear shift in how childcare is provided. Women in treated households reduce the amount of time spent on childcare as a primary activity by 2.7 pp, while not changing the time spent on childcare while engaged in other primary non-work activities. This suggests that women are more likely to work and to work longer hours, and one way to cope with this is to reduce the time spent on childcare as the main activity. As a result, secondary childcare becomes more prominent, with its share of total childcare time increasing by 8.4 pp relative to a pre-treatment mean of 64.5 percent.
Husbands, on average, do not exhibit significant changes in total time spent on household production or overall childcare. However, as shown in Figure A.8, their time spent on primary childcare increases significantly in most post-treatment years, while time spent on secondary childcare remains unchanged. In particular, treated husbands are no more likely to report engaging in childcare while working, which helps explain the decline in secondary childcare as a share of their overall caregiving. These patterns point to a key mechanism: as women reduce time spent on primary childcare, men appear to partially compensate—a finding consistent with the substitution channel. That said, the extent of this compensatory behavior is limited and does not fully offset the reductions observed among women.
Third, regarding personal time, women in treated households experience no significant change in time spent on personal care activities, such as sleeping and grooming. In contrast, they exhibit a 1.9 pp increase in time devoted to leisure activities, suggesting that the shift in labor supply does not fully crowd out their discretionary time. For husbands, however, there is a significant reduction in leisure time, amounting to 2.4 percentage points.
Table 3: Effect on Time Spent in Labor and Household Supply - Share of the Day
Time Spent in Labor Production Time Spent for Personal Care and Leisure
Working Main Work Main Work Main Work Secondary Work Commuting Personal Care Leisure
From Home Outside Home
Wives:
$Exposed_{o(H)}\times Post_{t}$ 0.075*** 0.074*** 0.054*** 0.021 -0.000 0.000 -0.013 0.019*
(0.018) (0.018) (0.014) (0.017) (0.001) (0.002) (0.011) (0.011)
Mean pre-treatment 0.172 0.170 0.025 0.145 0.001 0.012 0.374 0.112
Obs 7,989 7,989 7,989 7,989 7,989 7,989 7,989 7,989
$R^{2}$ 0.85 0.84 0.85 0.84 0.96 0.83 0.83 0.87
Husbands:
$Exposed_{o(H)}\times Post_{t}$ 0.024 0.023 0.047*** -0.024 0.005* -0.009** -0.001 -0.024*
(0.017) (0.017) (0.011) (0.019) (0.003) (0.004) (0.012) (0.014)
Mean pre-treatment 0.317 0.309 0.028 0.280 0.007 0.029 0.350 0.116
Obs 8,186 8,186 8,186 8,186 8,186 8,186 8,186 8,186
$R^{2}$ 0.85 0.85 0.88 0.85 0.82 0.83 0.84 0.85
Time Spent in Household Production
Housework Cooking Shopping Primary Secondary Total Total Prop. Secondary CC
Childcaring CC while working Secondary CC Childcaring in Total CC
Wives:
$Exposed_{o(H)}\times Post_{t}$ -0.001 -0.004 -0.006* -0.027** 0.022 0.025 -0.002 0.084***
(0.009) (0.006) (0.003) (0.011) (0.046) (0.021) (0.022) (0.029)
Mean pre-treatment 0.093 0.045 0.016 0.106 0.123 0.228 0.334 0.645
Obs 7,989 7,989 7,989 7,989 4,691 7,989 7,989 7,849
$R^{2}$ 0.86 0.83 0.85 0.84 0.92 0.84 0.84 0.85
Husbands:
$Exposed_{o(H)}\times Post_{t}$ 0.003 0.002 -0.004* -0.004 0.028 0.001 -0.003 -0.049
(0.009) (0.003) (0.002) (0.008) (0.027) (0.019) (0.020) (0.042)
Mean pre-treatment 0.041 0.016 0.007 0.051 0.042 0.127 0.177 0.649
Obs 8,186 8,186 8,186 8,186 7,113 8,186 8,186 7,568
$R^{2}$ 0.85 0.81 0.84 0.86 0.87 0.83 0.83 0.86
  • Notes: We use data from the the 2017-2024 ATUS panel datasets. We show results for the time spent in labor, household and personal activities available in ATUS. Time-use variables are recorded by one member in the household. Then, as Section  2 points out, we have two draws from ATUS: the men-respondent sample and the women-respondent sample. While the former has data on the time spent doing activities by husbands, the latter will have for wives. There is a first panel that has dependent variables that capture time spent in labor production, as a share of the whole day: working, working at their main work, working at their main work from home and outside home, working at their secondary work and commuting to work. Those that do not work should have a 0 imputed in all working variables. There is a panel with dependent variables related to personal care (i.e. sleeping or taking a shower) and leisure (ie. going to religious events or sport activities). Lastly there is a household panel with dependent variables that capture the time spent doing the following activities: housework, cooking, shopping, primary childcaring (report doing childcaring as main activity), childcaring while working, total childcaring as secondary activity (childcaring while doing primary activities), total childcaring (the sum of primary and secondary childcaring) and the share of secondary childcaring on total childcaring. We restrict our sample to weekdays and non-holiday days. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the occupation level.
The substitution in household production between spouses varies with the age of the youngest child. Table B.31 shows, to a certain degree, that when the youngest child is no longer a toddler, women are more likely to work longer hours - particularly outside home - and reduce time spent in household tasks. Moreover, secondary childcare increases in relative importance.
On the other hand, husbands in these households exhibit a significant reallocation of where their main job takes place, without changing their total working time. Specifically, the likelihood of working from home increases by 8.7 percentage points, accompanied by a 10 percentage point decline in working outside the home. Consistent with this shift, husbands report more time spent on household chore –although not fully significant as the multiple hypothesis test suggests – and a significant 5.5 percentage point rise in total childcare time, largely driven by primary childcare, though the latter component is not statistically significant. [^34] Those magnitudes come from the sum of the main and interaction effects. Nonetheless, these adjustments do not fully offset the decline in time contributed by their wives. In addition, husbands spend significantly less time on childcare while working.
In contrast, households with toddlers exhibit a different pattern. Among women, there is no significant change in overall working time. However, the share of time spent working from home increases by 6.9 percentage points, while work outside the home declines by 4.7 percentage points, consistent with a reduction in commuting time. These women also allocate more time to household tasks such as shopping, but show limited changes in childcare, with the exception of a significant rise in time spent caring for children while engaged in other activities.
Among husbands in toddler households, total working time increases, driven by a shift towards working from home. They are also significantly less likely to hold a secondary job. In terms of household production, there is no significant change in time spent on most domestic tasks, except for cooking, which increases. As expected given the age of their youngest children, these husbands spend more time caring for them while working. However, they do not report significant changes in either primary or total childcare time.
To reinforce the validity of the findings in this section, we perform several robustness checks. We begin by evaluating the comparability between the two ATUS samples and the CPS, given that the ATUS samples are non-overlapping subsamples drawn from the CPS. Table B.32 presents the first and second moments of the distribution for key characteristics of husbands, wives, and households across the CPS sample and both ATUS subsamples. The two ATUS samples exhibit nearly identical means and standard deviations across all variables, with the exception that individuals in the men-respondent sample are more likely to be classified as treated. Additionally, there are differences in the second moment of wages between these subsamples. The CPS sample displays modest differences relative to the ATUS samples, particularly in the likelihood of treatment and in measures of household income and education. These patterns suggest that the ATUS subsamples tend to capture relatively wealthier and more educated households.
A further concern is that the estimated effects might be driven by specific subsamples whose responses to the WFH shock diverge from those identified in the main CPS analysis in Table 2. We address this in two ways. First, both Table 3 and Figure A.7 show that the increase in women’s employment following the WFH shock is statistically significant and persistent. Second, despite the smaller sample sizes in ATUS, Table B.33 shows that the estimated effects on wives’ labor market outcomes are directionally consistent with those in the main CPS sample.
These exercises yield two key implications. First, although the women- and men-respondent samples are drawn independently, their similarity implies that results based on one can be generalized to the other. In other words, if time-use diaries were available for both spouses within the same household, we would expect patterns similar to those observed across the separate respondent samples. Second, the alignment between the ATUS and CPS results suggests that the findings in the ATUS subsamples can, to some extent, be extrapolated to the broader CPS sample.
Finally, we conduct a standard placebo test by assigning the onset of the post-treatment period to March 2018. Table B.34 reports the results. We find no significant changes along all the set of variables, further supporting the credibility of our baseline results.
Childcare in SIPP
We are also able to test the mechanisms using the SIPP childcare topical modules, which are available for all years in our 2018-2024 SIPP panels, and ask reference parents several (extensive margin) question about any agents involved in childcare activities for their children.
We present some results validating both mechanisms in the top panel of Table 4. We estimate our baseline cross-section specification on the 2018-2024 SIPP panel datasets on a battery of childcare related variables. Our results show that in households shocked by the occupation-level WFH treatment, when the child’s mother is working, the spouse is 6.4 pp more likely to be involved in childcare, or roughly a 13% increase over the baseline. The woman herself is not any more likely to be involved in childcare by 1.1 pp while she is at work. This pair of results are most consistent with the substitution hypothesis.
We also find that households treated by the WFH shock are significantly less likely to recur to the child’s grandparent - or any other relatives - or a daycare for day-to-day childcare. Further, households are 7.3 pp less likely to have paid for childcare and, although not significant, overall expenditure in childcare decreased by some 16%. This is consistent with the finding that male parents are more available for childcare, once again, consistent with the substitution channel hypothesis.
To validate the identification strategy underpinning these past estimations, we run a placebo test analogous to Table B.24, where we estimate the same specification on the 2017-2019 SIPP subsample and using artificial timing for the start of treatment (March 2018). We present this placebo test on the SIPP childcare and child activities outcomes on Table B.35. The falsification tests are largely reassuring: we find no significant effects on parental involvement in childcare during the wife’s working hours, nor on household expenditures on childcare. However, as a caveat, we note that the estimated effects on childcare provided by grandparents and other relatives are negative.
Table 4: Effect of Husbands’ WFH Occupation Shocks on Childcare Time and Child-Related Activities
Childcare by Childcare by Childcare by Childcare by Childcare by Childcare Paid for Log Childcare
Wife while Spouse while Grandparent Relative Non-relative by Daycare Childcare Expenditure
Working Wife Works
A. Entire sample
$Exposed_{o(H)}\times Post_{t}$ 0.004 0.064* -0.001 0.023 -0.009 -0.053* -0.073** -0.165
(0.032) (0.036) (0.031) (0.032) (0.021) (0.028) (0.031) (0.117)
Mean pre-treatment 0.330 0.485 0.416 0.454 0.186 0.266 0.484 4.710
Obs 60,264 60,264 85,379 85,379 85,379 85,379 68,047 32,910
$R^{2}$ 0.15 0.16 0.15 0.14 0.13 0.18 0.19 0.33
  • Notes: We use data from the the 2018-2024 SIPP panel datasets. We estimate our baseline cross-section specification which has state-by-month-year and spouse occupation fixed effects. We show results for childcare variables available in the SIPP, which ask households with minor of age children if in a typical week, the following are involved at least one day in looking after children: The wife and her spouse (when the wife is working), a grandparent, another relative (includes grandparent), a daycare. We also include a dummy on whether the family pays for daycare, and self-reported log expenditure in childcare. All regressions include fixed effects for state-by-time, and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Children’s Well-being
Finally, it is plausible that children may be worse off given that our main results suggest that women in our treatment sample are more likely to be working, spend more time at work and, as found in this section, there is an overall reduction in the household’s time spent child-caring as a primary activity. Although we cannot fully measure well-being with our current datasets, the SIPP has a handful related variables such as days that parents have dinner or go out with their child and the children’s involvement in extracurricular activities. In Table B.36, we estimate our baseline specification on these outcomes, and we find no significant changes for these variables. This evidence suggests that – to the extent we can measure well-being – children are not worse off by changes in household production and labor supply.
Summary
Overall, these results underscore the role of intra-household reallocation in explaining the observed improvements in women’s labor market outcomes. The substitution channel is the one that aligns more closely with our findings: women reduce their primary childcare time in order to increase their working time, and are modestly compensated by husbands. In addition, men are more likely to be at home and there is a decline in reliance on extended family or formal daycare services. In contrast, we find little support for a multitasking channel. These changes are not coming at the cost of any worsening the children’s well-being from the related variables we can observe.
7 Conclusion
In this paper, we provide novel evidence on the intra-household labor market and household production spillovers induced by large and permanent changes in WFH arrangements of male spouses in dual parent households. Exploiting plausibly exogenous variation in the post-pandemic shift in remote work options across occupations in the U.S., we show that increases in husbands’ WFH exposure have led to significant gains in their wives’ labor market outcomes - namely higher employment, labor force participation, and wages. These effects are concentrated among households with children of child-rearing age and are not driven by changes in group composition, assortative matching, and are widely robust to specification changes and several falsification test. Our analysis of time use and childcare arrangements indicates that the increase in women’s labor supply is partly facilitated by an intra-household reallocation of time. In particular, we find evidence of modest substitution in childcare responsibilities between spouses: women reduce the time they devote to primary childcare, while husbands slightly increase their involvement.
We highlight the importance of household-level dynamics when evaluating the labor market implications of job amenities like WFH. Our findings suggest that access to remote work arrangements for one partner can alleviate time constraints for the other, with implications for gender equality in labor force attachment and earnings. In future research, we hope to explore larger datasets covering the universe of U.S. households to better disentangle mechanisms underlying our results and follow-up on career progression for women entering the labor force.
Appendix
Appendix A Figures
Figure A.1: Evolution of WFH by industry (2 digits).
Notes: Author’s elaboration using WFH job postings data from Hansen et al. (2023). Classification based on NAICS 2022 2-digits.
Figure A.2: Adjusted Labor Market Trends by Treatment Group, 2017–2024
(a) Weekly Employment
(b) Weekly Labor Force Participation (LFP)
Notes: Each panel plots adjusted yearly trends in labor market outcomes for treated (blue line) and control (red line) groups. These estimates are the predicted values from separate regressions by group:
$y_{i(H),s,t}=f_{s}+f_{o(H)}+\sum_{\begin{subarray}{c}y=2017\\ y\neq 2019\end{subarray}}^{2024}\beta^{y}\cdot 1(\text{year}(t)=y)+X_{i(H),t}\delta+\epsilon_{i(H),s,t},$
where outcomes are residualized on state and husband-occupation fixed effects and controls for wife’s age, number of children, and both spouses’ education. The omitted year is 2019.
Figure A.3: Dynamic DiD Estimates of the Impact of Husbands’ WFH Shocks on Wives’ Labor Market Outcomes (cont’d)
(a) Usual Weekly Hours Worked
(b) Part-Time Status
(c) Log Wage (March Sample)
Notes: Each panel plots coefficients from a dynamic difference-in-differences specification comparing women whose husbands were in occupations with large WFH shocks to those in occupations with smaller changes. The omitted year is 2019. All regressions include state-by-time and husband-occupation fixed effects, and control for wife’s age, number of children, and education levels of both spouses. Confidence intervals are shown at the 95% level and are based on standard errors clustered at the husband-occupation level.
Figure A.4: Assortative Matching by Spousal WFH Exposure, Pre- and Post-COVID
(a) Pre-COVID (b) Post-COVID
Refer to caption Refer to caption
Notes: Each panel shows the distribution of wives across treatment quartiles conditional on their husband’s occupation being in a given WFH shock quartile, separately for the pre-COVID (panel a) and post-COVID (panel b) periods, using pooled CPS data from 2017 to 2024 with children of child-rearing age. Quartiles are based on the occupation-level change in WFH probability described in Subsection 2.3. The figure is designed to assess whether assortative matching by WFH exposure changed differentially following the pandemic.
Figure A.5: Changes in Fertility Outcomes by Husbands’ WFH Exposure
(a) Probability of Having a Child (b) Number of Children
Refer to caption Refer to caption
Notes: Each panel plots the change in fertility-related outcomes following the onset of COVID-19 across quartiles of the husband’s WFH exposure, defined by the occupation-level change in WFH probability as described in Subsection 2.3. Panel (a) displays the probability of having a child in the past 12 months; Panel (b) reports the average number of children in the household. The figure uses pooled CPS data of dual households from 2017 to 2024 and is intended to assess whether fertility patterns shifted differentially across exposure groups in the post-pandemic period.
Figure A.6: Kernel Density Plot Showing Common Support for Reweighting
Notes: This figure uses the 2017–2024 CPS sample of dual-parent households with children of child-rearing age. It plots the estimated probability that an observation is from the pre-COVID-19 period, obtained from a logit model on a set of covariates including spouse occupation, wife’s age and education, state, and race. The predicted values are derived from separate regressions by group:
$\Pr(\text{Pre-COVID})=f_{s}+f_{o(H)}+f_{e}+f_{a}+\gamma\text{Non-White}+\varepsilon,$
where $f_{s}$, $f_{o(H)}$, $f_{e}$, and $f_{a}$ represent fixed effects for state, husband’s occupation, wife’s education, and age bin, respectively. The densities indicate good overlap in covariate distributions before and after COVID, supporting the use of reweighting strategies such as DiNardo-Fortin-Lemieux (DiNardo et al. 1995) and inverse-propensity score weighting.
Figure A.7: Dynamic DiD of Share of the Day Spent in Labor Activities, by Year
(a) Working
Refer to caption Refer to caption
Wives Husbands
(b) Mainwork From Home
Refer to caption Refer to caption
Wives Husbands
(c) Mainwork Outside Home
Refer to caption Refer to caption
Wives Husbands
(d) Commuting
Refer to caption Refer to caption
Wives Husbands
Notes: Each panel from the first column plots coefficients from a dynamic difference-in-differences specification comparing women whose husbands were in occupations with large WFH shocks to those in occupations with smaller changes. Those in the second column compare husbands directly. The omitted year is 2019. All regressions include state-by-time and husband-occupation fixed effects, and control for wife’s age, number of children, and education levels of both spouses. Confidence intervals are shown at the 95% level and based on standard errors clustered at the husband’s occupation level.
Figure A.8: Dynamic DiD of Share of the Day Spent in Household Production, by Year
(a) Housework
Refer to caption Refer to caption
Wives Husbands
(b) Primary Childcare
Refer to caption Refer to caption
Wives Husbands
(c) Secondary Childcare
Refer to caption Refer to caption
Wives Husbands
(d) Ratio Secondary Childcare on Total Childcare
Refer to caption Refer to caption
Wives Husbands
Notes: Each panel from the first column plots coefficients from a dynamic difference-in-differences specification comparing women whose husbands were in occupations with large WFH shocks to those in occupations with smaller changes. Those in the second column compare husbands directly. The omitted year is 2019. All regressions include state-by-time and husband-occupation fixed effects, and control for wife’s age, number of children, and education levels of both spouses. Confidence intervals are shown at the 95% level and based on standard errors clustered at the husband’s occupation level.
Appendix B Tables
Table B.1: Top 15 CPS Occupations by Treatment Status
Treated Occupations Control Occupations
Managers, nec (including Postmasters) (10.3%) First-Line Supervisors of Sales Workers (6.5%)
Software Developers, Applications and Systems Software (5.2%) Driver/Sales Workers and Truck Drivers (6.4%)
Computer Scientists and Systems Analysts/Network systems Analysts/Web Developers (3.8%) Construction Laborers (4.4%)
Chief executives and legislators/public administration (3.1%) Carpenters (3.5%)
Police Officers and Detectives (3.0%) Elementary and Middle School Teachers (2.6%)
Constructions Managers (2.9%) Electricians (2.6%)
Lawyers, and judges, magistrates, and other judicial workers (2.8%) Physicians and Surgeons (2.5%)
General and Operations Managers (2.4%) Grounds Maintenance Workers (2.4%)
Sales Representatives, Wholesale and Manufacturing (2.4%) Retail Salespersons (2.4%)
Managers in Marketing, Advertising, and Public Relations (2.1%) Chefs and Cooks (2.3%)
Financial Managers (2.1%) First-Line Supervisors of Construction Trades and Extraction Workers (2.2%)
Farmers, Ranchers, and Other Agricultural Managers (1.9%) Laborers and Freight, Stock, and Material Movers, Hand (2.2%)
Accountants and Auditors (1.8%) Automotive Service Technicians and Mechanics (2.0%)
Computer and Information Systems Managers (1.8%) Janitors and Building Cleaners (1.9%)
Engineers, nec (1.8%) Other production workers including semiconductor processors and cooling and freezing equipment operators (1.7%)
  • Notes: This table lists the 15 most common occupations in the treated and control groups, respectively, based on the classification variable occ2010 from the CPS. The treatment dummy is assigned at the occupation level, as described in Subsection 2.3. An occupation is classified as treated if the change in the share of remote/hybrid job postings before and after COVID-19, as measured using data from Hansen et al. (2023), is above the median across all occupations. Frequencies are expressed as percentages of the total number of observations within each group. Occupation labels correspond to the value labels associated with occ2010.
Table B.2: Descriptive Statistics of Main Labor Market Outcomes, by Treatment Group and Pre/Post-COVID Period
Pre-COVID Post-COVID
Mean SD N Mean SD N
Panel A: Control Sample
Spouse Employed (Last Week) 1.00 0.00 89,805 1.00 0.00 34,832
Spouse LFP (Last Week) 1.00 0.00 89,805 1.00 0.00 34,832
Spouse Usual Work Hours 43.37 9.79 89,805 42.85 9.70 34,832
Spouse Part-time status 0.11 0.32 89,805 0.12 0.32 34,832
Spouse Employed (Last Year) 0.94 0.24 89,805 0.93 0.25 34,832
Spouse LFP (Last Year) 1.00 0.00 89,805 1.00 0.00 34,832
Spouse Part-time (Last Year) 0.04 0.19 89,805 0.05 0.21 34,832
Spouse Months Worked (Last Year) 11.68 1.33 89,805 11.72 1.22 34,832
Spouse Wage 41801.27 52925.01 89,805 43183.42 45112.61 34,832
Wife Employed (Last Week) 0.63 0.48 89,805 0.65 0.48 34,832
Wife LFP (Last Week) 0.64 0.48 89,805 0.67 0.47 34,832
Wife Usual Work Hours 36.74 10.50 89,805 36.87 10.06 34,832
Wife Part-time status 0.31 0.46 89,805 0.29 0.45 34,832
Wife Employed (Last Year) 0.63 0.48 89,805 0.66 0.47 34,832
Wife LFP (Last Year) 0.67 0.47 89,805 0.70 0.46 34,832
Wife Part-time (Last Year) 0.25 0.43 89,805 0.24 0.43 34,832
Wife Months Worked (Last Year) 10.85 2.62 89,805 11.00 2.40 34,832
Wife Wage 29377.36 36140.80 89,805 31438.11 30695.26 34,832
HH Wage 71203.43 66266.65 89,805 77014.98 60228.95 34,832
Panel B: Treated Sample
Spouse Employed (Last Week) 1.00 0.00 88,317 1.00 0.00 39,357
Spouse LFP (Last Week) 1.00 0.00 88,317 1.00 0.00 39,357
Spouse Usual Work Hours 44.38 9.11 88,317 43.44 8.75 39,357
Spouse Part-time status 0.08 0.27 88,317 0.09 0.29 39,357
Spouse Employed (Last Year) 0.95 0.21 88,317 0.96 0.19 39,357
Spouse LFP (Last Year) 1.00 0.00 88,317 1.00 0.00 39,357
Spouse Part-time (Last Year) 0.02 0.12 88,317 0.03 0.16 39,357
Spouse Months Worked (Last Year) 11.85 0.90 88,317 11.81 1.03 39,357
Spouse Wage 69080.29 63145.13 88,317 68379.72 62091.98 39,357
Wife Employed (Last Week) 0.69 0.46 88,317 0.73 0.44 39,357
Wife LFP (Last Week) 0.70 0.46 88,317 0.75 0.43 39,357
Wife Usual Work Hours 36.77 11.16 88,317 37.15 10.43 39,357
Wife Part-time status 0.30 0.46 88,317 0.29 0.45 39,357
Wife Employed (Last Year) 0.68 0.47 88,317 0.75 0.43 39,357
Wife LFP (Last Year) 0.72 0.45 88,317 0.78 0.41 39,357
Wife Part-time (Last Year) 0.24 0.43 88,317 0.20 0.40 39,357
Wife Months Worked (Last Year) 10.92 2.51 88,317 11.21 2.14 39,357
Wife Wage 40393.45 43102.21 88,317 46340.94 50705.32 39,357
HH Wage 105825.32 81284.33 88,317 113733.18 92238.76 39,357
  • Notes: This table reports descriptive statistics for key labor market outcomes across treated and control households, before and after the COVID-19 pandemic. The sample corresponds to the pooled 2017–2024 CPS basic monthly files, restricted to dual-parent households with children of child-rearing age. All estimates are computed using the CPS sampling weights.
Table B.3: Descriptive Statistics of Main Control Variables, by Treatment Group and Pre/Post-COVID Period
Pre-COVID Post-COVID
Mean SD N Mean SD N
Panel A: Control Sample
Spouse - Below HS Level 0.14 0.35 89,805 0.12 0.33 34,832
Spouse - HS Level 0.34 0.47 89,805 0.34 0.47 34,832
Spouse - Above HS Level 0.52 0.50 89,805 0.54 0.50 34,832
Spouse Age 36.61 7.12 89,805 37.29 6.94 34,832
Wife - Below HS Level 0.12 0.32 89,805 0.10 0.30 34,832
Wife - HS Level 0.25 0.43 89,805 0.25 0.43 34,832
Wife - Above HS Level 0.63 0.48 89,805 0.65 0.48 34,832
Wife Age 34.16 6.18 89,805 34.81 6.03 34,832
Number of own children in household 2.14 0.91 89,805 2.14 0.92 34,832
Age of Youngest Child 3.48 2.53 89,805 3.57 2.55 34,832
Age of Oldest Child 7.97 5.23 89,805 8.11 5.40 34,832
Youngest Child 0-4 0.65 0.48 89,805 0.63 0.48 34,832
Youngest Child 5-8 0.35 0.48 89,805 0.37 0.48 34,832
Panel B: Treated Sample
Spouse - Below HS Level 0.02 0.13 88,317 0.02 0.13 39,357
Spouse - HS Level 0.11 0.31 88,317 0.12 0.32 39,357
Spouse - Above HS Level 0.87 0.33 88,317 0.87 0.34 39,357
Spouse Age 38.03 6.30 88,317 38.55 6.17 39,357
Wife - Below HS Level 0.02 0.13 88,317 0.01 0.12 39,357
Wife - HS Level 0.10 0.30 88,317 0.10 0.30 39,357
Wife - Above HS Level 0.89 0.32 88,317 0.89 0.32 39,357
Wife Age 35.85 5.57 88,317 36.40 5.42 39,357
Number of own children in household 2.02 0.86 88,317 2.00 0.86 39,357
Age of Youngest Child 3.43 2.55 88,317 3.48 2.58 39,357
Age of Oldest Child 7.03 4.76 88,317 7.00 4.81 39,357
Youngest Child 0-4 0.65 0.48 88,317 0.64 0.48 39,357
Youngest Child 5-8 0.35 0.48 88,317 0.36 0.48 39,357
  • Notes: This table presents summary statistics for key household- and individual-level control variables across treated and control groups, before and after the COVID-19 pandemic. The statistics are based on the pooled 2017–2024 CPS basic monthly sample, restricted to dual-parent households with children of child-rearing age. All estimates are computed using the CPS sampling weights.
Table B.4: Heterogeneous Effects by Quartile of Husbands’ WFH Exposure
Monthly CPS sample March CPS sample
Employed LFP Usual work Part-time Employed LFP Part-time Months Log wages
hours status Status Worked
$Exposed_{o(H)}(Quartile2)\times Post_{t}$ -0.005 -0.007 0.138 -0.012 0.031* 0.031* 0.031 -0.264* -0.028
(0.014) (0.014) (0.319) (0.013) (0.017) (0.016) (0.026) (0.154) (0.059)
$Exposed_{o(H)}(Quartile3)\times Post_{t}$ 0.017 0.021 0.040 0.006 0.056*** 0.043** -0.007 -0.137 0.022
(0.015) (0.014) (0.252) (0.010) (0.020) (0.020) (0.022) (0.122) (0.050)
$Exposed_{o(H)}(Quartile4)\times Post_{t}$ 0.028* 0.030** 0.560** -0.008 0.041** 0.039** -0.030 0.174 0.125***
(0.015) (0.014) (0.263) (0.011) (0.020) (0.019) (0.021) (0.132) (0.048)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.665 0.708 0.247 10.897 9.992
Obs 252,309 252,309 162,882 161,095 38,637 38,637 25,920 25,920 25,920
$R^{2}$ 0.12 0.12 0.06 0.06 0.12 0.13 0.07 0.07 0.17
  • Notes: This table reports estimates from Equation (1), where treatment intensity is captured by quartiles of the husband’s occupation-level change in WFH probability, as defined in Subsection 2.3. The sample is the 2017–2024 CPS sample of dual-parent households with children of child-rearing age. The baseline (omitted) category is the bottom quartile of WFH exposure. The dependent variables include both weekly outcomes (employment, labor force participation, hours worked, part-time status) from the monthly CPS and annual outcomes (employment, labor force participation, months worked, part-time status, and log wages) from the March CPS (ASEC). All regressions control for the wife’s age, education, and number of children, as well as the husband’s education. We include state-by-time and husband-occupation fixed effects. Standard errors are clustered at the occupation level.
Table B.5: Effect of Husbands’ Industry-Level WFH Exposure on Wives’ Labor Market Outcomes
Monthly CPS sample March CPS sample
Employed LFP Usual work Part-time Employed LFP Part-time Months Log wages
hours status Status Worked
$Exposed_{n(H)}\times Post_{t}$ 0.019** 0.019* 0.264 -0.008 0.034** 0.026** 0.021 0.166* 0.036
(0.010) (0.010) (0.200) (0.008) (0.013) (0.013) (0.017) (0.093) (0.048)
Mean pre-treatment 0.671 0.686 36.661 0.313 0.669 0.713 0.246 10.910 9.995
Obs 257,733 257,733 166,132 164,375 36,483 36,483 24,596 24,596 24,596
$R^{2}$ 0.11 0.11 0.05 0.05 0.11 0.12 0.06 0.06 0.17
  • Notes: This table reports estimates from Equation (1), using the 2017–2024 CPS sample of dual-parent households with children of child-rearing age. Unlike in the main specification, treatment is defined using the husband’s industry (classified at the 2012 Census industry level), based on the change in remote job postings before and after the COVID-19 pandemic. Treated households are those in which the husband’s industry experienced a WFH probability increase above the median. We replace husband-occupation fixed effects with husband-industry fixed effects. All regressions include controls for the wife’s age, education, and number of children, as well as the husband’s education. State-by-time fixed effects are included throughout. Standard errors are clustered at the industry level.
Table B.6: Effects of Husbands’ Occupation- and Industry-Level WFH Shocks on Wives’ Labor Market Outcomes
Monthly CPS sample March CPS sample
Employed LFP Usual work Part-time Employed LFP Part-time Months Log wages
hours status Status Worked
$Exposed_{o(H)}\times Post_{t}$ 0.022 0.027* -0.167 0.015 0.057** 0.044** -0.018 0.088 0.095
(0.015) (0.014) (0.280) (0.012) (0.022) (0.020) (0.021) (0.154) (0.072)
$Exposed_{n(H)}\times Post_{t}$ 0.012 0.012 -0.147 -0.002 0.039** 0.029* 0.054** 0.183 0.046
(0.014) (0.014) (0.270) (0.010) (0.017) (0.016) (0.025) (0.146) (0.074)
$Exposed_{o(H)}\times Exposed_{n(H)}\times Post_{t}$ -0.004 -0.007 0.602 -0.013 -0.042 -0.030 -0.037 -0.045 -0.068
(0.017) (0.017) (0.385) (0.016) (0.027) (0.024) (0.031) (0.199) (0.091)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.669 0.713 0.246 10.910 9.995
Obs 252,309 252,309 162,882 161,095 36,473 36,473 24,577 24,577 24,577
$R^{2}$ 0.13 0.13 0.07 0.06 0.13 0.14 0.08 0.08 0.19
  • Notes: This table presents estimates from Equation (1), augmented to include both occupation- and industry-level treatment indicators and their interaction. Treatment is defined based on the change in the probability of WFH before and after the COVID-19 pandemic, using job postings data at the husband’s occupation ($o(H)$, Census 2010) and industry ($n(H)$, Census 2012) levels. The sample is the 2017–2024 CPS sample of dual-parent households with children of child-rearing age. All regressions include fixed effects for state-by-time, husband’s occupation, and husband’s industry. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.7: Effect of Husbands’ WFH Occupation Shocks on Wives’ Probability of Working Remotely
Cross-section Panel Panel with interaction
Spouse Restriction: All Employee Self-employed All Employee Self-employed All Employee Self-employed
$Exposed_{o(H)}\times Post_{t}$ 0.110*** 0.113*** 0.068*** 0.096*** 0.101*** 0.059*** 0.031*** 0.033*** 0.019
(0.014) (0.015) (0.025) (0.013) (0.014) (0.022) (0.008) (0.009) (0.023)
$Exposed_{o(i)}\times Post_{t}$ 0.158*** 0.160*** 0.141***
(0.011) (0.013) (0.024)
$Exposed_{o(H)}\times Exposed_{o(i)}\times Post_{t}$ 0.064*** 0.069*** 0.026
(0.015) (0.017) (0.033)
Mean pre-treatment 0.135 0.130 0.170 0.135 0.130 0.170 0.135 0.130 0.170
Obs 86,328 76,095 10,091 84,798 74,686 9,922 84,791 74,681 9,922
$R^{2}$ 0.16 0.17 0.29 0.65 0.66 0.67 0.67 0.67 0.68
  • Notes: We regress women’s self-reported teleworking against their spouses’ treatment status using our cross-section and panel specifications. Data comes from our household sample of the 2023-2024 monthly CPS. Respondents are asked if they work from home in their current occupation (post-Covid observation) and are asked whether they worked from home on February 2020 (pre-Covid observation). We use this two data points to build a pseudo-panel. It is implicitly assumed that they are working on same occupation post-Covid as they did pre-Covid. In the first three columns, we include fixed effects for month-year-state and spouse occupation, hence we label them as the cross-section regression. In the last three columns, we replace occupation with household fixed effects to exploit the panel dimension of the recalled data, hence we label them as the panel regressions. For each specification, we look at results in our regular sample of households, households where the male spouse is on an employee arrangement, and where they are self-employed.
Table B.8: Effects by Age of Children Distribution
Monthly CPS Sample March CPS Sample
Employed LFP Usual work Part-time Employed LFP Part-time Months Log wages
hours status Status Worked
A. All Households
$Exposed_{o(H)}\times Post_{t}$=1 0.030** 0.035*** 0.203 0.007 0.037** 0.037** -0.032 0.143 0.072*
(0.012) (0.011) (0.267) (0.011) (0.017) (0.016) (0.021) (0.119) (0.041)
$Exposed_{o(H)}\times Post_{t}\times 1(AgeYC\in[5,8])$ -0.016 -0.020 0.139 -0.012 -0.007 -0.033 0.000 -0.091 0.046
(0.018) (0.018) (0.405) (0.016) (0.030) (0.027) (0.039) (0.180) (0.072)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.664 0.708 0.247 10.896 9.993
$\beta_{1}+\beta_{2}=0$ (p-val) 0.436 0.378 0.229 0.655 0.257 0.856 0.296 0.725 0.068
Obs 252,303 252,303 162,875 161,086 38,593 38,593 25,859 25,859 25,859
$R^{2}$ 0.14 0.14 0.08 0.08 0.14 0.14 0.09 0.09 0.19
B. Households with Two or More Children
$Exposed_{o(H)}\times Post_{t}$=1 0.038*** 0.042*** 0.401 0.011 0.026 0.030 -0.037 0.206 0.091
(0.013) (0.013) (0.336) (0.015) (0.021) (0.021) (0.026) (0.149) (0.057)
$Exposed_{o(H)}\times Post_{t}\times 1(AgeYC\in[5,8])$ -0.028 -0.032 -0.285 -0.010 0.005 -0.029 0.005 -0.119 0.040
(0.022) (0.021) (0.515) (0.021) (0.036) (0.031) (0.048) (0.232) (0.094)
Mean pre-treatment 0.651 0.665 36.213 0.330 0.640 0.686 0.269 10.876 9.958
$\beta_{1}+\beta_{2}=0$ (p-val) 0.629 0.613 0.746 0.973 0.327 0.979 0.389 0.639 0.104
Obs 183,054 183,054 114,370 113,921 28,190 28,190 18,187 18,187 18,187
$R^{2}$ 0.15 0.15 0.11 0.10 0.15 0.15 0.11 0.11 0.21
C. Households with One Child
$Exposed_{o(H)}\times Post_{t}$=1 0.011 0.015 -0.383 0.014 0.065** 0.054* -0.009 -0.069 -0.017
(0.021) (0.021) (0.418) (0.016) (0.032) (0.030) (0.032) (0.168) (0.060)
$Exposed_{o(H)}\times Post_{t}\times 1(AgeYC\in[5,8])$ 0.021 0.021 1.568** -0.045 -0.021 -0.028 -0.017 0.292 0.189
(0.032) (0.030) (0.701) (0.027) (0.073) (0.065) (0.073) (0.309) (0.120)
Mean pre-treatment 0.728 0.744 37.618 0.272 0.732 0.768 0.198 10.944 10.076
$\beta_{1}+\beta_{2}=0$ (p-val) 0.219 0.125 0.041 0.199 0.514 0.665 0.696 0.377 0.109
Obs 69,059 69,059 48,126 46,807 10,251 10,251 7,485 7,485 7,485
$R^{2}$ 0.19 0.19 0.18 0.18 0.20 0.21 0.17 0.17 0.28
Notes: This table reports estimates of baseline outcomes, decomposing the effects by the age of the child. Panel A uses the sample of households with up to 4 children where the youngest child is between 0 and 8 years old. The coefficient on $Exposed_{o(H)}\times Post_{t}$ captures the effect for women in households where the youngest child is between 5 and 8 years old, while the coefficient on the interaction term $Exposed_{o(H)}\times Post_{t}\times 1(AgeYC\in[5,8])$ captures the differential effect for women whose youngest child is between 5 and 8 years old relative to those with a youngest child aged 0 to 4. Additionally, we interact all our baseline fixed effects by age of youngest child. Panels B and C repeat the same analysis, but panel B is estimated on a sample of at households with two or more children (i.e. households with at least 2 children but no more than 4) and panel C uses a sample of households with one child only. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level. We include a test for whether the sum of the coefficients is equal to 0.
Table B.9: Effect of Wives’ WFH Occupation Shocks on Their Own Labor Market Outcomes
Monthly CPS sample March CPS sample
Usual work Part-time Part-time Months Log wages
hours status Status Worked
$Exposed_{o(i)}\times Post_{t}$ -0.086 0.004 -0.025 0.077 0.051
(0.261) (0.009) (0.017) (0.076) (0.031)
Mean pre-treatment 36.839 0.306 0.227 11.225 10.100
Obs 175,724 173,862 25,006 25,006 25,006
$R^{2}$ 0.14 0.12 0.16 0.08 0.36
  • Notes: This table adapts our baseline cross-sectional specification to estimate the effect of wives’ own WFH shocks on their labor market outcomes, using the 2017–2024 CPS sample of dual-parent households with children of child-rearing age. Columns 1 and 2 report estimates for prior-week outcomes, using contemporaneous occupation-based WFH shocks as treatment. Columns 3 through 5 report estimates for previous-year outcomes (March CPS sample only), using occupation shocks measured for the previous calendar year.
Table B.10: Effect of Husband’s WFH Shock on Marriage, Divorce, and Fertility Outcomes
Married Divorced Had a child
in the last 12mm in the last 12mm in the last 12mm
$Exposed_{o(H)}\times Post_{t}$ 0.001 0.000 -0.007**
(0.002) (0.000) (0.003)
Mean pre-treatment 0.031 0.001 0.173
Obs 516,557 516,393 516,557
$R^{2}$ 0.04 0.00 0.32
  • Notes: This table reports estimates using data from the pooled ACS sample from 2017 to 2023. The outcomes—marriage, divorce, and fertility—are defined over the 12 months preceding the interview date. Occupation is measured only at the time of the interview, so we implicitly assume that individuals remained in the same occupation throughout the prior 12-month period.
Table B.11: Effect of Husband’s WFH Shock on Fertility Outcomes
Pr(At least one child) Number of Children
$Exposed_{o(H)}\times Post_{t}$ 0.002 -0.008
(0.005) (0.016)
Mean pre-treatment 0.738 1.551
Obs 593,624 593,624
$R^{2}$ 0.07 0.10
  • Notes: This table reports estimates from Equation (1) using the pooled 2017–2024 CPS data. The outcomes include the probability of having had a child in the past year and the number of children in the household. Treatment is defined based on the husband’s occupation-level WFH exposure, as detailed in Subsection 2.3. All regressions include controls for the wife’s age, number of children, and education, as well as the husband’s education. State-by-time and husband-occupation fixed effects are included. Standard errors are clustered at the occupation level.
Table B.12: Effect of Husband’s WFH Shock on Wives’ Labor Market Outcomes Using Reweighting Techniques
Monthly CPS sample March CPS sample
Employed LFP Usual work Part-time Employed LFP Part-time Months Log wages
hours status Status Worked
A. Using DFL Weights
$Exposed_{o(H)}\times Post_{t}$ 0.024** 0.028*** 0.199 0.004 0.036** 0.027* -0.032** 0.130 0.088**
(0.011) (0.010) (0.197) (0.008) (0.015) (0.015) (0.016) (0.100) (0.036)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.671 0.715 0.249 10.932 10.007
Obs 252,309 252,309 162,882 161,095 36,118 36,118 24,462 24,462 24,462
$R^{2}$ 0.12 0.12 0.06 0.06 0.12 0.13 0.07 0.07 0.18
B. Using IPW
$Exposed_{o(H)}\times Post_{t}$ 0.023** 0.027** 0.236 0.003 0.035** 0.026* -0.032** 0.137 0.096***
(0.011) (0.010) (0.196) (0.008) (0.015) (0.014) (0.016) (0.097) (0.037)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.671 0.715 0.249 10.932 10.007
Obs 252,309 252,309 162,882 161,095 36,118 36,118 24,462 24,462 24,462
$R^{2}$ 0.12 0.12 0.06 0.06 0.12 0.12 0.07 0.07 0.17
  • Notes: This table reports estimates from Equation (1), using the 2017–2024 CPS sample of dual-parent households with children of child-rearing age. Panel A applies reweighting using DiNardo-Fortin-Lemieux (DFL) weights (DiNardo et al., 1995), while Panel B uses inverse-propensity score weights (IPW). Both methods aim to correct for potential changes in group composition before and after the COVID-19 period. Details on the construction of these weights are provided in Section 5.
Table B.13: Test: Preliminary Results Using Alternative Definitions of WFH Shock
Monthly CPS Sample March CPS Sample
Employed LFP Usual work Part-time Employed LFP Part-time Months Log wages
hours status Status Worked
A. Shock: Difference in average 2023-2024 vs 2019 WFH vacancy rates at Census 2010 4d level (Original)
$Exposed_{o(H)}\times Post_{t}$ 0.024** 0.027*** 0.217 0.004 0.044*** 0.035** -0.030* 0.137 0.075**
(0.011) (0.010) (0.197) (0.008) (0.015) (0.015) (0.016) (0.097) (0.038)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.671 0.715 0.248 10.932 10.007
Obs 252,309 252,309 162,882 161,095 36,158 36,158 24,485 24,485 24,485
$R^{2}$ 0.12 0.12 0.06 0.06 0.12 0.12 0.07 0.07 0.17
B. Shock: Difference in 3 month moving average 2023-2024 vs 2019 WFH vacancy rates at Census 2010 4d level
$Exposed_{o(H)}\times Post_{t}$ 0.024** 0.027*** 0.217 0.004 0.044*** 0.035** -0.030* 0.137 0.075**
(0.011) (0.010) (0.197) (0.008) (0.015) (0.015) (0.016) (0.097) (0.038)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.671 0.715 0.248 10.932 10.007
Obs 252,309 252,309 162,882 161,095 36,158 36,158 24,485 24,485 24,485
$R^{2}$ 0.12 0.12 0.06 0.06 0.12 0.12 0.07 0.07 0.17
C. Shock: Difference in July 2023 vs July 2019 WFH vacancy rates at Census 2010 4d level
$Exposed_{o(H)}\times Post_{t}$ 0.024** 0.026** 0.177 0.002 0.050*** 0.040*** -0.028* 0.123 0.066*
(0.011) (0.010) (0.197) (0.008) (0.015) (0.015) (0.016) (0.099) (0.039)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.671 0.715 0.248 10.932 10.007
Obs 252,309 252,309 162,882 161,095 36,158 36,158 24,485 24,485 24,485
$R^{2}$ 0.12 0.12 0.06 0.06 0.12 0.12 0.07 0.07 0.17
D. Shock: WFH teleworkability a-la Dingel and Neumann at Census 2010 4d level
$Exposed_{o(H)}\times Post_{t}$ 0.042*** 0.044*** 0.542*** -0.014* 0.049*** 0.042*** -0.018 0.214** 0.052
(0.010) (0.009) (0.190) (0.008) (0.014) (0.014) (0.016) (0.096) (0.038)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.671 0.715 0.248 10.932 10.007
Obs 252,309 252,309 162,882 161,095 36,158 36,158 24,485 24,485 24,485
$R^{2}$ 0.12 0.12 0.06 0.06 0.12 0.12 0.07 0.07 0.17
E. Shock: Difference in average 2023-2024 vs 2019 WFH vacancy rates at SOC 3d level
$Exposed_{o(H)}\times Post_{t}$ 0.022** 0.025** 0.271 0.001 0.041*** 0.032** -0.026* 0.152 0.086**
(0.011) (0.011) (0.195) (0.008) (0.015) (0.015) (0.016) (0.097) (0.038)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.671 0.715 0.249 10.932 10.006
Obs 252,311 252,311 162,884 161,097 36,167 36,167 24,501 24,501 24,501
$R^{2}$ 0.11 0.11 0.05 0.05 0.10 0.11 0.05 0.05 0.16
F. Shock: Difference in 3 month moving average 2023-2024 vs 2019 vacancy rates at SOC 3d level
$Exposed_{o(H)}\times Post_{t}$ 0.022** 0.026** 0.222 0.002 0.043*** 0.034** -0.027* 0.147 0.086**
(0.011) (0.011) (0.196) (0.008) (0.015) (0.015) (0.016) (0.097) (0.038)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.671 0.715 0.249 10.932 10.006
Obs 252,311 252,311 162,884 161,097 36,167 36,167 24,501 24,501 24,501
$R^{2}$ 0.11 0.11 0.05 0.05 0.10 0.11 0.05 0.05 0.16
G. Shock: Difference in July 2023 vs July 2019 WFH vacancy rates at SOC 3d level
$Exposed_{o(H)}\times Post_{t}$ 0.020* 0.023** 0.195 -0.001 0.040*** 0.028* -0.021 0.139 0.071*
(0.011) (0.011) (0.198) (0.008) (0.015) (0.015) (0.016) (0.099) (0.039)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.671 0.715 0.249 10.932 10.006
Obs 252,311 252,311 162,884 161,097 36,167 36,167 24,501 24,501 24,501
$R^{2}$ 0.11 0.11 0.05 0.05 0.10 0.11 0.05 0.05 0.16
H. Shock: WFH teleworkability a-la Dingel and Neumann at SOC 3d level
$Exposed_{o(H)}\times Post_{t}$ 0.024** 0.027** 0.236 -0.005 0.039** 0.032** -0.028* 0.153 0.067*
(0.011) (0.011) (0.203) (0.008) (0.015) (0.015) (0.017) (0.102) (0.040)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.671 0.715 0.249 10.932 10.006
Obs 252,311 252,311 162,884 161,097 36,167 36,167 24,501 24,501 24,501
$R^{2}$ 0.11 0.11 0.05 0.05 0.10 0.11 0.05 0.05 0.16
  • Notes:: In panel A, an occupation is treated if the difference in average 2023 and 2024 vs 2019 remote job vacancy shares is above the median of all other Census 2010 4d level occupations. Panel B uses the 3 month moving average version of the WFH monthly variables. In panel C, an occupation is shocked if the difference in July 2023 vs July 2019 WFH or remote job vacancy shares is above the median of all other Census 2010 4d level occupations. In panel D, an occupation is shocked if its Dingel and Neumann teleworkability index is above the median of all other Census 2010 4d level occupations. Panels E, F, G, H are analogous to A, B, C, D respectively, with the change that occupations are compared at the SOC 2018 3d level. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.14: Robustness of CPS Results to Alternative Fixed Effects Specifications
Monthly CPS sample March CPS sample
Employed LFP Usual work Part-time Employed LFP Part-time Months Log wages
hours status Status Worked
A. FE: Month-year + Spouse Occupation (Group of 16) + State
$Exposed_{o(H)}\times Post_{t}$ 0.028*** 0.031*** 0.148 0.001 0.031** 0.022 -0.033** 0.134 0.128***
(0.010) (0.010) (0.208) (0.008) (0.014) (0.014) (0.015) (0.094) (0.034)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.671 0.715 0.249 10.932 10.007
Obs 252,311 252,311 162,884 161,097 36,128 36,128 24,477 24,477 24,477
$R^{2}$ 0.10 0.10 0.03 0.03 0.09 0.10 0.04 0.04 0.14
B. FE: Month-year-state + Spouse Occupation (Group of 16)
$Exposed_{o(H)}\times Post_{t}$ 0.027*** 0.029*** 0.105 0.002 0.030** 0.022 -0.032** 0.141 0.125***
(0.010) (0.010) (0.201) (0.008) (0.014) (0.014) (0.015) (0.095) (0.033)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.671 0.715 0.249 10.932 10.007
Obs 252,311 252,311 162,884 161,097 36,128 36,128 24,477 24,477 24,477
$R^{2}$ 0.11 0.11 0.04 0.04 0.10 0.10 0.05 0.05 0.15
C. FE: Month-year-state + Spouse Occupation (Group of 16)-by-state
$Exposed_{o(H)}\times Post_{t}$ 0.029*** 0.031*** 0.080 0.003 0.031** 0.022 -0.035** 0.118 0.113***
(0.010) (0.010) (0.203) (0.008) (0.014) (0.014) (0.016) (0.095) (0.034)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.671 0.715 0.249 10.932 10.007
Obs 252,307 252,307 162,883 161,095 36,119 36,119 24,456 24,456 24,456
$R^{2}$ 0.12 0.12 0.06 0.06 0.12 0.12 0.08 0.08 0.18
D. FE: Month-year + Spouse Occupation (4d) + State
$Exposed_{o(H)}\times Post_{t}$ 0.026** 0.029*** 0.267 0.002 0.034** 0.027* -0.031** 0.101 0.079**
(0.011) (0.011) (0.204) (0.008) (0.015) (0.014) (0.016) (0.094) (0.036)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.665 0.708 0.247 10.897 9.992
Obs 252,309 252,309 162,882 161,095 38,637 38,637 25,920 25,920 25,920
$R^{2}$ 0.11 0.11 0.04 0.04 0.12 0.12 0.06 0.06 0.17
E. FE: Month-year-state + Spouse Occupation (4d) (Baseline)
$Exposed_{o(H)}\times Post_{t}$ 0.024** 0.027*** 0.217 0.004 0.035** 0.027* -0.030* 0.120 0.078**
(0.011) (0.010) (0.197) (0.008) (0.015) (0.014) (0.016) (0.095) (0.036)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.665 0.708 0.247 10.897 9.992
Obs 252,309 252,309 162,882 161,095 38,637 38,637 25,920 25,920 25,920
$R^{2}$ 0.12 0.12 0.06 0.06 0.12 0.13 0.07 0.07 0.17
F. FE: Month-year-state + Spouse Occupation (4d)-by-state
$Exposed_{o(H)}\times Post_{t}$ 0.025** 0.028** 0.234 0.004 0.045*** 0.037** -0.034* 0.196* 0.079*
(0.012) (0.011) (0.225) (0.010) (0.016) (0.015) (0.018) (0.110) (0.043)
Mean pre-treatment 0.672 0.686 36.635 0.313 0.661 0.705 0.248 10.909 10.005
Obs 251,222 251,222 161,717 159,944 35,040 35,040 22,501 22,501 22,501
$R^{2}$ 0.24 0.24 0.22 0.20 0.26 0.26 0.26 0.25 0.33
  • Notes: This table presents estimates from alternative versions of Equation (1) using the 2017–2024 CPS sample of dual-parent households with children of child-rearing age, varying the fixed effects (FE) structure. Panel E reports the baseline specification, which includes state-by-time and husband-occupation fixed effects. Panel A includes separate fixed effects for month-year, state, and broad occupation groups (collapsed into 16 categories such as Business/Finance, Education Professionals, Healthcare, Sales, Construction, etc.). Panel B adds state-by-time fixed effects, and Panel C includes state-by-time and broad occupation group fixed effects. Panels D to F replicate Panels A to C, respectively, but use detailed 4-digit Census occupation fixed effects instead of the broader groupings. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.15: Robustness of CPS Results to Sequential Removal of Demographic Controls
Control Set 1 Control Set 2 Control Set 3 Control Set 4 Control Set 5
(No controls) (Baseline)
Outcome: Weekly Employment
$Exposed_{o(H)}\times Post_{t}$ 0.020* 0.019* 0.021* 0.024** 0.024**
(0.011) (0.011) (0.011) (0.011) (0.011)
Mean pre-treatment 0.672 0.672 0.672 0.672 0.672
Obs 252,309 252,309 252,309 252,309 252,309
$R^{2}$ 0.06 0.08 0.09 0.12 0.12
Outcome: Weekly LFP
$Exposed_{o(H)}\times Post_{t}$ 0.024** 0.022** 0.024** 0.027*** 0.027***
(0.010) (0.010) (0.011) (0.010) (0.010)
Mean pre-treatment 0.686 0.686 0.686 0.686 0.686
Obs 252,309 252,309 252,309 252,309 252,309
$R^{2}$ 0.06 0.08 0.09 0.12 0.12
Control: Number of children No Yes Yes Yes Yes
Control: Women’s age (5y sbin) No No Yes Yes Yes
Control: Women’s education level No No No Yes Yes
Control: Husband’s education level No No No No Yes
  • Notes: This table reports results from estimating Equation (1) on the 2017–2024 CPS sample of dual-parent households with children of child-rearing age. The baseline specification, shown in Column 5, includes controls for the wife’s age (in 5-year bins), number of children in the household, and both spouses’ education levels. Columns 1–4 sequentially remove these controls to assess the sensitivity of the results to the inclusion of covariates. All specifications include state-by-time and husband-occupation fixed effects.
Table B.16: Robustness of CPS Results to Adding Controls for Children’s Age
Control Set 1 Control Set 2 Control Set 3 Control Set 4 Control Set 5
Outcome: Weekly Employment
$Exposed_{o(H)}\times Post_{t}$ 0.024** 0.025** 0.025** 0.026** 0.025**
(0.011) (0.011) (0.011) (0.011) (0.011)
Mean pre-treatment 0.672 0.672 0.672 0.672 0.672
Obs 252,309 252,309 252,309 252,309 252,309
$R^{2}$ 0.12 0.12 0.12 0.13 0.12
Outcome: Weekly LFP
$Exposed_{o(H)}\times Post_{t}$ 0.027*** 0.028*** 0.029*** 0.029*** 0.029***
(0.010) (0.010) (0.011) (0.011) (0.011)
Mean pre-treatment 0.686 0.686 0.686 0.686 0.686
Obs 252,309 252,309 252,309 252,309 252,309
$R^{2}$ 0.12 0.12 0.12 0.13 0.12
Control: Baseline controls Yes Yes Yes Yes Yes
Control: Age youngest child No Yes No Yes Yes
Control: Age oldest child (raw) No No Yes Yes No
Control: Age oldest child (4y bins) No No No No Yes
  • Notes: This table reports estimates from Equation (1) using the 2017–2024 CPS sample of dual-parent households with children of child-rearing age. Column 1 reproduces the baseline specification, which includes controls for the wife’s age (in 5-year bins), number of children in the household, and both spouses’ education levels. In Columns 2 through 5, we sequentially add controls for the age of the youngest and oldest child, both as continuous variables and in four-year bins. All regressions include state-by-time and husband-occupation fixed effects.
Table B.17: Robustness to Alternative Definitions of the Post-COVID Period
Monthly CPS sample March CPS sample
Employed LFP Usual work Part-time Employed LFP Part-time Months Log wages
hours status Status Worked
A. Drop March 2020 - June 2023
$Exposed_{o(H)}\times Post_{t}$ 0.027** 0.031*** 0.217 0.004 0.036** 0.027* -0.024 0.156 0.078**
(0.011) (0.010) (0.201) (0.009) (0.016) (0.014) (0.016) (0.096) (0.038)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.658 0.702 0.251 10.855 9.970
Obs 242,010 242,010 155,875 154,220 31,429 31,429 20,968 20,968 20,968
$R^{2}$ 0.12 0.12 0.06 0.06 0.12 0.13 0.07 0.08 0.18
B. Drop March 2020 - March 2023 (Baseline)
$Exposed_{o(H)}\times Post_{t}$ 0.024** 0.027*** 0.217 0.004 0.036** 0.027* -0.024 0.156 0.078**
(0.011) (0.010) (0.197) (0.008) (0.016) (0.014) (0.016) (0.096) (0.038)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.658 0.702 0.251 10.855 9.970
Obs 252,309 252,309 162,882 161,095 31,429 31,429 20,968 20,968 20,968
$R^{2}$ 0.12 0.12 0.06 0.06 0.12 0.13 0.07 0.08 0.18
C. Drop March 2020 - December 2022
$Exposed_{o(H)}\times Post_{t}$ 0.026** 0.029*** 0.254 -0.001 0.038*** 0.033*** -0.026** 0.168** 0.088***
(0.010) (0.010) (0.195) (0.008) (0.013) (0.012) (0.012) (0.074) (0.033)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.658 0.702 0.251 10.855 9.970
Obs 265,074 265,074 171,602 169,789 37,800 37,800 25,415 25,415 25,415
$R^{2}$ 0.12 0.12 0.06 0.06 0.12 0.13 0.07 0.07 0.18
D. Drop March 2020 - September 2022
$Exposed_{o(H)}\times Post_{t}$ 0.027*** 0.030*** 0.223 -0.001 0.038*** 0.033*** -0.026** 0.168** 0.088***
(0.010) (0.010) (0.194) (0.008) (0.013) (0.012) (0.012) (0.074) (0.033)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.658 0.702 0.251 10.855 9.970
Obs 275,405 275,405 178,523 176,700 37,800 37,800 25,415 25,415 25,415
$R^{2}$ 0.12 0.12 0.06 0.06 0.12 0.13 0.07 0.07 0.18
E. Drop March 2020 - June 2022
$Exposed_{o(H)}\times Post_{t}$ 0.026*** 0.030*** 0.274 -0.007 0.038*** 0.033*** -0.026** 0.168** 0.088***
(0.010) (0.009) (0.198) (0.008) (0.013) (0.012) (0.012) (0.074) (0.033)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.658 0.702 0.251 10.855 9.970
Obs 285,943 285,943 185,410 183,305 37,800 37,800 25,415 25,415 25,415
$R^{2}$ 0.12 0.12 0.06 0.05 0.12 0.13 0.07 0.07 0.18
F. Drop March 2020 - March 2022
$Exposed_{o(H)}\times Post_{t}$ 0.027*** 0.030*** 0.291 -0.007 0.038*** 0.033*** -0.026** 0.168** 0.088***
(0.009) (0.009) (0.193) (0.008) (0.013) (0.012) (0.012) (0.074) (0.033)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.658 0.702 0.251 10.855 9.970
Obs 296,868 296,868 192,437 190,217 37,800 37,800 25,415 25,415 25,415
$R^{2}$ 0.12 0.12 0.06 0.05 0.12 0.13 0.07 0.07 0.18
G. Drop March 2020 - December 2021
$Exposed_{o(H)}\times Post_{t}$ 0.028*** 0.031*** 0.336* -0.010 0.033*** 0.027*** -0.032*** 0.192*** 0.090***
(0.009) (0.009) (0.183) (0.007) (0.011) (0.010) (0.010) (0.066) (0.029)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.658 0.702 0.251 10.855 9.970
Obs 310,380 310,380 201,248 198,979 44,484 44,484 29,953 29,953 29,953
$R^{2}$ 0.12 0.12 0.06 0.05 0.12 0.12 0.07 0.07 0.18
H. Drop March 2020 - September 2021
$Exposed_{o(H)}\times Post_{t}$ 0.030*** 0.032*** 0.372** -0.011 0.033*** 0.027*** -0.032*** 0.192*** 0.090***
(0.009) (0.008) (0.177) (0.007) (0.011) (0.010) (0.010) (0.066) (0.029)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.658 0.702 0.251 10.855 9.970
Obs 321,251 321,251 208,280 205,970 44,484 44,484 29,953 29,953 29,953
$R^{2}$ 0.12 0.12 0.06 0.05 0.12 0.12 0.07 0.07 0.18
  • Notes: This table reports estimates from Equation (1) using the 2017–2024 CPS sample of dual-parent households with children of child-rearing age. We examine the sensitivity of our baseline results to alternative definitions of the post-COVID period. Panel A drops all observations between March 2020 and June 2023. Panel B reports our baseline specification, which excludes data from March 2020 to March 2023. Panels C through G progressively shift the post-COVID window earlier by three-month increments, ending with a definition where post-COVID begins in September 2021. All regressions include state-by-time and husband-occupation fixed effects, and control for the wife’s age, number of children, and both spouses’ education levels.
Table B.18: Robustness to Excluding Households with Self-Employed Husbands
Monthly CPS sample March CPS sample
Employed LFP Usual work Part-time Employed LFP Part-time Months Log wages
hours status Status Worked
$Exposed_{o(H)}\times Post_{t}$ 0.033*** 0.036*** 0.245 0.004 0.040*** 0.032** -0.031* 0.192* 0.092**
(0.011) (0.011) (0.213) (0.009) (0.014) (0.014) (0.018) (0.104) (0.039)
Mean pre-treatment 0.670 0.685 36.688 0.308 0.668 0.706 0.244 10.884 9.988
Obs 223,112 223,112 144,361 142,372 33,835 33,835 22,833 22,833 22,833
$R^{2}$ 0.12 0.12 0.06 0.06 0.13 0.13 0.07 0.07 0.18
  • Notes: This table reports results from Equation (1) using the 2017–2024 CPS sample of dual-parent households with children of child-rearing age. We restrict the sample to households where the husband is classified as an employee (i.e., excluding households in which the husband reports being self-employed). The treatment is defined based on the husband’s occupation-level WFH shock, as in the baseline specification. All regressions include state-by-time and husband-occupation fixed effects, as well as controls for the wife’s age, number of children, and both spouses’ education levels.
Table B.19: Robustness of Wage Results to Trimming High-Income Households (March CPS)
Employed LFP Part-time Months Log wages
Status Worked
A. No income trim (Baseline)
$Exposed_{o(H)}\times Post_{t}$ 0.035** 0.027* -0.030* 0.120 0.078**
(0.015) (0.014) (0.016) (0.095) (0.036)
Mean pre-treatment 0.665 0.708 0.247 10.897 9.992
Obs 38,637 38,637 25,920 25,920 25,920
$R^{2}$ 0.12 0.13 0.07 0.07 0.17
B. Trim top 1% income earners
$Exposed_{o(H)}\times Post_{t}$ 0.034** 0.026* -0.030* 0.113 0.073**
(0.015) (0.014) (0.016) (0.096) (0.035)
Mean pre-treatment 0.663 0.706 0.249 10.889 9.969
Obs 38,403 38,403 25,686 25,686 25,686
$R^{2}$ 0.12 0.13 0.07 0.07 0.17
C. Trim top 5% income earners
$Exposed_{o(H)}\times Post_{t}$ 0.032** 0.024* -0.027* 0.111 0.057
(0.015) (0.014) (0.016) (0.097) (0.035)
Mean pre-treatment 0.654 0.699 0.256 10.852 9.905
Obs 37,442 37,442 24,724 24,724 24,724
$R^{2}$ 0.12 0.13 0.07 0.07 0.15
  • Notes: This table reports robustness checks on women’s log wage outcomes using the 2017–2024 March CPS sample of dual-parent households with children of child-rearing age. We examine whether our baseline wage results are sensitive to the exclusion of households at the top of the income distribution. Panel A reports the baseline specification without trimming, included for comparison. Subsequent panels show results after progressively trimming the top percentiles of the income distribution. The treatment is based on the husband’s occupation-level WFH shock, as described in Subsection 2.3. All regressions include state-by-time and husband-occupation fixed effects, as well as controls for the wife’s age, number of children, and both spouses’ education levels.
Table B.20: Robustness Checks Using Alternative WFH Shocks and Teleworkable Occupation Overlap
Monthly CPS Sample March CPS Sample
Employed LFP Usual work Part-time Employed LFP Part-time Months Log wages
hours status Status Worked
A. Shock: Difference in average 2023-2024 vs 2019 WFH vacancy rates at Census 2010 4d level (Original)
$Exposed_{o(H)}\times Post_{t}$ 0.024** 0.027*** 0.217 0.004 0.044*** 0.035** -0.030* 0.137 0.075**
(0.011) (0.010) (0.197) (0.008) (0.015) (0.015) (0.016) (0.097) (0.038)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.671 0.715 0.248 10.932 10.007
Obs 252,309 252,309 162,882 161,095 36,158 36,158 24,485 24,485 24,485
$R^{2}$ 0.12 0.12 0.06 0.06 0.12 0.12 0.07 0.07 0.17
B. Shock: Restricting to occupations that original treatment definition and the one from the Teleworkability Index overlap
$Exposed_{o(H)}\times Post_{t}$ 0.039*** 0.042*** 0.466** -0.006 0.054*** 0.045*** -0.030* 0.208* 0.075*
(0.011) (0.010) (0.196) (0.009) (0.016) (0.016) (0.018) (0.107) (0.042)
Mean pre-treatment 0.661 0.675 36.496 0.317 0.660 0.703 0.254 10.903 9.988
Obs 212,610 212,610 135,192 133,635 30,378 30,378 20,240 20,240 20,240
$R^{2}$ 0.12 0.12 0.06 0.06 0.12 0.12 0.08 0.07 0.17
C. Shock: Difference in average 2023-2024 vs 2019 WFH vacancy rates at SOC 2018 3d level
$Exposed_{o(H)}\times Post_{t}$ 0.022** 0.025** 0.271 0.001 0.041*** 0.032** -0.026* 0.152 0.086**
(0.011) (0.011) (0.195) (0.008) (0.015) (0.015) (0.016) (0.097) (0.038)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.671 0.715 0.249 10.932 10.006
Obs 252,311 252,311 162,884 161,097 36,167 36,167 24,501 24,501 24,501
$R^{2}$ 0.11 0.11 0.05 0.05 0.10 0.11 0.05 0.05 0.16
D. Shock: Same as B, at SOC 2018 3d level
$Exposed_{o(H)}\times Post_{t}$ 0.026** 0.030*** 0.320 -0.003 0.045*** 0.036** -0.033* 0.173 0.081**
(0.012) (0.011) (0.203) (0.009) (0.017) (0.016) (0.018) (0.109) (0.041)
Mean pre-treatment 0.661 0.676 36.499 0.318 0.661 0.706 0.254 10.904 9.991
Obs 223,332 223,332 142,064 140,697 31,991 31,991 21,400 21,400 21,400
$R^{2}$ 0.11 0.11 0.05 0.05 0.10 0.11 0.05 0.05 0.16
  • Notes: Panel A defines treated occupations as those where the increase in the share of remote job postings between 2019 and 2023 exceeds the median across all Census 2010 4-digit occupations. Panel B uses the same criterion but limits the sample to occupations classified as teleworkable by Dingel and Neiman (2020). Panels C and D repeat the analyses in Panels A and B, respectively, using broader occupation definitions based on SOC 2018 3-digit codes. All regressions include fixed effects for state-by-time, husband’s occupation, and husband’s industry. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.21: Test: Leave-One Occupation Out using Census 2d
Monthly CPS Sample March CPS Sample
Employed LFP Usual Work Part-time Employed LFP Part-time Months Log wages
hours Status Status Worked
$Exposed\_{o(H)}\times Post\_{t}$ 0.024 0.027 0.217 0.004 0.035 0.027 -0.030 0.120 0.078
Average Bootstrap Estimate 0.024 0.027 0.217 0.004 0.035 0.027 -0.030 0.120 0.078
Average Bootstrap SE 0.011 0.010 0.198 0.008 0.015 0.014 0.016 0.096 0.036
P-Statistic 0.051 0.000 1.000 1.000 0.030 0.604 0.842 1.000 0.040
  • Notes: The table reports the main estimates of the analysis for the labor outcomes of women in the non-ASEC and ASEC sample. Average Boostrap Estimate is the average from the estimates obtained from dropping in every iteration a different occupation at a time. The Average Standard Error is the mean from the standard errors in each iteration from leaving a different occupation out from estimations. The P-statistic captures the percentage of estimates from the iterations that were not significant at the 5% level. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors from each iteration are clustered at the husband’s occupation level.
Table B.22: Placebo Test: Effect of Husbands’ WFH Shocks on Wives Labor Market Outcomes in Households with Older or No Children
Monthly CPS sample March CPS sample
Employed LFP Usual work Part-time Part-time Months Log wages
hours status Status Worked
A. Youngest child is in-between 12 and 18 y.o.
$Exposed_{o(H)}\times Post_{t}$ 0.005 0.006 -0.334 0.010 -0.021* 0.010 0.036
(0.012) (0.012) (0.320) (0.011) (0.012) (0.063) (0.027)
Mean pre-treatment 0.787 0.802 38.463 0.249 0.233 11.055 10.017
Obs 89,499 89,499 66,916 67,740 41,980 41,980 41,980
$R^{2}$ 0.11 0.11 0.09 0.08 0.05 0.06 0.15
B. Youngest child is 18+ y.o.
$Exposed_{o(H)}\times Post_{t}$ -0.018 -0.019 0.696 -0.027 -0.005 -0.129 0.178
(0.022) (0.021) (0.550) (0.020) (0.044) (0.220) (0.117)
Mean pre-treatment 0.773 0.788 39.407 0.213 0.154 11.424 10.001
Obs 33,342 33,342 24,670 24,850 2,158 2,158 2,158
$R^{2}$ 0.16 0.16 0.19 0.17 0.26 0.23 0.27
C. No children sample
$Exposed_{o(H)}\times Post_{t}$ -0.008 -0.004 -0.139 0.008 0.044*** -0.153 -0.041
(0.011) (0.010) (0.212) (0.008) (0.017) (0.120) (0.045)
Mean pre-treatment 0.819 0.837 39.890 0.184 0.130 11.395 10.109
Obs 141,102 141,102 110,607 112,145 12,792 12,792 12,792
$R^{2}$ 0.08 0.09 0.06 0.06 0.09 0.07 0.19
  • Notes: This table reports placebo tests using the 2017–2024 CPS dual-parent household sample. We re-estimate the baseline cross-sectional specification from Equation (1) on subsamples of households where the treatment should have less relevance. Specifically, Panel A restricts to households where the youngest child is between 12 and 18 years old; Panel B to households where the youngest child is older than 18; and Panel C to households without any children. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.23: Effect of Wives’ WFH Exposure on Their Husbands’ Labor Market Outcomes
Occupation shock Industry shock
Employed LFP Usual work Part-time Log wages Employed LFP Usual work Part-time Log wages
hours status hours status
$\text{Exposed (Wife)}\times Post_{t}$ 0.003 0.003 0.092 -0.004 0.011 -0.002 0.001 -0.222 0.006 -0.010
(0.005) (0.004) (0.191) (0.006) (0.033) (0.004) (0.004) (0.200) (0.004) (0.035)
Mean pre-treatment 0.938 0.955 43.930 0.100 10.600 0.936 0.955 43.961 0.101 10.599
Obs 182,081 182,081 160,954 165,928 22,973 189,558 189,558 167,125 172,339 24,641
$R^{2}$ 0.04 0.04 0.05 0.03 0.20 0.04 0.03 0.04 0.03 0.19
  • Notes: This table reports estimates from the cross-sectional specification described in Equation (1), using the 2017–2024 CPS sample of dual-parent households with children of child-rearing age. We examine how increases in wives’ WFH exposure—measured at the occupation or industry level—affect their husbands’ labor market outcomes. The first five columns use occupation-level WFH shocks for wives as the treatment variable; the final five columns use industry-level shocks. All regressions include standard demographic controls (wife’s age, number of children, spouses’ education), as well as fixed effects for state-by-time and the wife’s occupation (or industry, as appropriate). Standard errors are clustered at the wife’s occupation or industry level.
Table B.24: Placebo Test: Effect of Husbands’ WFH Shocks on Wives Labor Market Outcomes Using Fake COVID Timing (March 2018)
Occupation shock Industry shock
Employed LFP Usual work Part-time Log wages Employed LFP Usual work Part-time Log wages
hours status hours status
$\text{Exposed (Husband)}\times Post_{t}$ 0.005 0.004 0.343 -0.014 -0.002 -0.001 0.002 0.456* -0.019* -0.042
(0.009) (0.008) (0.233) (0.010) (0.038) (0.008) (0.008) (0.246) (0.010) (0.038)
Mean pre-treatment 0.664 0.679 36.562 0.326 9.971 0.662 0.679 36.599 0.326 9.976
Obs 178,120 178,120 112,966 111,993 21,507 182,084 182,084 115,314 114,370 20,680
$R^{2}$ 0.12 0.12 0.06 0.06 0.17 0.11 0.11 0.06 0.05 0.16
  • Notes: This table reports results from a placebo test in which the COVID-19 treatment period is artificially shifted to begin in March 2018, three years before the actual onset of the pandemic. The specification follows the baseline cross-sectional model, estimated using the 2017–2019 CPS subsample. The first five columns use occupation-level WFH exposure shocks as treatment, while the last five use industry-level shocks. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.25: Effect of Husbands’ WFH Shock on Wives’ Labor Market Outcomes — ACS
Last Week Last 12 Months
Employment LFP Self-employed Employment Log(wage) Months worked Usual Work Hours Part-time
$Exposed_{o(H)}\times Post_{t}$ 0.019*** 0.022*** -0.006*** 0.024*** 0.053*** 0.124*** 0.566*** -0.035***
(0.005) (0.005) (0.002) (0.005) (0.012) (0.027) (0.117) (0.006)
Mean pre-treatment 0.679 0.699 0.080 0.698 9.956 10.890 35.916 0.480
Obs 541,204 541,204 463,235 516,557 364,841 214,127 382,429 516,557
$R^{2}$ 0.09 0.09 0.01 0.08 0.15 0.03 0.03 0.08
  • Notes: This table reports estimates from Equation (1) using the 2017–2023 ACS sample of dual-parent households with children of child-rearing age. The dependent variables include weekly labor market outcomes—employment, labor force participation, and self-employment status—as well as annual labor market outcomes over the past 12 months, including employment, log wages, months worked, usual weekly hours, and part-time status. All intensive margin outcomes are defined only for employed individuals and are coded as missing for those not working. The treatment indicator is defined at the occupation level based on the change in the probability of working from home between the pre-pandemic period (2019) and the post-pandemic period (2023–2024), as described in Subsection 2.3. All regressions control for the wife’s age, education, and number of children, as well as the husband’s education. Specifications include state-by-time and husband-occupation fixed effects. Standard errors are clustered at the occupation level.
Table B.26: Effects by Age of Children Distribution — ACS
Last Week Last 12 Months
Employment LFP Self-employed Employment Log(wage) Months worked Usual Work Hours Part-time
A. All Households
$Exposed_{o(H)}\times Post_{t}$=1 0.012* 0.014** 0.008*** 0.016*** 0.124*** 0.216*** 0.522*** -0.037***
(0.006) (0.006) (0.002) (0.005) (0.018) (0.022) (0.108) (0.006)
$Exposed_{o(H)}\times Post_{t}\times 1(AgeYC\in[5,8])$ -0.019*** -0.023*** 0.004 -0.019*** -0.061*** -0.128*** -0.294 0.022***
(0.005) (0.005) (0.003) (0.005) (0.018) (0.036) (0.195) (0.006)
Mean pre-treatment 0.679 0.699 0.080 0.698 9.956 10.890 35.916 0.480
$\beta_{1}+\beta_{2}=0$ (p-val) 0.278 0.170 0.000 0.668 0.000 0.007 0.217 0.056
Obs 541,206 541,206 463,237 516,559 364,843 214,127 382,431 516,559
$R^{2}$ 0.07 0.07 0.01 0.06 0.12 0.02 0.02 0.06
B. Households with Two or More Children
$Exposed_{o(H)}\times Post_{t}$=1 0.013* 0.013* 0.009*** 0.015** 0.123*** 0.215*** 0.694*** -0.038***
(0.007) (0.007) (0.003) (0.007) (0.020) (0.029) (0.152) (0.007)
$Exposed_{o(H)}\times Post_{t}\times 1(AgeYC\in[5,8])$ -0.020*** -0.023*** 0.003 -0.017*** -0.065*** -0.135*** -0.485** 0.026***
(0.005) (0.005) (0.004) (0.006) (0.020) (0.045) (0.226) (0.007)
Mean pre-treatment 0.655 0.675 0.087 0.672 9.916 10.854 35.338 0.514
$\beta_{1}+\beta_{2}=0$ (p-val) 0.260 0.156 0.001 0.718 0.000 0.027 0.293 0.149
Obs 387,991 387,991 322,382 370,494 252,018 147,578 265,315 370,494
$R^{2}$ 0.07 0.07 0.00 0.06 0.11 0.02 0.01 0.05
C. Households with One Child
$Exposed_{o(H)}\times Post_{t}$=1 0.009 0.013** 0.006* 0.017*** 0.122*** 0.211*** 0.213 -0.035***
(0.007) (0.006) (0.003) (0.005) (0.018) (0.033) (0.140) (0.006)
$Exposed_{o(H)}\times Post_{t}\times 1(AgeYC\in[5,8])$ -0.011 -0.016** 0.005 -0.018* -0.037 -0.089 0.098 0.006
(0.007) (0.007) (0.006) (0.009) (0.030) (0.066) (0.310) (0.010)
Mean pre-treatment 0.738 0.758 0.064 0.766 10.046 10.969 37.232 0.393
$\beta_{1}+\beta_{2}=0$ (p-val) 0.804 0.676 0.031 0.911 0.003 0.039 0.257 0.009
Obs 153,215 153,215 140,855 146,065 112,825 66,549 117,116 146,065
$R^{2}$ 0.05 0.05 0.01 0.04 0.14 0.03 0.02 0.05
  • Notes: his table reports estimates of baseline outcomes using the 2017–2023 ACS sample of dual-parent households with children of child-rearing age, decomposing the effects by the age of the child. Panel A uses the sample of households with up to 4 children where the youngest child is between 0 and 8 years old. The coefficient on $Exposed_{o(H)}\times Post_{t}$ captures the effect for women in households where the youngest child is between 5 and 8 years old, while the coefficient on the interaction term $Exposed_{o(H)}\times Post_{t}\times 1(AgeYC\in[5,8])$ captures the differential effect for women whose youngest child is between 5 and 8 years old relative to those with a youngest child aged 0 to 4. Additionally, we interact all our baseline fixed effects by age of youngest child. Panels B and C repeat the same analysis, but panel B is estimated on a sample of at households with two or more children (i.e. households with at least 2 children but no more than 4) and panel C uses a sample of households with one child only. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level. We include a test for whether the sum of the coefficients is equal to 0.
Table B.27: Effect of Husbands’ WFH Occupation Shocks on Wives Labor Market Outcomes in the SIPP Data: Cross-Section Specification
Employed Employed LFP Avg work hours Part-time Self-employed Log wages Work at home
1+ weeks all weeks last month (Main Job) (+1 job) (Main job) (Main Job) (Main Job)
A. Baseline estimation
$Exposed_{o(H)}\times Post_{t}$ 0.049** 0.043* 0.055** 0.508 0.007 -0.008 -0.037 -0.015
(0.023) (0.022) (0.023) (0.912) (0.016) (0.021) (0.066) (0.028)
Mean pre-treatment 0.668 0.646 0.682 34.452 0.080 0.102 7.722 0.088
Obs 85,383 85,383 85,383 55,488 57,525 55,488 51,810 55,488
$R^{2}$ 0.20 0.19 0.19 0.18 0.15 0.19 0.28 0.18
B. Interactions by age of youngest child
$Exposed_{o(H)}\times Post_{t}$=1 0.023 0.023 0.029 0.534 0.056* -0.026 -0.102 -0.059
(0.040) (0.040) (0.041) (1.349) (0.032) (0.040) (0.114) (0.043)
$Exposed_{o(H)}\times Post_{t}\times 1(AgeYC\in[0,4])$ 0.033 0.027 0.030 0.247 -0.043 0.038 0.087 0.068
(0.060) (0.058) (0.060) (1.816) (0.041) (0.052) (0.139) (0.055)
Mean pre-treatment 0.667 0.646 0.682 34.450 0.080 0.102 7.722 0.088
Obs 85,461 85,461 85,461 55,614 57,648 55,614 51,947 55,614
$R^{2}$ 0.19 0.18 0.19 0.16 0.14 0.18 0.26 0.17
  • Notes: In Panel A, we estimate the baseline cross-section specification using data from the 2018-2024 SIPP panels. The data is at the the monthly level, and the sample is composed of dual parent households with youngest child aged 0 to 8 years old. Estimations exclude data from March 2020 to March 2021 to partial out Covid-19 temporary shocks. Outcomes include employment and labor force status during the recall month, average work hours, self-employment, log wages and working arrangements in the main job, and whether the respondent has a second job. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level. In Panel B we add interactions between post-treatment status and a dummy for whether the youngest child is 0 to 4 years old; all fixed effects are interacted with this dummy.
Table B.28: Effect of Husbands’ WFH Occupation Shocks on Wives Labor Market Outcomes in the SIPP Data: Panel Specification
Employed Employed LFP Avg weekly Part-time Self-employed Log wages Work at home
1+ weeks all weeks last month work hours (+1 job) (Main job) (all jobs) (Main Job)
A. All SIPP panel samples
$Exposed_{o(H)}\times Post_{t}$ 0.025 0.028 0.044 2.317 0.139* 0.023 0.376*** -0.178*
(0.046) (0.046) (0.048) (1.694) (0.081) (0.016) (0.137) (0.094)
Mean pre-treatment 0.667 0.646 0.682 35.993 0.076 0.101 7.726 0.086
Obs 77,012 77,012 77,012 51,667 52,077 50,252 50,016 50,252
$R^{2}$ 0.85 0.81 0.86 0.79 0.74 0.97 0.86 0.86
B. 2020 SIPP panel sample
$Exposed_{o(H)}\times Post_{t}$ 0.019 0.028 0.036 3.853* 0.163 0.018 0.343** -0.171
(0.055) (0.056) (0.057) (2.005) (0.101) (0.020) (0.155) (0.113)
Mean pre-treatment 0.698 0.677 0.708 35.015 0.065 0.095 7.750 0.107
Obs 7,646 7,646 7,646 5,287 5,370 5,172 5,089 5,172
$R^{2}$ 0.89 0.86 0.88 0.85 0.75 0.99 0.87 0.84
C. Balanced 2019-2022 SIPP panel sample
$Exposed_{o(H)}\times Post_{t}$ -0.079 -0.063 -0.065 3.737 0.183 -0.005 0.492** -0.126
(0.059) (0.061) (0.061) (2.766) (0.148) (0.007) (0.228) (0.148)
Mean pre-treatment 0.724 0.710 0.724 32.970 0.057 0.144 7.775 0.136
Obs 1,978 1,978 1,978 1,383 1,429 1,331 1,349 1,331
$R^{2}$ 0.87 0.83 0.87 0.79 0.57 0.99 0.84 0.70
D. Balanced 2019-2021 SIPP panel sample
$Exposed_{o(H)}\times Post_{t}$ 0.000 0.013 -0.010 3.144 0.076 0.013 0.303 0.158*
(0.052) (0.056) (0.047) (4.229) (0.100) (0.014) (0.197) (0.091)
Mean pre-treatment 0.686 0.675 0.686 33.471 0.049 0.146 7.774 0.151
Obs 2,765 2,765 2,765 1,809 1,860 1,776 1,713 1,776
$R^{2}$ 0.85 0.83 0.87 0.75 0.47 0.99 0.84 0.69
  • Notes: The panel specification replaces cross-section FE with household-level and time-level (i.e. month-year) FE. In addition we also ensure that treatment status has not change from treated to untreated (or viceversa) in the sample (i.e. we enforce that spouses must have been in similarly exposed occupations across any waves they appear). Panel A runs the panel specification on the entire SIPP dataset using cross-section weights. Panel B runs the panel specification with cross-section weights on the sample of household that joined the SIPP on 2020 and thus were covered during reference years 2019 to 2022. Panel C runs the panel specification on a balanced panel of households that appeared every year between 2019 and 2022 using 4-year panel weights. Panel D runs the panel specification on a balanced panel of households that appeared every year between 2019 and 2021 using 3-year panel weights. All regressions include controls for the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.29: Main Results by Age of Youngest Child in the SIPP sample
Employed Employed LFP Avg work hours Part-time Self-employed Log wages Work at home
1+ weeks all weeks last month (Main Job) (+1 job) (Main job) (Main Job) (Main Job)
A. All parents
$Exposed_{o(H)}\times Post_{t}$=1 0.056* 0.050 0.059* 0.781 0.014 0.012 -0.015 0.009
(0.033) (0.032) (0.033) (1.195) (0.022) (0.027) (0.080) (0.035)
$Exposed_{o(H)}\times Post_{t}\times 1(AgeYC\in[5,8])$ -0.033 -0.027 -0.030 -0.247 0.043 -0.038 -0.087 -0.068
(0.060) (0.058) (0.060) (1.816) (0.041) (0.052) (0.139) (0.055)
Mean pre-treatment 0.667 0.646 0.682 34.450 0.080 0.102 7.722 0.088
$\beta_{1}+\beta_{2}=0$ (p-val) 0.566 0.563 0.482 0.693 0.080 0.512 0.373 0.169
Obs 85,461 85,461 85,461 55,614 57,648 55,614 51,947 55,614
$R^{2}$ 0.19 0.18 0.19 0.16 0.14 0.18 0.26 0.17
B. At least second time parents
$Exposed_{o(H)}\times Post_{t}$=1 0.071 0.066 0.069 1.813 0.036 0.012 0.028 0.008
(0.045) (0.044) (0.044) (1.398) (0.029) (0.035) (0.096) (0.044)
$Exposed_{o(H)}\times Post_{t}\times 1(AgeYC\in[5,8])$ -0.057 -0.050 -0.047 -1.899 0.022 -0.046 -0.195 -0.101*
(0.073) (0.072) (0.072) (1.975) (0.053) (0.060) (0.161) (0.061)
Mean pre-treatment 0.646 0.625 0.660 33.780 0.083 0.113 7.681 0.091
$\beta_{1}+\beta_{2}=0$ (p-val) 0.758 0.736 0.635 0.955 0.152 0.472 0.195 0.042
Obs 64,767 64,767 64,767 40,957 42,398 40,957 38,003 40,957
$R^{2}$ 0.22 0.21 0.22 0.18 0.15 0.21 0.28 0.21
C. First time parents
$Exposed_{o(H)}\times Post_{t}$=1 0.058 0.033 0.074 -1.946 -0.030 -0.012 -0.049 -0.049
(0.058) (0.058) (0.058) (2.221) (0.039) (0.043) (0.133) (0.066)
$Exposed_{o(H)}\times Post_{t}\times 1(AgeYC\in[5,8])$ 0.060 0.071 0.021 7.322** 0.051 -0.018 0.202 0.112
(0.127) (0.128) (0.126) (3.462) (0.093) (0.088) (0.305) (0.132)
Mean pre-treatment 0.728 0.704 0.745 36.344 0.066 0.062 7.845 0.081
$\beta_{1}+\beta_{2}=0$ (p-val) 0.313 0.385 0.411 0.043 0.759 0.705 0.579 0.589
Obs 19,810 19,810 19,810 13,901 14,474 13,901 13,248 13,901
$R^{2}$ 0.30 0.29 0.30 0.29 0.38 0.34 0.42 0.34
  • Notes: We replicate the heterogeneity analysis by age of youngest child from Table B.8, but applied on the full 2018-2024 SIPP sample.
Table B.30: Placebo Tests on Main Labor Market Outcomes in the SIPP
Employed Employed LFP Avg work hours Part-time Self-employed Log wages Work at home
1+ weeks all weeks last month (Main Job) (+1 job) (Main job) (Main Job) (Main Job)
A. Placebo: Effect of female spouses WFH occupation shock on their husbands
$Exposed_{o(W)}\times Post_{t}$ 0.018 0.015 0.018 0.253 0.001 -0.043 0.071 0.056**
(0.016) (0.017) (0.014) (0.746) (0.023) (0.027) (0.056) (0.024)
Mean pre-treatment 0.939 0.925 0.954 42.339 0.088 0.124 8.156 0.058
Obs 61,795 61,795 61,795 55,938 57,997 55,938 53,053 55,938
$R^{2}$ 0.15 0.14 0.16 0.14 0.14 0.18 0.27 0.21
B. Youngest child is 12 to 18 years old
$Exposed_{o(H)}\times Post_{t}$ 0.138** 0.136** 0.129** 2.062 -0.014 -0.107* 0.064 0.053
(0.055) (0.056) (0.052) (2.032) (0.043) (0.063) (0.154) (0.050)
Mean pre-treatment 0.780 0.764 0.793 35.889 0.088 0.107 7.706 0.072
Obs 20,614 20,614 20,614 15,682 16,134 15,682 14,767 15,682
$R^{2}$ 0.25 0.24 0.25 0.22 0.15 0.25 0.33 0.24
C. Household has no children
$Exposed_{o(H)}\times Post_{t}$ 0.086** 0.082* 0.069* -5.280*** -0.027 0.007 -0.062 -0.001
(0.043) (0.043) (0.041) (1.347) (0.038) (0.042) (0.102) (0.050)
Mean pre-treatment 0.826 0.809 0.844 38.562 0.098 0.089 7.814 0.074
Obs 25,904 25,904 25,904 20,585 21,420 20,585 19,433 20,585
$R^{2}$ 0.18 0.17 0.18 0.18 0.19 0.23 0.28 0.25
D. Placebo: Moving Covid shock to March 2018
$Exposed_{o(H)}\times Post_{t}$ 0.020 0.018 0.019 -0.696 0.030** 0.018 -0.057 0.021
(0.020) (0.020) (0.020) (0.704) (0.015) (0.017) (0.049) (0.014)
Mean pre-treatment 0.654 0.632 0.671 34.841 0.078 0.099 7.722 0.064
Obs 63,833 63,833 63,833 41,139 42,570 41,139 38,394 41,139
$R^{2}$ 0.20 0.20 0.20 0.19 0.17 0.19 0.31 0.17
  • Notes: Data comes from the 2018-2024 SIPP panel datasets. In Panel A we estimate of female spouses’ WFH occupation shocks on their labor market outcomes, akin to Table B.23. In Panels B and C we estimate our main effect on a sample of households whose youngest children are teenagers, or have no children, respectively, akin to Table B.22. In Panel D, we artificially move the treatment shock as starting in March 2018, as in Table B.24. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.31: Mechanism Effects by Age of Children Distribution
Differential Effect: Youngest Child is 0-4 vs. 5-8 years old
Time Spent in Labor Production Time Spent for Personal Care and Leisure
Working Main Work Main Work Main Work Secondary Work Commuting Personal Care Leisure
From Home Outside Home
Wives:
$Exposed_{o(H)}\times Post_{t}$=1 0.019 0.022 0.069*** -0.047* -0.003* -0.008*** -0.020 0.009
(0.024) (0.024) (0.015) (0.026) (0.002) (0.003) (0.017) (0.015)
$Exposed_{o(H)}\times Post_{t}\times 1(AgeYC\in[5,8])$ 0.099* 0.095 -0.019 0.113** 0.005** 0.014** 0.016 -0.009
(0.059) (0.058) (0.034) (0.046) (0.002) (0.006) (0.019) (0.022)
Mean pre-treatment 0.172 0.170 0.025 0.145 0.001 0.012 0.374 0.113
$\beta_{1}+\beta_{2}=0$ (p-val) 0.021 0.022 0.123 0.067 0.271 0.265 0.753 0.995
Obs 7,964 7,964 7,964 7,964 7,964 7,964 7,964 7,964
$R^{2}$ 0.93 0.93 0.94 0.93 0.99 0.92 0.92 0.94
Husbands:
$Exposed_{o(H)}\times Post_{t}$=1 0.046** 0.057** 0.019* 0.038 -0.006** -0.004 -0.017 -0.020
(0.023) (0.024) (0.011) (0.025) (0.003) (0.006) (0.012) (0.019)
$Exposed_{o(H)}\times Post_{t}\times 1(AgeYC\in[5,8])$ -0.064 -0.073* 0.068* -0.141*** 0.005 -0.006 0.021 0.053
(0.039) (0.038) (0.040) (0.043) (0.005) (0.009) (0.016) (0.039)
Mean pre-treatment 0.317 0.308 0.029 0.280 0.007 0.029 0.350 0.116
$\beta_{1}+\beta_{2}=0$ (p-val) 0.577 0.591 0.025 0.001 0.738 0.254 0.738 0.304
Obs 8,165 8,165 8,165 8,165 8,165 8,165 8,165 8,165
$R^{2}$ 0.94 0.94 0.95 0.94 0.95 0.93 0.94 0.94
Time Spent in Household Production
Housework Cooking Shopping Primary Secondary Total Total Prop. Secondary CC
Childcaring CC while working Secondary CC Childcaring in Total CC
Wives:
$Exposed_{o(H)}\times Post_{t}$=1 0.046*** -0.000 0.007* -0.016 0.005 0.052* 0.036 0.107**
(0.017) (0.008) (0.004) (0.015) (0.074) (0.028) (0.028) (0.044)
$Exposed_{o(H)}\times Post_{t}\times 1(AgeYC\in[5,8])$ -0.104*** -0.020* -0.014** -0.003 -0.024 -0.035 -0.039 0.009
(0.030) (0.012) (0.006) (0.024) (0.089) (0.047) (0.058) (0.061)
Mean pre-treatment 0.093 0.045 0.016 0.106 0.123 0.228 0.334 0.645
$\beta_{1}+\beta_{2}=0$ (p-val) 0.005 0.018 0.120 0.247 0.674 0.628 0.952 0.004
Obs 7,964 7,964 7,964 7,964 4,669 7,964 7,964 7,824
$R^{2}$ 0.95 0.93 0.95 0.93 0.97 0.93 0.93 0.94
Husbands:
$Exposed_{o(H)}\times Post_{t}$=1 -0.003 0.007** -0.000 -0.018 0.088* -0.003 -0.021 -0.044
(0.013) (0.003) (0.003) (0.012) (0.048) (0.028) (0.029) (0.054)
$Exposed_{o(H)}\times Post_{t}\times 1(AgeYC\in[5,8])$ 0.020 -0.000 0.003 0.022 -0.196** 0.054 0.076 0.092
(0.017) (0.007) (0.004) (0.013) (0.077) (0.044) (0.047) (0.090)
Mean pre-treatment 0.041 0.016 0.007 0.051 0.042 0.127 0.177 0.648
$\beta_{1}+\beta_{2}=0$ (p-val) 0.120 0.233 0.263 0.575 0.029 0.111 0.090 0.500
Obs 8,165 8,165 8,165 8,165 7,093 8,165 8,165 7,549
$R^{2}$ 0.92 0.94 0.92 0.93 0.93 0.92 0.92 0.95
  • Notes: We use data from the the 2017-2024 ATUS panel datasets. We estimate our baseline cross-section specification which has state-by-month-year and spouse occupation fixed effects. We show results for the time spent in labor, household and personal activities available in ATUS. Time-use variables are recorded by one member in the household. Then, as Section 2 points out, we have two draws from ATUS: the men-respondent sample and the women-respondent sample. While the former has data on the time spent doing activities by husbands, the latter will have for wives. There is a first panel that has data on the time spent in labor production, capturing the share of the whole day - over 24 hours - that respondents spent: working, working at their main work, working at their main work from home and outside home, working at their secondary work and commuting to work. There is a panel on personal care (i.e. sleeping or taking a shower) and leisure (ie. going to religious events or sport activities). Lastly there is household panel with the time spent doing the following activities: housework, cooking, shopping, primary childcaring (report doing childcaring as main activity), childcaring while working, total childcaring as secondary activity while doing other activities, total childcaring (the sum of primary and secondary childcaring) and the share of secondary childcaring on total childcaring. We restrict our sample to weekdays and non-holiday days. We run Equation 1 adding interactive effects by the age of the youngest child as in Table B.8. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. We include a test for whether the sum of the coefficients is equal to 0. Standard errors are clustered at the husband’s occupation level.
Table B.32: Mean and SD from CPS Compared to Women-Respondent and Men-Respondent in the ATUS Sample
CPS ATUS
Sample Women-Resp Sample Men-Resp Sample
Mean SD Mean SD Mean SD
Panel A: Husband Characteristics
Spouse In Shocked Occupation (Treatment) 0.51 0.50 0.56 0.50 0.61 0.49
Spouse Employed (Last Week) 0.93 0.25 0.95 0.23 0.95 0.22
Spouse LFP (Last Week) 0.96 0.21 0.96 0.19 0.97 0.18
Spouse Usual Work Hours 43.51 9.35 43.88 9.40 43.62 9.05
Spouse Part-time status 0.10 0.31 0.10 0.30 0.12 0.32
Spouse Wage 56103.91 60645.22 62109.25 68648.95 61607.77 55584.11
Spouse - Below HS Level 0.08 0.27 0.05 0.22 0.04 0.21
Spouse - HS Level 0.23 0.42 0.18 0.38 0.15 0.36
Spouse - Above HS Level 0.69 0.46 0.77 0.42 0.81 0.40
Spouse Age 37.67 6.82 37.92 6.33 37.93 6.43
Panel B: Wife Characteristics
Wife Employed (Last Week) 0.66 0.47 0.69 0.46 0.69 0.46
Wife LFP (Last Week) 0.68 0.47 0.71 0.46 0.71 0.45
Wife Usual Work Hours 37.13 10.58 37.44 11.03 36.79 10.83
Wife Part-time status 0.29 0.46 0.30 0.46 0.32 0.47
Wife Wage 37475.39 44037.64 42048.27 47212.46 41085.61 55670.88
Wife - Below HS Level 0.06 0.24 0.05 0.21 0.03 0.18
Wife - HS Level 0.17 0.38 0.11 0.32 0.10 0.30
Wife - Above HS Level 0.76 0.42 0.84 0.37 0.86 0.34
Wife Age 35.23 5.93 35.52 5.35 35.84 5.50
Panel C: HH Characteristics
HH Wage 92792.16 80799.47 99228.70 82494.59 99112.51 74113.41
Number of own children in household 2.08 0.89 2.01 0.86 2.01 0.84
Age of Youngest Child 3.48 2.54 3.39 2.54 3.47 2.55
Age of Oldest Child 7.54 5.10 6.89 4.54 6.87 4.54
Youngest Child 0-4 0.64 0.48 0.66 0.47 0.65 0.48
Youngest Child 5-8 0.36 0.48 0.34 0.47 0.35 0.48
  • Notes: The table reports mean and standard deviations between the different samples: the benchmark CPS sample from the main findings and the ATUS samples. There are two ATUS samples: women respondents sample and men respondents sample. The former is when wives are those that report their time diary of their activities, while the latter is when husbands report them. Panel A captures variables related to husband characteristics, Panel B is related to wives characteristics and Panel C are related to household variables.
Table B.33: Effect of Husbands’ WFH Shock on Wives’ ATUS Labor Market Outcomes
Women-Reported Sample Men-Reported Sample
Employed LFP Usual work Part-time Log wages Employed Usual work Part-time Log wages
hours Status hours Status
$Exposed_{o(H)}\times Post_{t}$ 0.047* 0.034 4.404*** -0.117*** 0.085 0.073** 1.962** -0.046 0.017
(0.027) (0.026) (0.896) (0.031) (0.055) (0.029) (0.860) (0.034) (0.052)
Mean pre-treatment 0.722 0.730 35.894 0.300 10.160 0.666 35.517 0.294 10.266
Obs 7,989 7,989 5,476 5,780 5,274 8,175 5,506 5,688 4,831
$R^{2}$ 0.84 0.86 0.89 0.90 0.93 0.85 0.91 0.92 0.95
  • Notes: We use data from the the 2017-2024 ATUS panel datasets. Time-use variables are recorded by one member in the household. Then, as Section 2 points out, we have two draws from ATUS: the men-respondent sample and the women-respondent sample. We use Labor Market Outcomes of wives from ATUS data, different from the ones from the CPS surveys. While in the women-respondent sample, wives report their own labor outcomes, in the men-respondent, their spouses respond for them. We exploit a fake timing as in Table B.24. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.34: Placebo Timing: Effect on Time Spent in Labor and Household Supply - Share of the Day
Time Spent in Labor Production Time Spent for Personal Care and Leisure
Working Main Work Main Work Main Work Secondary Work Commuting Personal Care Leisure
From Home Outside Home
Wives:
$Exposed_{o(h)}\times Post_{t}$ -0.024 -0.018 -0.003 -0.015 -0.001 -0.006 -0.008 -0.013
(0.025) (0.025) (0.015) (0.030) (0.002) (0.005) (0.012) (0.018)
Mean pre-treatment 0.166 0.165 0.025 0.140 0.001 0.013 0.373 0.108
Obs 4,488 4,488 4,488 4,488 4,488 4,488 4,488 4,488
$R^{2}$ 0.86 0.86 0.88 0.85 0.86 0.84 0.87 0.88
Husbands:
$Exposed_{o(h)}\times Post_{t}$ -0.024 -0.025 0.004 -0.029 0.002 -0.010 0.014 0.031
(0.032) (0.030) (0.015) (0.032) (0.007) (0.008) (0.014) (0.026)
Mean pre-treatment 0.324 0.312 0.026 0.286 0.009 0.036 0.353 0.113
Obs 4,483 4,483 4,483 4,483 4,483 4,483 4,483 4,483
$R^{2}$ 0.88 0.88 0.85 0.87 0.86 0.83 0.87 0.86
Time Spent in Household Production
Housework Cooking Shopping Primary Secondary Total Total Prop. Secondary CC
Childcaring CC while working Secondary CC Childcaring in Total CC
Wives:
$Exposed_{o(h)}\times Post_{t}$ 0.023 0.011 -0.007 0.027 0.062 -0.001 0.026 -0.037
(0.016) (0.007) (0.004) (0.019) (0.097) (0.032) (0.034) (0.045)
Mean pre-treatment 0.097 0.047 0.017 0.110 0.104 0.237 0.347 0.641
Obs 4,488 4,488 4,488 4,488 2,605 4,488 4,488 4,407
$R^{2}$ 0.89 0.86 0.85 0.85 0.93 0.87 0.86 0.86
Husbands:
$Exposed_{o(h)}\times Post_{t}$ -0.012 -0.002 0.003 0.010 -0.031 0.023 0.033 0.026
(0.010) (0.005) (0.003) (0.012) (0.055) (0.039) (0.045) (0.059)
Mean pre-treatment 0.035 0.014 0.007 0.051 0.031 0.115 0.166 0.629
Obs 4,483 4,483 4,483 4,483 3,932 4,483 4,483 4,122
$R^{2}$ 0.86 0.84 0.88 0.84 0.84 0.85 0.85 0.86
  • Notes: We use data from the the 2017-2024 ATUS panel datasets. We show results for the time spent in labor, household and personal activities available in ATUS. Time-use variables are recorded by one member in the household. Then, as Section 2 points out, we have two draws from ATUS: the men-respondent sample and the women-respondent sample. While the former has data on the time spent doing activities by husbands, the latter will have for wives. There is a first panel that has data on the time spent in labor production, capturing the share of the whole day - over 24 hours - that respondents spent: working, working at their main work, working at their main work from home and outside home, working at their secondary work and commuting to work. There is a panel on personal care (i.e. sleeping or taking a shower) and leisure (ie. going to religious events or sport activities). Lastly there is household panel with the time spent doing the following activities: housework, cooking, shopping, primary childcaring (report doing childcaring as main activity), childcaring while working, total childcaring as secondary activity while doing other activities, total childcaring (the sum of primary and secondary childcaring) and the share of secondary childcaring on total childcaring. We restrict our sample to weekdays and non-holiday days. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level. We exploit a fake timing as in Table B.24.
Table B.35: Placebo Test: Effect of Husbands’ WFH Occupation Shocks on Childcare and Child Well-Being Outcomes (COVID Shock Moved to March 2018)
Childcare by Childcare by Childcare by Childcare by Childcare Paid for Log Childcare
Wife while Spouse while Grandparent Relative by Daycare Childcare Expenditure
Working Wife Works
$Exposed_{o(H)}\times Post_{t}$ 0.025 -0.012 -0.037* -0.046** -0.007 0.028 -0.006
(0.028) (0.030) (0.022) (0.022) (0.021) (0.027) (0.095)
Mean pre-treatment 0.319 0.487 0.422 0.462 0.255 0.478 4.636
Obs 44,531 44,531 63,829 63,829 63,829 50,889 24,551
$R^{2}$ 0.14 0.17 0.16 0.15 0.20 0.20 0.36
  • Notes: We replicate the analysis from Table 4 but we only use the 2017-2019 SIPP data, and artificially change the post-Covid definition to be after March 2018. All regressions include fixed effects for state-by-time, husband’s occupation, and husband’s industry. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.36: Effect of Husbands’ WFH Occupation Shocks on Child Well-Being and Activities
Days wife Days spouse Days wife Days spouse Child is in Child takes Child plays
has dinner has dinner takes child takes child club/activity lessons after sports in/out
with child with child on outings on outings after school school of school
A. Entire sample
$Exposed_{o(H)}\times Post_{t}$ -0.029 0.082 0.175 0.080 0.056 0.042 -0.085
(0.086) (0.180) (0.234) (0.245) (0.059) (0.062) (0.060)
Mean pre-treatment 6.532 5.817 3.533 2.677 0.353 0.486 0.535
Obs 44,620 44,620 31,873 31,873 27,705 27,705 27,705
$R^{2}$ 0.18 0.23 0.25 0.24 0.27 0.32 0.30
  • Notes: We use data from the the 2018-2024 SIPP panel datasets. We estimate our baseline cross-section specification which has state-by-month-year and spouse occupation fixed effects. We show results for childcare variables available in the SIPP, which ask households with minor of age children if in a typical week, the following are involved at least one day in looking after children: The wife and her spouse (when the wife is working), a grandparent, another relative (includes grandparent), a daycare. We also include a dummy on whether the family pays for daycare, and self-reported log expenditure in childcare.
Appendix C Discussion on Identification Strategy
Our empirical design estimates the causal effect of an increase in the husband’s work-from-home (WFH) feasibility on the wife’s labor market outcomes. However, in many datasets, including the CPS, occupational information is only recorded for currently employed individuals. This constraint introduces two primary limitations. First, we are unable to systematically observe the wife’s own occupation when she is not employed, preventing us from identifying her exposure to WFH exposure in non-employment spells. Second, this data limitation complicates the disentangling of the husband’s WFH-induced shock from any contemporaneous effects driven by changes in the wife’s own WFH exposure. While the latter concern can be partially addressed by restricting the sample to wives who are employed – allowing us to examine adjustments at the intensive margin – the former limitation precludes a comprehensive analysis of women’s own responses to shifts in their WFH exposure, as this measure is defined only for the employed subsample.
To understand the potential consequences of this limitation, consider a simplified version of our baseline specification:
$y_{i(H)t}=\alpha+\beta_{0}D^{o(H)}+\gamma D_{t}+\beta_{h}D^{o(H)}\times D_{t}+u_{i(H)t}$ (3)
where $D^{o(H)}=1$ if the husband of woman $i$ is in a treated occupation (i.e., one with a large WFH shock), and $D_{t}=1$ for post-COVID observations.
Suppose that unobserved variation in the wife’s own WFH exposure enters the error term:
$u_{i(H)t}=\beta_{1}D^{o(i)}+\beta_{w}D^{o(i)}\times D_{t}+\beta_{2}D^{o(i)}\times D^{o(H)}+\beta_{hw}D^{o(i)}\times D^{o(H)}\times D_{t}+\varepsilon_{i(H)t}$ (4)
where $D^{o(i)}=1$ if woman $i$ herself is in a treated occupation.
Even if we assume that the cross-occupation interaction terms $\beta_{2}$ and $\beta_{hw}$ are zero, the difference-in-differences estimator $\hat{\beta}_{h}$ from  (3) still identifies:
[…]
Even under the assumption that assortative matching patterns are stable pre- and post-COVID (i.e., $\Delta P(D^{o(i)}=1|D^{o(H)}=1)=\Delta P(D^{o(i)}=1|D^{o(H)}=0)=0$), the estimator will still capture some direct effect through $\beta_{w}$.
While we cannot directly estimate or control for the wife’s WFH exposure across the full sample, we show in Section 5 that the observed effects are unlikely to be driven by changes in the wife’s occupational flexibility. For instance, we find that the husband’s WFH shock leads to changes in the wife’s intensive labor supply, whereas her own WFH exposure does not. This supports the interpretation of our estimates as capturing intra-household spillovers, rather than direct responses to the wife’s own occupational exposure.