### v2

# Remote Work and Women’s Labor Supply: The New Gender Division at Home

Abstract
We study how increases in remote work opportunities for men affect their spouses’ labor supply. Exploiting variation in the change in work-from-home (WFH) exposure across occupations before and after the COVID-19 pandemic, we find that increases in husbands’ WFH exposure lead to sizable improvements in wives’ labor-market outcomes: annual employment rises by roughly 2.5 percentage points (from a 69% pre-treatment mean), earnings increase by about 5%, weekly hours worked rise by roughly half an hour, weeks worked increase by about 1.3%, and the likelihood of part-time work declines by approximately 9%. Evidence from time-use diaries and childcare questionnaires suggests these effects are driven by intra-household reallocation of child-caring time: women are less likely to engage in primary childcare activities, while men working at home partially compensate by covering more for their spouse. These results highlight the role of collective household decision-making in shaping the labor market consequences of remote work.
JEL Codes: D13, J13, J22, J16, J31
Keywords: work-from-home, workplace flexibility, labor supply, intra-household allocation, gender gaps, time use, childcare
1 Introduction
Over the past century, gender gaps in education, human capital, and labor market outcomes have narrowed substantially in developed countries (Olivetti and Petrongolo, 2016). Nonetheless, a persistent divergence remains in labor market trajectories following parenthood (Kleven et al., 2024). After childbirth, women frequently reduce their labor force participation or exit employment altogether (Kleven, 2022), and those who remain employed often face large earnings penalties relative to men (Bertrand et al., 2010). These gender disparities are largely driven by (i) the unequal division of childcare responsibilities within households, and (ii) the fact that many jobs - including high-paying white collar positions - have historically lacked flexibility options to allow women to balance both their careers and motherhood (Goldin, 2014).
This paper examines whether the expansion of one aspect of workplace flexibility - namely, remote work opportunities - can shift these intra-household dynamics and encourage greater labor force participation among women with children. We focus on U.S. households with young children and combine occupation-level data on remote work vacancies with rich household-level survey data from the ACS and other national datasets. Leveraging quasi-experimental variation in remote work exposure across occupations before and after the pandemic, we estimate whether large-scale changes in remote work availability altered household labor supply decisions, particularly by enabling mothers to remain in or return to the labor force.
Our motivation builds on the idea, emphasized by Goldin (2014), that the last chapter of gender convergence may depend on restructuring work to allow for greater flexibility. Workplace flexibility has been shown to help parents, especially mothers, accommodate both predictable and unexpected caregiving demands (Angelici and Profeta, 2024). Evidence also shows that women and parents of young children place a high value on flexible work arrangements (Mas and Pallais, 2017; Bloom et al., 2022; Aksoy et al., 2022), and are more averse to commuting (Le Barbanchon et al., 2021). While not synonymous with flexibility, remote work is strongly associated with it, particularly in allowing job tasks to be performed from home, which could enable a more equitable division of household production between spouses.
We propose a household model of labor supply and childcare with remote work options to guide our empirical analysis. Crucially, the opportunity to look after the child while working poses a technological change in household production function which expands the spouses’ choice set and favors increasing labor supply while making childcare division more equitable. Despite this potential, there is limited causal evidence on how remote work affects women’s labor supply, particularly via intra-household channels. Our paper fills this gap by documenting how the expansion of remote work affects their spousal employment and time use.
The backdrop for our analysis is the COVID-19 pandemic, which triggered a rapid and persistent shift in remote work norms. In the U.S., only 7% of full paid workdays were conducted from home in 2019. By June 2023, this share had quadrupled to 28% (Barrero et al., 2023). However, the magnitude of this shift varied widely across occupations: while some sectors (e.g., computer science and professional services) adopted remote work at scale, others (e.g., transport, healthcare, education) largely returned to in-person arrangements. In Figure 1, we illustrate this heterogeneity using data from Hansen et al. (2023) on remote work job postings between 2019 and 2024. We use this variation to study whether and how women’s labor market outcomes respond to changes in household exposure to remote work opportunities, shaped by the husband’s occupation.
Our empirical strategy compares the labor market outcomes of women with young children whose husbands experienced a large increase in the availability of remote work, measured as the change in the share of job vacancies offering WFH in their occupation, to those whose husbands experienced little or no change. Our main hypothesis is that an increase in the husbands’ available WFH options leads to an increase in the wife’s engagement in the labor market. This design offers two main advantages. First, the shock plausibly expands the household’s set of feasible work–care arrangements, creating opportunities for women to enter or re-enter in the labor market. Second, defining treatment through the husband’s occupation avoids a major selection problem: in most cross-sectional surveys, a woman’s occupation is only observed if she is employed. Using changes in WFH exposure in the wife’s occupation as treatment would condition on labor force participation, precluding identification of extensive margin effects – our primary outcome of interest.
Figure 1: Evolution of WFH by occupation (2 digits).
Notes: Author’s elaboration using WFH job postings data from Hansen et al. (2023). Classification based on SOC 2018 2-digits.
Our first-stage estimates indicate that occupation-level increases in WFH probability after COVID translated into substantial increases in men’s realized WFH: take-up rose by about 20 percentage points on average, nearly tripling relative to pre-pandemic levels. In turn, wives’ labor-market outcomes improve along several margins: annual employment rises by roughly 2.5 percentage points (from a 69% pre-treatment mean) or 4%, weekly hours worked rise by roughly half an hour, weeks worked increase by about 1.3%, and the likelihood of part-time work declines by approximately 9%. As a result of this changes in labor supply, overall labor earnings for these women increased by about 5% - with no significant changes in their hourly wage rates. Using a simple decomposition, we show that half of this effect stems from the extensive margin increase and the remainder from the intensive margin.
We then move on to conduct heterogeneity analysis to understand which households have benefited the most. The increases in women’s employment and hours are sharply concentrated among couples with stronger ex ante attachment to remote-capable work: effects are largest for college-educated women (and for women married to college-educated men) and are most pronounced for women ages 25–40, consistent with heightened time constraints during prime childrearing years. Heterogeneity by household composition reinforces a childcare-based channel: responses peak in households with young children—especially when the youngest child is preschool-aged—and attenuate substantially in childless households or those with older children. Finally, we parse out the effect of men’s remote work onto their wives’ labor supply while holding wives’ exposure to remote work fixed through several tests. These show that spousal dynamics and remote work spillovers from one spouse onto the other are economically meaningful, as even women that are unlikely to work from home given their occupation or education degree experienced significant increase in overall labor supply and earnings.
To validate the credibility of our identification strategy, we conduct a broad set of robustness checks that address potential threats to our findings. We examine the possibility of changes in household composition or selection into treatment following the pandemic, including tests for differential fertility and marriage patterns and reweighting strategies that align pre- and post-COVID covariate distributions. We evaluate the sensitivity of our estimates to alternative fixed effect structures, and demographic controls with particular attention to differential trends across major occupational groups, including differences in educational trajectories. Additionally, we explore the robustness of our findings to varying definitions of the post-treatment period and sample restrictions. Across all these exercises, our main results remain consistent, reinforcing the interpretation that increased WFH options for men played a causal role in raising women’s labor market engagement.
Finally, we study how shifts in the household’s allocation of labor supply, childcare responsibilities, and household production activities could be explaining our employment effects. To do this, we rely on mainly on two data sources: the time-use diary data from the American Time Use Survey (ATUS) and the topical modules on childcare from the Survey of Income and Program Participation (SIPP). We complement this data on commuting and remote work from the ACS. Our findings reveal that wives of men that saw large increases in work from home saw a large increase increase their share of day spent working, with no significant change in overall household production. Women themselves are also much more likely to work remote, with approximately a 75% rise relative to the pre -treatment mean, suggesting that remote work enables labor supply for both spouses. Interestingly, we observe that women reduce time spent on childcare as a primary activity and husbands partially with increased childcare while working. Households are also less likely to pay for daycare services. We also find evidence that these adjustments in women’s employment and the allocation of childcare are not coming at the expense of the children’s or women’s well-being.
Overall, our findings suggest that increases in men’s access to remote work can influence their spouses’ labor supply decisions, particularly in households with young children. Our analysis of time-use and childcare patterns offers tentative evidence that the effects may stem from shifts in household responsibilities: women are reducing their primary childcare time and households spend less in daycare services, while husbands modestly seem to provide more primary childcaring and take care of children while women work. Our findings can be rationalized through the lens of a household model where remote work options change the childcare technology and expands households’ feasibility set. Taken together, the results point to the value of analyzing remote work not only as an individual labor market arrangement, but also as a household-level shock, offering new perspective on how WFH access may affect gender gaps in employment.
Related Literature
Our main contribution is to the extensive literature on gender inequality in the labor market (Waldfogel, 1998; Anderson et al., 2002; Ejrnæs and Kunze, 2013; Kleven et al., 2024). Goldin (2014) emphasizes that closing gender gaps requires structural changes in the labor market that enhance flexibility. More recently, Goldin (2022) has argued that the rise of WFH should eventually benefit employment of college-educated mothers. It has been difficult to disentangle net employment effects in the immediate aftermath of the Covid pandemic due to a shortage in the supply of childcare services (Heggeness et al., 2021). By exploiting the long-run effect of Covid on the availability of remote work options, we are among the first papers to find that WFH can lead to closing gender gaps in employment, even in developed country. Our result resembles most closely those by Jalota and Ho (2024), which find in an RCT in India that increased remote work options doubles female labor force participation among data-entry workers.
Second, we contribute to the growing literature on how working arrangements affect labor market dynamics. While existing research has primarily focused on the direct effects of flexible work arrangements on individual outcomes – highlighting their growing value for both workers and firms (Bloom et al., 2015, 2022; Angelici and Profeta, 2024; Choudhury et al., 2024; Mas and Pallais, 2017) – relatively few studies have examined how changes in work arrangements reshape decision-making at the household level. Prior work shows that women and parents place especially high value on flexibility (Mas and Pallais, 2020; Aksoy et al., 2022), yet evidence on intra-household responses remains limited. We contribute to this emerging strand by leveraging a large-scale, quasi-experimental shift in remote work feasibility to study how increased WFH access for men affects household labor supply and home production. [^1] Closest to our work is Harrington and Kahn (2023), who examine how increases in remote work opportunities, as proxied by women’s college attainment, affect mothers’ employment and earnings trajectories In contrast, we estimate the spillover effects of men’s remote work access on their spouses’ labor market behavior and on the intra-household time allocation.
Finally, our paper contributes to the literature on the labor supply effects of parental leave and temporal flexibility policies for parents. Prior work generally finds that such policies have limited, and sometimes even negative, effects on women’s long-term labor market outcomes (Kleven et al., 2024; Ginja et al., 2023, 2020; Canaan, 2022). At best, fathers’ access to temporal flexibility has been shown to improve maternal postpartum health, with little impact on female labor supply (Persson and Rossin-Slater, 2024). How, then, do we reconcile these findings with our own evidence of sizable and persistent effects from increased remote work? We argue that the scale and perceived permanence of the WFH shock fundamentally differ from traditional parental leave policies. Rather than offering short-term accommodations around childbirth, the pandemic-driven shift toward remote work restructured household production possibilities and gave families a broader set of tools to jointly optimize labor supply and caregiving responsibilities over the longer run.
Outline
The remainder of the paper is structured as follows. Section 2 presents a simple household model that rationalizes the effects of increasing remote work options on household labor supply and the division of childcare. Section 3 describes our main data sources and outlines the sample restrictions used to construct comparable treatment and control groups. Section 4 presents our empirical strategy and discusses the identification assumptions underlying our analysis. In Section 5, we report the main results on women’s labor market outcomes. Section 6 provides a series of robustness checks and placebo tests to validate our identification strategy. In Section 7, we explore time-use and childcare data to shed light on the intra-household market and non-market supply changes that can help to explain the observed effects on the lens of a collective problem. Section 8 concludes.
2 A Household Model of Remote Work and Labor Supply
To guide the empirical analysis, we present a simple collective household model that rationalizes household decision-making when the option to work remotely becomes available for households with binding childcare constraints. The model formalizes the idea that work from home expands the household’s feasible set of work-care arrangements by allowing some childcare to be produced jointly with market work, thereby relaxing childcare constraints in a collective decision problem.
2.1 Set-up
Consider a dual-parent household composed of a female ($f$), a male ($m$) and a child. Each spouse allocates one unit of time across market work and primary childcare. For the wife, market work is denoted by $h_{f}$ and primary childcare by $n_{f}$. For the husband, total market work is $h_{m}$, which can be performed either on-site ($h_{m}^{O}$) or remotely ($h_{m}^{R}$), and primary childcare is $n_{m}$. Remote work is subject to a technological/feasibility constraint. Specifically, the husband can perform remotely at most a fraction $\bar{r}\in[0,1]$ of his total market work. [^2] We interpret an increase in $\bar{r}$ as an expansion in remote-work opportunities. This is the main source of variation in our empirical analysis. We allow work hours to be bounded above by some idiosyncratic factor $\bar{H}$.
The child in the household is young and their influence in the household comes solely through the demand for childcare. Specifically, the household must meet a minimum childcare time requirement of $\bar{N}$. Primary childcare produced by either spouse contributes one-for-one. In addition, when the husband works remotely, he can provide secondary childcare concurrently with market work at a rate $\bar{\theta}\in[0,1]$. The childcare requirement is thus
$n_{f}+n_{m}+\bar{\theta}h_{m}^{R}\geq\bar{N}.$ (1)
Consumption $c$ is financed by labor earnings from both spouses. Men and women earn wages $w_{m}$ and $w_{f}$, respectively, whether they work from home or not. The consumption good is a numeraire public good. Both spouses have symmetric separable preferences over consumption, market work, and primary childcare:
$U_{i}=u(c)-\varphi(h_{i})-\chi(n_{i}),\qquad i\in\{f,m\},$ (2)
where $u(\cdot)$ is increasing and concave, and $\varphi(\cdot)$ and $\chi(\cdot)$ are increasing and convex. Importantly, disutility from childcare is tied to primary childcare $n_{i}$; secondary childcare generated while working remotely does not involve additional disutility. [^3] This assumption simplifies the analysis and will lead to a 1-to-1 uptake of remote work option. However, as we show in an extension it is not essential to produce our main qualitative predictions. The household solve a standard collective problem, wherein they maximize the weighted sum $\lambda_{f}U_{f}+\lambda_{m}U_{m}$ subject to the aforementioned budget, remote work, childcare and work hours constraints. In what follows, the Lagrange multipliers for these restrictions are $\beta,\mu,\gamma$ and $\alpha_{i}$, respectively.
2.2 Labor and childcare supply
Because primary childcare enters utility with positive marginal disutility, the household has no incentive to supply childcare by any amount higher than $\bar{N}$. Moreover, given our assumptions, remote work weakly dominates on-site work as it generates childcare without incurring in any disutility. The household exhausts available remote work, so $h_{m}^{R}=\bar{r}h_{m}$. An increase in $\bar{r}$ therefore maps one-for-one into higher remote-work hours for the husband, holding total work $h_{m}$ fixed, and mechanically increases secondary childcare $\bar{\theta}h_{m}^{R}$.
The first-order condition with respect to total male labor $h_{m}:=h_{m}^{R}+h_{m}^{O}$ can be written as
$\lambda_{m}\Big[\varphi^{\prime}(h_{m})-\chi^{\prime}(1-h_{m})\Big]-\gamma+\mu\bar{r}+\alpha_{m}+\beta w_{m}=0.$ (3)
An increase in $\bar{r}$ raises the marginal value of male labor through the term $\mu\bar{r}$: additional market work can now be performed remotely which further relaxes childcare constraint. With additional free time, and absent any cap in work hours, the model predicts an increase in total male labor supply. If the male spouse is constrained in work hours, then changes in $\bar{r}$ do not translate into changes in their labor supply, and the primary margin of adjustment is within male labor: the household shifts work towards $h_{m}^{R}$ but keeping $h_{m}$ and $n_{m}$ constant.
From the binding childcare constraint in (1), total primary childcare $n_{f}+n_{m}$ decreases in WFH capacity $\bar{r}$. This time must be reallocated to market work across spouses and the intra-household allocation between spouses is governed by the ratio of marginal utilities, the Pareto weights and the relative wages. However, when the husband is hours-constrained, the adjustment occurs fully through the wife’s primary childcare, substitute $\bar{\theta}\bar{r}\bar{H}$ away from childcare and into their labor supply.
2.3 Extensions
Labor force participation
To capture potential extensive-margin responses, suppose the wife faces a fixed utility cost $F>0$ from entering in the labor market due to, say, commuting/setup costs, job search frictions, or coordination costs of arranging care. Let $V^{work}(\bar{r})$ be the value of the collective problem when the wife works ($h_{f}>0$), and $V^{out}(\bar{r})$ the value when she does not ($h_{f}=0$). When the husband is hours-constrained, increases in $\bar{r}$ raise the marginal benefit of wives’ participation because the household can convert the additional secondary childcare into reduced primary childcare burden only if the wife reallocates time toward work while the husband maintains fixed hours. Moreover, convexities in $\varphi(\cdot)$ and $\chi(\cdot)$ imply gains from diversification across spouses in market and home production. Consequently, the gap $V^{work}(\bar{r})-V^{out}(\bar{r})$ is increasing in $\bar{r}$, making it more likely that the wife enters employment as remote-work opportunities expand.
Remote work for women
Suppose both male and female spouses can work remote a fraction $\bar{r}_{m}$ and $\bar{r}_{f}$ of their overall labor supply, and that their shocks are correlated so that $r_{i}=s_{i}+\rho$. In the absence of any differences in wages or Pareto weights, an idiosyncratic shock on the female’s remote work options, $s_{i}$, raises their labor supply proportionally more than the male. If the shock affects both spouses homogeneously through $\rho$ - say, if both spouses work in the same occupation - then the time gains from remote work that are shifted into labor are split evenly between both spouses. Overall, the effect of remote work shocks on female labor supply increases when both spouses have the option to work remotely.
Remote work for women can also directly change extensive margin of female labor supply if it reduces the fixed cost of women going into the labor market for the first time or after a long unemployment spell. Households will find it less costly to send the female spouse on the labor market if transportation frictions and costs associated with finding and adapting to a new job are significantly reduced. This can be achieved if fixed costs introduced in the previous extension are some decreasing function of the spouse’s common remote work option, $F=F(\rho)$, or are negatively correlated with it.
Demand for daycare
Suppose spouses can hire daycare services $d$ which contributes 1-for-1 towards the childcare restriction at a price of $p$. For rich households that can afford to hire daycare to fully offest primary childcare - i.e. hire $d=\bar{N}$ -, remote work directly reduces their demand for daycare since, if anything, childcare while working substitutes daycare and frees up resources foregone on consumption. [^4] If increasing hours worked is not optimal for the household, then they will reduce demand for daycare by at least the amount that of work pre-shock can be used for childcare, $\bar{\theta}\bar{r}h^{R}_{m}$, and spend it on consumption. For households that cannot afford enough daycare to fully offset daycare - i.e. $d<\bar{N}$ - the overall effect of remote work on daycare demand is ambiguous. Increasing daycare when the technology to do childcare while working is available is inefficient on its own as the household may be providing more childcare than the requirement $\bar{N}$. However, the household may decide to substitute primary childcare for daycare since labor supply (and consequently, earnings) increase following remote work shocks and childcare generates disutility.
Disutility from secondary childcare
The assumption that secondary childcare generates no additional disutility simplifies the main analysis by ensuring full remote work uptake, i.e. $h^{R}_{m}=\bar{r}h_{m}$. We could relax this assumption, for example, by allowing additional secondary childcare to enter the childcare disutility function as $\chi(n_{m}+\bar{\theta}\bar{r}h^{R}_{m})$. Unlike primary childcare, secondary childcare generates labor earnings. Then the overall choice of remote work depends on whether the utility value of the increase in earnings from switching primary childcare into secondary childcare offsets the marginal disutility of secondary childcare. However, the continuity and convexity in the utility functions ensures that households find it optimal to do some positive amount of remote work. [^5] In other words, it cannot be optimal for the household to supply the same amount of labor and childcare without engaging in any remote work as the potential gains in earnings from working a marginal unit of time remotely have not been redistributed across the household. This would only be an option if there was a discontinuous jump in disutility from the first time unit of secondary childcare. Hence the predictions from the baseline model hold qualitatively.
3 Data
3.1 Data Sources
We draw on several public-use U.S. household surveys for our analysis. Our primary results are based on the American Community Survey (ACS), with complementary evidence from the Current Population Survey (CPS) and the Survey of Income and Program Participation (SIPP). To investigate underlying mechanisms, we use detailed time-use information from the American Time Use Survey (ATUS), with additional evidence from SIPP. To measure changes in WFH exposure at the occupation level, we rely on job postings data on remote and hybrid work arrangements (Hansen et al., 2023), as well as established measures of WFH feasibility by occupation from Dingel and Neiman (2020). We describe each data source in greater detail below.
American Community Survey (ACS)
We use public-use microdata from ACS (Ruggles et al., 2025) samples from 2017 to 2023. The ACS provides detailed information on all household members, including the presence and age of children, as well as occupational codes for both husbands and wives. [^6] Occupation codes at the 4-digit level are based on the Census Occupation Classification and are already harmonized to the 2010 coding scheme. Critically, the ACS provides information on current weekly employment status and labor force participation, as well as retrospective measures covering the previous 12 months. [^7] This contrasts with the CPS, where annual outcomes refer to the previous calendar year. These include indicators for employment, usual hours worked per week, number of weeks worked, part-time status, wages and total salary income (earnings). [^8] Throughout the paper, we define intensive margin outcomes such as wages conditional on employment.
WFH Job Postings Data (Hansen et al., 2023)
Our preferred measure of WFH exposure is derived from job postings data on remote and hybrid work arrangements from 2019 to 2024. These data were constructed by a team of researchers in collaboration with Lightcast, using a large language model (LLM) trained to classify job advertisements based on whether the position allows remote work at least one day per week. The resulting measures are available at the month-year–occupation level [^9] Occupations are classified using the Standard Occupational Classification 2018 (SOC) at 3 digits. and at the month-year- industry level. [^10] Industry codes are classified using North American Industry Classification System (NAICS) 2022 at 3 digits. We use this data to build WFH probability changes by occupation, which will be key for our definition of treatment, as we discuss in Subsection 3.3.111111In Section 6, we also use an alternative measure capturing WFH feasibility, that is the teleworkability index à la Dingel and Neiman (2020).
American Time Use Survey (ATUS)
We use the ATUS (Flood et al., 2023) samples from 2017 to 2024. The ATUS draws its sample from households completing their final CPS interview, ensuring consistency in demographic and labor market variables across surveys. Each ATUS respondent reports detailed time use for a designated 24-hour period, allowing us to examine outcomes such as time spent on childcare, labor and household production. We can also assess the time doing secondary activities, such as the time spent childcaring while doing another activity. These measures are central to investigating the mechanisms underlying the intra-household spillovers we study. Importantly, the ATUS interviews only one respondent per household. By merging the ATUS with the CPS using unique person identifiers, we obtain time-use data for either the wife or the husband, but never both spouses within the same household.
Current Population Survey (CPS) and Survey of Income and Program Participation (SIPP)
We replicate our main findings with public-use microdata from the CPS (Flood et al., 2024) spanning the years 2017–2024, and the 2018-2024 SIPP panel datasets. For the CPS, we integrate data from both the basic monthly files and the Annual Social and Economic Supplement (ASEC), also known as the March files. While weekly measures of employment are available in both the basic monthly and March files, annual employment/labor force participation and income variables are available exclusively in the March files. [^12] See Appendix B of Kleven (2024) for a detailed discussion of differences between the CPS basic monthly and March CPS files, and between weekly and annual variables. In contrast with ACS, annual variables refer to the previous calendar year. In the case of the the SIPP, we exploit the 4-year rotational panel survey. [^13] In 2018 the SIPP redesigned the survey that make it difficult to compare the 2014-2017 SIPP panels and post 2018 panels. The 2018-2024 SIPP panels cover information for reference years 2017-2023. Households are interviewed once a year and asked to recall their detailed monthly employment status and income during the previous year. After the year they join the SIPP sample, households are followed and interviewed for the next three years, after which they are dropped from the sample. Each year new households join the sample, so up to 4 different household cohorts populate each year’s SIPP sample. Importantly, the 2019 SIPP sample experienced significant logistical difficulties, which led the US Census to prioritize resources on following through with the 2018 sample. As a result, the sample of new households that joined in 2019 was not adequately surveyed and the decision was made to not follow up on them in subsequent surveys. Its panel structure allows us to incorporate household fixed effects into our baseline specification to account for time-invariant unobserved heterogeneity. Also, we use SIPP childcare topical modules, which are available for all years in our 2018-2024 SIPP panels, to complement mechanism analysis.
3.2 Sample Restrictions
We restrict our primary analysis, based on the ACS, to dual-parent households in which both the husband and wife reside together and have at least one child of child-rearing age. Specifically, we focus on households with children between the ages of 0 and 8 and exclude households with more than four children. Additionally, we restrict the sample to women aged 20 to 50 and men aged 18 to 64.141414We adopt standard age restrictions for women, as is common in the literature (Kleven, 2024). We additionally impose age restrictions on men to ensure they are of typical labor-force age.
In terms of time coverage, we use ACS files spanning 2017 to 2023, the most recent year of data. The files contain both weekly labor force variables — comparable to those in the basic monthly files of CPS — and additional variables such as annual employment, weeks worked, usual weekly hours worked, part-time status and earnings, which refer to the previous 12 months. As a result, the effective data coverage for regressions related to these annual outcomes is 2016 to 2022.151515Note that occupation is measured/reported at the time of the survey, while annual outcomes refer to the previous 12 months. Therefore, in regressions using these outcomes, we implicitly assume that individuals’ occupation at the time of the survey is the same as in the previous 12 months, which may introduce some measurement error.
3.3 Treated and Control Group Definitions
We use job postings data from Hansen et al. (2023) to construct occupation-level measures of the change in WFH probability before and after the COVID-19 pandemic. Specifically, we compute the difference in the share of hybrid and remote-work job postings between the post-pandemic period (2023–2024) and the pre-pandemic period (2019). [^16] As an alternative, we also construct a measure based on the change in WFH share between two specific months: July 2023 and July 2019. Figure A.1 shows the distribution of our WFH exposure measure.
We convert occupation codes from the original SOC 3-digit classification used in the job postings data to the occupation codes used in the ACS. Occupations with a WFH probability change at or above the median are classified as treated, and those below the median as control. [^17] As a robustness check, we alternatively define treatment based on the median split within the original SOC 2018 3-digit classification. These occupation-level measures capture the pandemic-induced shift in WFH feasibility and are merged to our household survey data based on the husband’s occupation at the time of the survey. [^18] Ideally, we would match individuals to their pre-pandemic occupation to avoid concerns about endogenous selection into more WFH-intensive occupations on the basis of unobservables.
Appendix Table B.1 lists the 15 most common occupations in the treated and control groups. As expected, occupations that experienced the largest increases in WFH probability after COVID-19 include those in the tech sector, law, and management,such as computer and software engineers, lawyers, and various types of managers. In contrast, occupations with minimal changes in WFH feasibility are concentrated in sales, construction, and retail, as well as in sectors like healthcare (e.g., physicians and surgeons) and education (e.g., teachers), where in-person presence remains a key job requirement. We use both the 2019 Occupational Employment Statistics (OES) data (Table B.2) and 2019 ACS data (Table B.3) to compute characteristics of treated and control occupations. As expected, treated occupations are associated with higher wages, both on average and at any percentile of the distribution, and tend to employ more educated workers with longer schooling and longer work hours. Somewhat surprisingly, women appear slightly more represented in treated occupations in the ACS sample, although the difference is modest (about 3 percentage points) and not economically significant.
Using the flexibility indexes from Goldin (2014) - updated with the latest O*NET data - we find little correlation between our treatment definition and these flexibility measures (Figure A.2, Figure A.3), easing concerns about pre-existing occupational differences. As shown in Table B.4, treated occupations are somewhat more WFH-suitable (e.g., less physical proximity, more irregular schedules) but not necessarily more “flexible” in the sense of Goldin’s indexes, which are statistically higher for treated occupations. [^19] Higher index values indicate lower worker substitutability and greater job flexibility. This distinction underscores that while remote work is related to flexibility, the two are not equivalent.
Appendix Table B.5 and Table B.6 present descriptive statistics for baseline labor market outcomes and household-level covariates, respectively. On average, treated households are more likely to have both spouses employed, [^20] For spouse employment and labor force participation measured in the previous week, this pattern is mechanical: the variable equals 1 for all observations because occupation—and hence treatment status—is only recorded for individuals who are currently employed or in the labor force. with both husbands and wives working at higher rates than in control households. As expected, treated households are more likely to have both spouses highly educated and less likely to have attained only a high school degree or less. Treated wives are also slightly older, on average, than their counterparts in the control group. In terms of family composition, treated households tend to have fewer children and younger children on average. Additionally, labor market outcomes for women in the treated group exhibit larger average changes post-COVID compared to women in the control group, though this pattern is purely descriptive. In all regressions, we control for the education levels of both spouses, the wife’s age, and the number of children in the household.
4 Empirical Strategy
Baseline Difference-in-Differences
We exploit variation in the change in WFH probability across occupations induced by the COVID-19 pandemic. Our identification strategy compares labor market outcomes of women whose husbands were employed in occupations that experienced larger shifts in WFH probability, before and after the pandemic. As explained in the previous section, change in WFH exposure is collapsed at the occupation level and reflects the difference in demand-driven WFH probabilities pre- and post-COVID.
Our baseline specification is:
$y_{ist}=(\beta_{1}Post_{t}+\beta_{2}Post_{t}\times During_{t})\times\text{High-}\Delta\text{WFH}_{\,o(h(i))}+X^{\prime}_{it}\delta+\gamma_{st}+\Theta_{o(h(i))}+\varepsilon_{ist}.$ (4)
where $y_{ist}$ denotes a labor market outcome for woman $i$, observed in state $s$ and time tt. [^21] Time is defined at the month-year level in CPS, ATUS, and SIPP. In ACS, where month of interview is unavailable, time is defined at the year level. $\text{High-}\Delta\text{WFH}_{\,o(h(i))}$ is a dummy for treated occupations, whereas $During_{t}$ and $Post_{t}$ are time-dummies for the years 2020-2021 and for years 2020 onwards, respectively. The coefficient on the first interaction term, $\beta_{1}$, captures the treatment effect of interest: it equals one for observations after March 2020 in which the husband’s occupation $o(h(i))$ belongs to the set of occupations that experienced an above-median increase in the probability of WFH between the pre- and post-pandemic periods. The coefficient on the second interaction termm, $\beta_{2}$, captures the effect only during the covid period, from 2020 to 2021.  [^22] In the ACS sample, where data are available only at the annual level from 2017 to 2023, we define the post-period as years from 2020 to 2023, due to the absence of month-level timing and in order to have a sufficient sample size to run our regressions. The during period is set for the 2020 and 2021 period. In the ATUS sample, a subsample from CPS, where we have data from 2017 to 2024, we define March 2020 as the starting point of the post-treatment period and March 2022 as the end date of the during period, in order to have sufficient sample size.
The control vector $X_{i,t}$ includes the woman’s age, education, and number of children, as well as the husband’s education. All regressions include state-by-time fixed effects $\gamma_{s,t}$ to account for local labor market shocks and policy variation, and occupation-of-husband fixed effects $\Theta_{o(h(i))}$ to absorb time-invariant differences across husband’s occupations.
When leveraging the panel structure of one of our datasets, the SIPP, we estimate a panel version of Equation (4) by including household fixed effects. As a result, we remove any control or fixed effect that is time-invariant, differencing out all unobserved and observed factors that do not vary over time within households and/or individuals.
When conducting heterogeneity analysis by age of children or other relevant categories, we interact the treatment term with the categorical variables of interest, omitting the base category, and similarly interact controls and fixed effects to allow for differential trends across groups.
We also conduct industry-level specifications, where we construct a treatment indicator based on whether the husband’s Census 2012 industry ($n(h(i))$) experienced an above-median increase in WFH probability. In these regressions, we replace the occupation-level treatment term in Equation (4) with the industry-level indicator interacted with the post and during-COVID periods. As with the main specification, we include state-by-time and spouse occupation and/or industry fixed effects, along with the standard set of demographic controls.
Identification Assumption
Under a standard parallel trends assumption — namely, that in the absence of the pandemic, the labor market outcomes of women whose husbands worked in occupations that experienced large increases in WFH probability would have followed similar trends to those whose husbands worked in occupations with minimal changes — the coefficient $\beta_{1}$ identifies the causal effect of interest. Specifically, $\beta_{1}$ captures any effect of the husband’s exogenous shift in WFH probability, induced by the pandemic, on the wife’s labor market outcomes.
Disentangling Spillovers from Matching Effects
We categorize the effects on women’s labor outcomes into two categories. First, spillover effects operating through the husband’s increased WFH exposure — for instance, via reallocation of childcare or household production — which can shift the wife’s labor supply even holding her own WFH exposure fixed. Second, correlated direct effects arising from assortative matching in marriage markets, whereby wives of men in highly WFH-feasible occupations may themselves hold (or seek) WFH-feasible jobs, mechanically correlating the husband’s exposure with the wife’s own exposure. A key empirical limitation is that most household surveys record occupation only for the employed; the wife’s WFH exposure is therefore missing precisely when she is non-employed. This constrains our ability to condition on, or interact with, the wife’s occupation in the full sample. [^23] See Appendix C for a detailed discussion of these measurement and identification issues and their implications for separating spillovers from direct effects.
Given these constraints, in Section 5 we adopt complementary strategies to separate and assess the contributions of spillover effects (operating through the husband’s WFH exposure) and correlated direct effects (arising from assortative matching) to the overall estimated effect.
Dynamic Difference-in-Differences
To assess pre-trends and the dynamic evolution of treatment effects, we estimate a dynamic version of Equation (4). Specifically, we estimate the following specification:
$y_{ist}=\sum_{t\neq 2019}\beta_{t}\,\text{High-}\Delta\text{WFH}_{\,o(h(i))}\times D_{t}+X^{\prime}_{it}\delta+\gamma_{st}+\Theta_{o(h(i))}+\varepsilon_{ist}.$ (5)
where 2019 serves as the omitted (base) year. The coefficients $\beta_{t}$ capture the differential evolution of outcomes for women whose husbands were in occupations with large WFH shifts, relative to those in occupations with smaller or no changes, for each year from 2016 to 2023 (excluding 2019). All notation follows as in Equation (4).
5 Main Results
First-Stage: WFH Shock Exposure and Husbands’ WFH uptake
Our main identification strategy exploits how occupation-level shocks in women’s spouses increased women’s labor market participation by plausibly increasing spouse’s time availability for household production. Consequently, our first stage analysis consists of verifying that large shocks on occupation-level remote work options available to fathers of young children in our ACS sample led to sizable increase in their uptake of remote work. By and large we can confirm this claim. Figure 2 shows that men in treated occupation increase their uptake of remote work by 20 p.p with respect to the control group and relative to pre-COVID year 2019. Although the effect decreases over time, it clearly does not to return to pre-covid level and is still 10 p.p.in 2023242424We use as measure of remote work in ACS the following question: How did you usually get to work LAST WEEK. Mark (X) ONE box for the method of transportation used for most of the distance. As discussed in Buckman et al. (2025), this question design yields a measure for the incidence of fully remote work, and aligns with comparable measures from the Survey on Working Arrangements SWAA and other surveys.. With a pre-pandemic average WFH rate of 4%, rates for men have tripled well after the pandemic in our sample households. This increase roughly matches the large increases in share of remote work job vacancies for high exposed occupations that we documented in Figure 1.
Figure 2: Dynamic Effect of Husbands’ WFH Shock on Own Remote Work Uptake
Notes: The figure plots coefficients from a dynamic difference-in-differences specification comparing husbands who were in occupations with large WFH shocks to those in occupations with smaller changes. The omitted year is 2019. All regressions include state-by-time and husband-occupation fixed effects, and control for wife’s age, number of children, and education levels of both spouses. Confidence intervals are shown at the 95% level and are based on standard errors clustered at the husband-occupation level.
Table 1 reports spouses’ WFH uptake estimates from the baseline DiD in (4). On average, husbands’ WFH probability increases by about 13 percentage points (pp) in the post-pandemic period, with larger gains concentrated in the COVID years (2020–2021), as shown by the post-Covid and treatment interaction. We also break down WFH uptake by employment status of the husband. The increase is markedly larger for employees — roughly a doubling relative to pre-COVID levels — whereas the self-employed show a smaller but still sizeable rise of about 40 percent from a higher pre-period baseline of 14 percent. In the final column, we examine commuting time (minutes): men in more exposed occupations reduce commuting by about 5 and a half minutes, roughly 18% of the pre-treatment mean.
Overall, the increase in WFH uptake among men was as large as the shock in available jobs with remote work options available in their occupations. Although the effects are much smaller by 2023 - which corresponds with more companies increasingly mandating return to office among their employees - the shocks have had lasting effects, well after Covid. [^25] ACS captures only fully remote work, but nowadays hybrid options are more prevalent. Moreover, these shocks did not immediately increase men’s intensive margin of labor supply. In Table B.7, we find that the remote work shock did not increase working hours or wages among men, and if anything, men spend marginally less hours of the week working. From the household’s perspective, the shock initially led to a higher and long-term availability of men at households with young children.
Table 1: Effect of Occupation-Level WFH Shocks on Husbands’ Remote Work Uptake
Fully Remote Work Commuting
All Employee Self-employed Time To Work
$Exposed_{o(h(i))}\times Post_{t}$ 0.133*** 0.139*** 0.053*** -5.384***
(0.017) (0.018) (0.012) (0.647)
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ 0.038*** 0.039*** 0.017 -1.247***
(0.007) (0.007) (0.019) (0.265)
Mean pre-treatment 0.044 0.038 0.136 29.688
$\beta_{1}+\beta_{2}=0$ (p-val) 0.000 0.000 0.000 0.000
Obs 704,497 659,974 44,491 704,408
$R^{2}$ 0.17 0.18 0.14 0.06
  • Notes: Estimates from Eq. (4) using husbands’ self-reported telework status as the outcome. The sample is the ACS, 2017–2023. Outcomes refer to the week prior to the interview. We report results for (i) the full sample, (ii) households in which the husband is an employee, and (iii) households in which he is self-employed. The WFH measure is constructed from the commuting-mode question—“How did you usually get to work LAST WEEK? Mark (X) ONE box for the method of transportation used for most of the distance.”—and is interpreted as the incidence of fully remote work; see Buckman et al. (2025). All specifications control for the wife’s age, education, and number of children, as well as the husband’s education. State-by-time and husband-occupation fixed effects are included. Standard errors are clustered at the occupation level.
Effects on Labor Market Outcomes
We hypothesize that expanded WFH opportunities for men relax constraints on household production — both temporal and physical — thereby easing wives’ participation in the labor market while raising children. Table 2 presents estimates from Eq. (4) and shows that shocks on available remote work options for women’s husbands increase women’s labour supply both in the extensive and intensive margins. [^26] We find similar results if we split occupations by quartiles of the treatment exposure. Effects are concentrated and largely similar among the third and fourth quartiles as shown in Table B.8. We find that the probability a wife is employed increases by about 2.5 percentage points — roughly a 4% rise relative to the 69% pre-treatment mean. Total annual earnings increase by around 5%. Weekly hours rise by approximately 0.5 hours, and the number of weeks worked increases by 0.66. Consistent with these intensive-margin responses, the likelihood of part-time work falls by about 2.5 percentage points (9% relative to the pre-treatment mean), while we find no statistically significant effect on hourly wages. [^27] We also evaluate how our effects would change if we defined treatment at the male spouse’s industry occupation-level. Figure A.4 shows time series evolution of industry-level WFH trends between 2017 and 2023, showing largely analogous effects to the occupation-level trends shown in Figure 1. In Table B.10, we present estimates of our baseline specification but using the husband’s industry-level WFH exposure as the treatment variable. The results go in the same direction as those based on occupational exposure, but are lower in magnitude and not as significant. This is consistent with our prior that employer industry may not be as precise as occupation to identify available remote work options. Within a given industry or firm - e.g. hospitals or education -, certain jobs are much more likely to be teleworkable than others - e.g. office admin vs health practitioner.
Because log earnings are observed only for individuals with positive earnings, our earnings, hours, and wage regressions are estimated on the sample of individuals observed working in a given year. In a repeated cross section, this sample includes both individuals who were already employed and individuals who enter employment in the post period, though we cannot distinguish between these groups. To relate these intensive-margin estimates to an unconditional earnings concept, we use a standard two-part decomposition. Unconditional mean earnings can be written as $\mathbb{E}[A]=\Pr(E=1)\cdot\mathbb{E}[A\mid E=1]$. Using a first-order approximation, the treatment effect on log mean earnings is
$\Delta\log\mathbb{E}[A]\approx\Delta\log\Pr(E=1)+\Delta\log\mathbb{E}[A\mid E=1].$ (6)
We approximate the first term by $\beta_{E}/\bar{E}$, where $\beta_{E}$ is the estimated employment effect and $\bar{E}$ is the baseline employment rate in the control group, and the second term using the coefficients from the log earnings decomposition estimated among individuals observed working. [^28] Because earnings among working individuals can be written as $A=w\times H\times WW$, where $w$ denotes the hourly wage, $H$ usual hours worked per week, and $WW$ weeks worked during the year, we decompose changes in log earnings conditional on employment as $\Delta\log\mathbb{E}[A\mid E=1]\approx\beta_{\log w}+\beta_{\log H}+\beta_{\log WW}$. When hours or weeks are measured in levels, coefficients are converted to log changes using a first-order approximation evaluated at baseline means.
Applying this decomposition, we find that intensive-margin responses account for approximately 2.9 log points (46%) of the overall earnings effect, reflecting contributions from hourly wages (0.4 log points), usual hours worked (1.3 log points), and weeks worked (1.2 log points), while the extensive margin contributes an additional 3.4 log points (54%) through increased employment. Together, these components imply an overall earnings effect of about 6.3 log points, which exceeds the coefficient from the conditional log earnings regression. This difference is consistent with newly employed individuals having lower earnings than incumbent workers, so that changes in the composition of the employed sample attenuate the estimated effect on mean log earnings. [^29] Because the log earnings regression is conditional on positive earnings, the composition of the employed sample changes when employment increases. In a repeated cross section, the post-period sample of earners includes both pre-existing workers and new entrants, whom we cannot separately identify. As a result, the decomposition is descriptive and need not sum exactly to the coefficient from the conditional log-earnings regression, even though the earnings identity holds at the individual level.
Appendix Table B.9 reports a complementary specification in which earnings and its components are defined for all individuals, with zeros imputed for non-employed individuals. In this case, the earnings identity holds mechanically and the decomposition is exact. The resulting estimates show that the “wage” component closely mirrors the overall earnings effect, reflecting transitions from zero to positive wages associated with increased employment rather than changes in hourly pay among workers, while changes in hours and weeks play a more limited role.
Table 2: Effect of Husbands’ WFH Shock on Wives’ Labor Market Outcomes
Employment Log Earnings Log Wage Usual Weekly Hours Weeks Worked Part-time
Worked Status
$Exposed_{o(h(i))}\times Post_{t}$ 0.024*** 0.052*** 0.004 0.472*** 0.666*** -0.025***
(0.005) (0.012) (0.009) (0.114) (0.114) (0.005)
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ -0.009* -0.009 -0.005 0.082 -0.236* 0.001
(0.005) (0.012) (0.009) (0.113) (0.120) (0.005)
Mean pre-treatment 0.698 9.956 2.695 36.390 46.060 0.272
$\beta_{1}+\beta_{2}=0$ (p-val) 0.000 0.000 0.857 0.000 0.000 0.000
Obs 729,120 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.08 0.15 0.20 0.03 0.03 0.04
  • Notes: This table reports estimates from Equation (4) using the 2017–2023 ACS sample of dual-parent households with children of child-rearing age. The dependent variables include annual employment, log annual earnings, log (hourly) wage, usual hours worked in the week, number of weeks worked, and part-time status. All intensive margin outcomes are defined only for employed individuals and are coded as missing for those not working. The treatment indicator is defined at the occupation level based on the change in the probability of WFH between the pre-pandemic period (2019) and the post-pandemic period (2023–2024), as described in Subsection 3.3. All regressions control for the wife’s age, education, and number of children, as well as the husband’s education. Specifications include state-by-time and husband’s occupation fixed effects. Standard errors are clustered at the husband’s occupation level.
Figure 3 plots the annual coefficients from the dynamic DiD specification in Equation (5) for our main labor-market outcomes. Employment shows mild pre-trends in 2016–2017, while earnings, hours worked, weeks worked, and part-time status are comparatively flat in the pre-treatment period. Consistent with the null results in the tables, the log-wage event-study exhibits no discernible pre- or post-treatment effects. Overall, the evidence largely supports the parallel trends assumption and indicates that treatment effects strengthen in the post-COVID period. [^30] To assess whether these effects reflect improvements in treated households or declines in control households, Figure A.5 plots adjusted labor-market trends separately for treated and control groups. Employment declines for both groups in 2020; by 2022, however, employment rises sharply among treated households, while the control group returns to pre-pandemic levels. Earnings increase only for the treated group. Hours worked and weeks worked increase – and part-time employment declines – for both groups, with effects larger in magnitude among treated households. Log wages exhibit a temporary narrowing of the treated–control differential between 2018 and 2020, which explains the dip observed in panel (c) of Figure 3. Moreover, effects are persistent in the post-Covid era which matches the lasting effects on the uptake of remote work found in the first stage analysis.
Figure 3: Dynamic Effect of Husbands’ WFH Shock on Women’s Labor Market Outcomes
(a) Annual Employment (b) Log Earnings
Refer to caption Refer to caption
(c) Log Wages (d) Usual Weekly Work Hours
Refer to caption Refer to caption
(e) Weeks Worked (f) Part-time Status
Refer to caption Refer to caption
Notes: Each panel plots coefficients from the dynamic DiD specification shown in Equation 5 in the 2017-2023 ACS sample. The omitted year is 2019. All regressions include state-by-time and husband-occupation fixed effects, and control for wife’s age, number of children, and education levels of both spouses. Confidence intervals are shown at the 95% level and are based on standard errors clustered at the husband-occupation level.
Heterogeneity by Household Demographics
We explore heterogeneity across standard household demographics to shed light on which couples drive the positive effects of spousal WFH shocks on women’s labor supply. In Table B.11, we re-estimate the baseline specification separately by spouses’ education levels. The effect of spousal WFH shocks on women’s labor supply is strongest and most precisely estimated among college-educated women, a group that constitutes the majority of our sample and whose point estimates are closest to the baseline results. When conditioning on husbands’ education, we similarly find that effects are concentrated among women whose spouse has a college degree. This pattern is consistent with college-educated men being more likely to have access to remote work and to use it effectively. Earnings effects for women follow a similar pattern, with the largest and most baseline-consistent estimates observed when either spouse holds a college degree. In addition, in Table B.12 we conduct a parallel heterogeneity analysis by women’s age. The labor supply responses are most pronounced for women between ages 25 and 40, aligning closely with the baseline estimates and consistent with our child-rearing mechanism, as this age range corresponds to peak caregiving responsibilities.
Spousal WFH Shocks and Household Composition
Within our sample of dual-parent households, we investigate whether differences in household composition can give a sense on how the responsibilities of household allocation shape the effect of spousal WFH shocks on women’s labour supply. First, it is sensible to believe that results would differ depending on how young children are within our sample. In Figure A.6, we examine this possibility by estimating heterogeneous effects by age of the youngest child in the household. We estimate our baseline specification on separate ACS subsamples defined by the age of the youngest child in the household. Employment and hours worked effects are largest for households with children aged 2 to 5.313131Effects on weeks worked do not follow a clear monotonic pattern, with larger estimates for households with children aged 0–1 and 4–5. This timing coincides with the period when daycare, pre-K, and early schooling become more salient options. Part-time work effects display a mild U-shape, peaking for households with children aged 2 to 5. Overall, these patterns are consistent with childcare constraints being most binding during the preschool years, easing once children enter school, and being particularly difficult to offset during infancy.
Second, while it is plausible to see women labor supply effects for women with older children or even among couples without children, we hypothesize that in couples without young children we should see smaller effects of spousal WFH shocks on women’s labor supply. We test this in panels Table B.13 where we look at women labor supply effects for households without any children, where the youngest child is teenager or an adult. Effects on employment are at most half the size, and wages, work hours and weeks worked do not change significantly for households without children or where the youngest is an adult. We do, however, see significant employment and work hour effects among households where the youngest child is roughly a teenager (12 to 17 years old) that are fairly similar to our main sample effects, but no changes in earnings or weeks worked throughout the year.
Effects of Spousal WFH Shocks on Men
Our preferred specification uses husbands’ changes in their occupation WFH rates following Covid-19 as an exogenous shock that impacts their wives’ labor market decisions. However, it could wee be that when their wives’ face an increased availability of remote work options, men also respond in their labor supply. While increasing women’s availability at home is not bound to change division of home production, it may still increase the intensive margin of men’s labor supply. To test for potential spillover effects in the opposite direction, Table B.14 inverts the baseline specification and estimates the impact of wives’ WFH shocks on husbands’ labor market outcomes using ACS data. We find no evidence of changes in men’s hours worked, while their wages increase by approximately 2 percent. This pattern is consistent with men moving into better-paying jobs when their spouses experience increased work flexibility. Importantly, we find no evidence of negative income effects for husbands, alleviating concerns that declines in husbands’ labor supply or earnings might drive the observed increases in wives’ labor supply.
Disentangling Spillovers from Matching Effects
As discussed in Section 4, our baseline coefficients combine two channels: the intra-household response generated by men’s WFH exposure (spillover effects) and the component operating through women’s own WFH exposure that is correlated with the former spouse’s shock (correlated direct effects).
As a first step toward isolating women’s own shocks, we augment Equation (4) with women’s occupation fixed effects. In this specification, $\beta$ captures the effect of husbands’ exposure to high-WFH occupations, holding constant wives’ exposure through their own occupation. [^32] By construction, this specification conditions on employed wives, so the estimates pertain to intensive margins only. Table B.15 indicates economically meaningful spillovers, with point estimates very similar to those in Table 2.333333Consistent to the main findings, there is still no effect on hourly wages. Second, we estimate the effect of women’s WFH shocks on their own labor market outcomes. Table B.16 indicates that women do experience an increase in their intensive margin outcomes, mainly in the number of weeks worked. [^34] Due to survey design, we are unable to estimate extensive-margin own-elasticities. Occupation is only reported for individuals who are currently employed or in the labor force, which makes occupation-based WFH exposure mechanically collinear with participation and employment status. As a result, it is not possible to separately identify the causal effect of own WFH exposure on the decision to enter the labor force or become employed.
Neither of the previous strategies allows us, however, to separate correlated direct and spillover effects on the extensive margin of women’s labor supply, nor do they quantify the relative contribution of each channel. To overcome this, we adapt Harrington and Kahn (2023)’s strategy, and leverage women’s college major as an ex ante proxy for WFH probability. For each major, we compute pre- and post-pandemic changes in WFH uptake and define $\text{High-}\Delta\text{WFH}_{e(i)}$ as above the median. [^35] Women without a reported major form a separate bin. We then estimate:
$y_{ist}=$ $\beta_{1}\big(Post_{t}\times\text{High-}\Delta\text{WFH}_{\,o(h(i))}\big)+\beta_{2}\big(Post_{t}\times\text{High-}\Delta\text{WFH}_{\,e(i)}\big)$ (7)
$\quad+\beta_{3}\big(Post_{t}\times\text{High-}\Delta\text{WFH}_{\,o(h(i))}\times\text{High-}\Delta\text{WFH}_{\,e(i)}\big)$
$\quad+X^{\prime}_{it}\delta+\gamma_{st}+\Theta_{o(h(i)),e(i)}+\varepsilon_{ist}.$
In this setup, $\beta_{1}$ identifies intra-household spillovers for women who are ex ante unlikely to WFH (low-exposed majors), $\beta_{2}$ captures women’s direct exposure holding husbands unexposed, and $\beta_{3}$ captures the incremental effect when both channels are turned on, disentangling the correlated direct effects of women.
Table B.17 shows that there exist pure spillovers effects explaining part of the increase in the labor supply of women. Treated households with women with low-exposed majors experience an increase in employment of 1.1 p.p (about 1.5% of the pre-period mean), a 2.6 p.p reduction in part-time likelihood, and increase their working time by 43 minutes and 0.65 weeks. There is a 3.5% increase in earnings, while, as in 5, there are no effects in wages. Among high-exposed majors, the joint effect of the spillover effects and the assortative ones imply a 3.9 p.p. employment increase, a 3.1 pp reduction of part-time and more than an additional hour of work. Together, the evidence indicates that the spillover effects play a central role in explaining the increase in women’s labor market outcomes, while the residual effects operate through correlated direct effects arising from assortative matching.
6 Robustness and Validation
Group Composition and Selection
We conduct a series of robustness checks to assess whether our main results are driven by changes in group composition between treated and control households before and after the pandemic. Given the cross-sectional nature of the ACS, one potential concern is that our findings may not reflect intra-household adjustments—for example, a woman increasing her labor supply in response to her husband’s WFH shift—but instead reflect changes in assortative matching. Specifically, it is possible that, post-COVID, men in high–WFH-shock occupations are more likely to be partnered with women who are already more attached to the labor force.
We address this concern by examining changes in assortative matching over time. Figure A.7 in the Appendix provides suggestive evidence that changes in spousal composition are unlikely to be driving our main results. While there is clear evidence of assortative matching, i.e., households in which the husband is in the highest quartile of WFH exposure are more likely to have wives in similarly high-exposure occupations than households in which the husban is in the first quartile, this pattern appears stable over time. Specifically, the distribution of wives’ exposure conditional on their husbands’ treatment quartile does not shift meaningfully between the pre- and post-COVID periods. As discussed in Subsection 3.3, occupational WFH exposure is strongly correlated with wives’ income and labor force attachment. The stability of the matching pattern across periods is therefore reassuring and suggests that selection into treatment via assortative matching is not biasing our estimates. In addition, we examine potential changes in key demographic outcomes - namely, marriage, divorce, and fertility. Table B.18 reports estimates from specification (4), where the dependent variable is an indicator for whether the woman got married, divorced, or had a child within the 12 months preceding the interview date. Across all outcomes, we find no statistically significant effects of the husband’s WFH shock, suggesting no differential selection into or out of marriage, nor into fertility, between treated and control groups following the pandemic.
Another potential concern is that the composition of households with young children may have shifted after the pandemic. If couples choosing to have children post-COVID differ systematically in income, education, or labor-market attachment, our estimates could partly reflect selection rather than causal effects. To assess this, Figure A.8 plots changes in the probability of having a child and in the number of children across quartiles of husbands’ WFH exposure, before and after COVID. The figure shows that households in the lower two quartiles of husbands’ exposure became slightly more likely to have a child and to have more children than those in the upper quartiles, but the differences are very economically small. Consistently, Table B.19 reports regression results from Equation (4) showing no meaningful effects on fertility outcomes, and column (3) of Table B.18 indicates no statistically significant change in the probability of having a child in the past 12 months.
To address concerns about endogenous changes in group composition over time, we implement reweighting strategies that adjust the post-COVID sample to match the distribution of pre-COVID covariates. We estimate a single logit model for the probability of an observation belonging to the pre-COVID period, using spouse occupation, wife’s age and education, and state as predictors. [^36] The predicted probabilities are obtained from the following specification: $\Pr(\text{Pre-COVID})=f_{s}+f_{o(H)}+f_{e}+f_{a}+\varepsilon,$ where $f_{s}$, $f_{o(H)}$, $f_{e}$, and $f_{a}$ denote state, husband’s occupation, education, and age fixed effects, respectively. We then plot the predicted probabilities for pre- and post-COVID observations to verify common support (see Figure A.9). We construct DiNardo et al. (1995) (DFL) weights as p/(1p)p/(1-p) and inverse-propensity score (IPW) weights as $1/p$, where $p$ is the predicted probability of being in the pre-COVID period. [^37] Both sets of weights are normalized to sum to one and scaled by ACS sampling weights. As shown in Table B.20, the reweighted estimates remain consistent with our baseline results, alleviating concerns that shifting sample composition drives our findings.
Treatment Definition Robustness
A potential concern is that our baseline treatment definition — based on changes in the average within-occupation WFH job postings rate between the post-covid period in 2023 and 2024 and the pre-covid period in 2019 at the Census 2010 4-digit occupation level — may be driving the results. To address this, we examine the robustness of our findings to a range of alternative treatment definitions in Table B.21.
The estimates reported in Panel B confirm that smoothing the share of remote work job vacancies using a three-month moving average does not affect the treatment assignment, and henceforth, the resulting coefficients. Similarly, the results in Panels C, E, F, and G, which use alternative classification schemes and different measures of WFH change, yield estimates that are closely aligned with those obtained under the baseline treatment.
As an alternative measure of remote work potential, we incorporate the teleworkability index developed by Dingel and Neiman (2020) in Panel D and H, which classifies occupations based on their suitability for being performed from home. [^38] The index is constructed using responses from two ONET surveys: the Work Context Questionnaire and the Generalized Work Activities Questionnaire. These surveys assess the importance of physical presence in task performance, allowing for an evaluation of whether specific tasks within each occupation could feasibly be performed remotely. The final occupation-level indicator is calculated as the average teleworkability score across tasks within each occupation. [^39] Consistent with our baseline job-posting measure, we define treatment as an indicator equal to one if the husband’s occupation is at or above the median of the teleworkability distribution. In contrast to our primary measure based on changes in job postings, this index captures the inherent feasibility of telework by occupation, rather than shifts in employer demand for remote work arrangements. [^40] Importantly, this classification does not account for demand-side frictions or social norms that may affect the practical implementation of remote work. For instance, while Dingel and Neiman (2020) classify teachers as capable of WFH, post-pandemic trends suggest that in-person instruction remains the preferred mode, largely due to parental demand (Barrero et al., 2023). The panels yield nearly indistinguishable estimates from those based on the benchmark treatment. Finally, Table B.22 replicates our main results using an alternative WFH measure based on the share of husbands reporting WFH in the ACS at the occupation–year level. Also in this case, the results are quantitatively and qualitatively unchanged.
Taken together, these findings indicate that our results are not sensitive to how treatment is defined, and that the estimated effects are robust across a range of reasonable alternatives.
Specification and Sample Robustness
We test in several ways the robustness of our main women employment results to minor changes in the specification and sampling choices. First, we test the robustness to the choice of fixed effects. By design, the correct specification should include at least an after Covid-19 dummy and a treatment dummy. In our baseline specification, we generalize these to year-by-state fixed effects and spouse occupation fixed effects to gain more precision. In Table B.23, we make minor alterations to the fixed effect structure in our main estimations for the ACS sample. We change the occupation group definitions, the geographical area, the level of interactions between state fixed effects, and others. By comparing all other results to Panel E - our baseline specification - a clear pattern emerges in that coefficients do not change significantly with the fixed effects interactions.
Second, our results may be driven by differential trends across broad industries or occupations. Table B.24 addresses this issue by incorporating year-by-broad-occupation and year-by-broad-industry fixed effects. The estimates remain statistically significant in nearly all specifications, with annual employment being the only outcome that exhibits a modest attenuation in some cases. In addition, Table B.25 controls for heterogeneous educational trends by comparing occupations within the same quintile of average years of schooling. The treatment effects persist, with annual employment being the one with the most attenuation. Taken together, these results indicate that our findings are not driven by differential educational trends across occupations, nor by other unobserved characteristics within broad occupations or industries.
Third, we perform other exhaustive analysis: (a) varying the demographic controls of the households in Table B.26 and Table B.27; (b) restricting the sample to non self-employers and trimming top income households in Table B.28 and Table B.29 respectively; and, (c) dropping major occupation groups in Table B.30. Overall, all these findings indicate that our main results are not sensitive to the choice of controls, or sample restrictions both at the household and occupation level.
Placebo tests
We conduct a standard timing falsification test by artificially assigning the onset of the COVID shock to 2018 and estimating the baseline specification on the 2017–2020 ACS subsample, with reference years 2016–2019. We define the pre- and post-shock periods as 2016–2017 and 2018–2019, respectively. Because this subsample predates the COVID period and the shock was largely unanticipated, we would not expect to observe treatment effects in this placebo setting. Results from this falsification exercise are reported in Table B.31. We find no evidence of effects on earnings, hours worked, weeks worked, or part-time status around this artificial treatment date. We do observe a small effect on log wages, as well as on employment, which are consistent with the mild pre-trends and patterns documented in the dynamic DiD plots.
Evidence from Alternative Datasets
Estimate from alternative datasets generally support our results. First, we replicate our analysis using the 2017-2024 CPS sample, as described in Section 3. Although the CPS offers a smaller sample than the ACS, it provides comparable labor-market measures that allow an independent check of our main findings. [^41] Because CPS outcomes are measured monthly, we define the “during” and “post” periods as described in Section 3 for the ATUS sample. Namely, we define March 2020 as the starting point of the post-treatment period and March 2022 as the end date of the during period. Consistent with this, Table B.32 reports reduced-form estimates for wives that closely mirror the ACS results. Women in higher-exposure households are 2.7–3.5 percentage points more likely to be employed, depending on whether employment is measured over the previous week or year. They report higher weekly hours, are roughly 3 percentage points less likely to work part-time, and earn approximately 8 percent higher wages.
Second, we replicate several of our main employment results for women using SIPP data and exploit its four-year panel structure to estimate panel regressions. Specifically, we replace the baseline fixed effects with month–year and household fixed effects. Results for several samples are reported in Table B.33.424242In addition to restricting the sample to dual-parent households with children aged 0–8, we further limit the analysis to households that do not switch treatment status across panel interviews. Although the SIPP contains information from 2017 to 2023, we focus on the 2020 SIPP panel, which follows households from reference years 2019 to 2022 and is the only panel spanning both the pre- and post-COVID periods. [^43] Due to logistical constraints, the SIPP did not continue following households newly added in the 2019 sample.
Across specifications, the estimated effects of spousal WFH shocks on women’s labor supply have the same sign as in the ACS and CPS, though the estimates are substantially noisier. Extensive-margin effects are similar in magnitude but less precisely estimated, while hours worked and earnings effects are statistically significant but larger than those observed in the other datasets. We interpret these inflated magnitudes cautiously, as they are likely driven by limited statistical power: the panel analysis is based on a much smaller sample, with approximately 500 unique households.
Summary
Taken together, these robustness and falsification checks support the validity of our identification strategy. The estimated effects are not driven by differential selection, sample restrictions, or alternative treatment definitions. Main estimates across different nationally representative datasets are similar either sign, although panel estimates are noisy.
7 Household Specialization
The remote work shock may have changed how households coordinate their market and home production decisions. Understanding and evaluating the changes in household specialization and time use is central to reconcile not only women’s increase labor supply, but also interpret the overall effect of the WFH shock on households.
In this section, we argue that the large employment effects we observe for women with young children can be explained by changes in the household production function, particularly in the burden and organization of childcare duties. As discussed in Section 5, these changes may arise from two forces: (i) spillovers from husbands’ increased presence at home and (ii) women’s own access to work from home through assortative matching.
To explore how the intra-household allocation of time adjusts in response to the WFH shock, we estimate Equation (4) using variables from the ATUS, ACS and SIPP that capture both the intensive and extensive margins of labor supply and household production. Our goal is to assess how time spent on household activities and childcare is redistributed within households, explaining with the lens of a collective perspective the observed increase in labor market participation among women in treated households. [^44] For the ATUS, we restrict the sample to respondents who completed the time-use diary on a weekday and non-holiday day, in order to capture household dynamics on a typical workday. [^45] We evaluate the comparability between the two ATUS samples and the CPS, given that the ATUS samples are non-overlapping subsamples drawn from the CPS. Table B.38 presents the first and second moments of the distribution for key characteristics of husbands, wives, and households across the CPS sample and both ATUS subsamples. The CPS sample displays modest differences relative to the ATUS samples, particularly in the likelihood of treatment and in measures of household income and education. These patterns suggest that the ATUS subsamples tend to capture relatively wealthier and more educated households. [^46] The estimated effects might be driven by specific subsamples whose responses to the WFH shock diverge from those identified in the ACS results in Table 2 or the CPS results in Table B.32. We address this in two ways. First, both Table B.34 and Figure 4 show that the increase in women’s employment following the WFH shock is statistically significant and persistent. Second, despite the smaller sample sizes in ATUS (mainly on the ASEC variables like wages, earnings and weeks worked), Table B.39 shows that the estimated effects on wives’ labor market outcomes are directionally consistent with those in CPS sample, which are also aligned with our main findings using the ACS sample.
Market Production Time Use
First, husbands experience a substantial change in their work location, consistent with being the directly treated group as Figure A.10 illustrates. These sheds more light on the validity of the first stage. They increase time working from home by 4.7 percentage points and appear to reduce time working outside the home. Although the latter estimate is noisy due to limited sample size, Figure 4 shows a negative and relatively stable effect across all post-treatment years. These shifts in work location are persistent over time, unlike the more transitory patterns observed for women. Notably, men do not exhibit any change in total hours worked. However, treated men consistently report reduced commuting time by 5 minutes on average. Efficiency-wise these results mean that men are working the same amount of time, but avoiding part of the sunk costs from commuting. [^47] For the ATUS sample, the reduction is of around 17 minutes.
Wives experience an increase in the time allocated to market work, which aligns with the labor supply effects Section 5 documents. As shown in Figure A.10, treated women increase on average the share of their day spent working after the pandemic by 7.5 percentage points, i.e., equivalent to nearly two additional hours. [^48] A full day includes 24 hours. This rise is driven mainly by a 5 percentage point increase in working from home, which is consistent with the 6.5 pp. increase in WFH uptake we report in Figure 4. Consistent with the fact that most of increase in working is coming from work from home uptake, there is a significant reduction in the time spent commuting for most of the post-treatment years by almost 10% - 2.5 minutes. [^49] In Table B.34, the ATUS commuting variable is not significant, which it is not striking. It is more difficult to detect an effect on a substantially smaller sample size and on an outcome whose variation is small, relative to others like share of time working. By performing the disentanglement of spillovers from correlated direct effects as in Section 5, we observe that most of women’s increase in WFH uptake and reduction in commuting comes from the correlated direct effects, as shown in Table B.35. This result is alligned with the assumption proposed in the extension of Section 2 in which women can also hold remote work jobs.
Household Production Time Use
This change in the joint labor market choices came alongside with new patterns in the collective household production, as Figure 4 documents. Treated households exhibit meaningful shifts in the allocation of childcare responsibilities.
Although the total time spent caring for children does not change significantly, there is a clear shift in how childcare is provided. Women in treated households reduce the amount of time spent on childcare as a primary activity by 2.7 pp - an approximately 25 percent decrease relative to the average pre-treatment values. While they are not reporting to be spending significant more time child-caring while working, although they are spending more time WFH. Additionally, there is no reduction in secondary childcare, as Table B.34 shows. In other words, women mainly cut back on primary childcare. As a result, secondary childcare becomes more prominent: its share of total childcare time rises by 8.4 p.p. relative to a pre-treatment mean of 64.5 percent.
Husbands, on average, do not exhibit significant changes in total time spent on overall childcare. However, as shown in Figure 4, their time spent on childcaring while working increases significantly by 1.5 percentage points. This is a 200% increase relative to baseline. This aligns with the fact that men are more likely to work from home due to their shock. These patterns point that women reduce time spent on primary childcare, and men appear to partially compensate while working. In terms of household division, there is a more equal distribution in the time spent childcaring between spouses. The gender gap in childcare time was of approximately 0.5 in the pre-treatment, as Table B.34 reflects, meaning that women spent the double of time childcaring compared to men. Treated households narrow the gap to approximately 0.75.505050On the household production side, both men and women do not spend significantly any change in time in housework. Although this estimate is not precise for women, it consistently points in a negative direction, while virtually zero for men. Childcare might be more substitutable and proportionally more important in time to be able to detect significant changes, compared to other household tasks. This might explain the lack of precise effects on the other less prominent household activities.  [^51] We conduct a standard placebo test by assigning the onset of the post-treatment period to March 2018. Table B.40 reports the results. We find no significant changes.
Figure 4: Dynamic Effect of Husbands’ WFH Shock on Women’s WFH Uptake and Households’ Time-Use
(a) Women WFH Uptake (ACS)
Women Men
(b) Commuting (ACS)
Refer to caption Refer to caption
(c) Primary Childcare (ATUS)
Refer to caption Refer to caption
(d) Childcare while Working (ATUS)
Refer to caption Refer to caption
Notes: The top center panel shows effect of spousal WFH shock on women’s WFH uptake. Starting from the second row of subfigures, each panel to the left plots coefficients from a dynamic difference-in-differences specification comparing women whose husbands were in occupations with large WFH shocks to those in occupations with smaller changes. Panels to the right compare husbands directly. The omitted year is 2019. All regressions include state-by-time and husband-occupation fixed effects, and control for wife’s age, number of children, and education levels of both spouses. Confidence intervals are shown at the 95% level and based on standard errors clustered at the husband’s occupation level.
Childcare in SIPP
Using the SIPP childcare topical modules, we are able to provide complementary evidence how household change their sources of childcare and hiring of daycare following shocks on spousal WFH options. We present some results in Table B.36. We estimate our baseline cross-section specification on the 2018-2024 SIPP panel datasets on a battery of childcare related variables. Our results show that in households shocked by the occupation-level WFH treatment, when the child’s mother is working, the spouse is 6.4 pp more likely to be involved in childcare, or roughly a 13% increase over the baseline. On the other hand, the woman herself is not any more likely to be involved in childcare.
We also find that households exposed to the WFH shock are significantly less likely to rely on grandparents or other relatives, or on daycare, for day-to-day childcare. Treated households are 7.3 pp less likely to have any paid childcare, and although imprecisely estimated, their total childcare expenditure falls by about 16%. The dynamic DiD estimates in Figure A.11 show clear post-2019 declines in both the probability of paying for childcare and log childcare spending. Overall, these patterns are consistent with fathers becoming more available for childcare in treated households.
Household’s Well-being
As Table B.34 points out, women in treated households experience no significant change in time spent on personal care activities, such as sleeping and grooming. In contrast, they exhibit a 1.9 pp increase in time devoted to leisure activities, suggesting that the shift in labor supply and the new household specialization decisions do not fully crowd out their discretionary time. This may suggest that women are not worse-off in the sense that they are not taking more responsabilites by crowding out their personal time. For husbands, however, there is a significant reduction in leisure time, amounting to 2.4 percentage points.
Finally, it is plausible that children may be worse off given that our main results suggest that women in our treatment sample are more likely to be working, spend more time at work and, as found in this section, there is an overall reduction in the household’s time spent child-caring as a primary activity. Although we cannot fully measure well-being with our current datasets, the SIPP has a handful related variables such as days that parents have dinner or go out with their child and the children’s involvement in extracurricular activities. In Table B.37, we estimate our baseline specification on these outcomes, and we find no significant changes for these variables. This evidence suggests that – to the extent we can measure well-being – children are not worse off by changes in household production and labor supply.
Summary
Overall, these results underscore the role of intra-household reallocation in explaining the observed improvements in women’s labor market outcomes. Couples seem to be collectively opting for women to reduce their primary childcare time in order to increase their working time, and husbands modestly compensate them as they are spending more time at home. In addition, there is a decline in reliance on extended family or formal daycare services. These changes are not coming at the cost of any worsening the children’s or women’s well-being from the related variables we can observe.
8 Conclusion
In this paper, we provide novel evidence on the intra-household labor market and household production spillovers induced by large and permanent changes in WFH arrangements of male spouses in dual parent households. Exploiting plausibly exogenous variation in the post-pandemic expansion of remote work across U.S. occupations, we show that increases in husbands’ WFH exposure lead to sizable gains in wives’ labor market outcomes—higher employment, wages, and hours worked, and a lower likelihood of part-time work. These effects are concentrated among households with children of child-rearing age and mostly highly educated spouses and are not driven by changes in group composition, assortative matching, and are widely robust to specification changes and several falsification test. Our analysis of time use and childcare arrangements indicates that the increase in women’s labor supply is partly facilitated by an intra-household reallocation of time. In particular, we find evidence of modest substitution in childcare responsibilities between spouses: women reduce the time they devote to primary childcare, while husbands slightly increase their involvement.
We highlight the importance of household-level dynamics when evaluating the labor market implications of job amenities like WFH. Our findings suggest that access to remote work arrangements for one partner can alleviate time constraints for the other, with implications for gender equality in labor force attachment and earnings. In future research, we hope to explore larger datasets covering the universe of U.S. households to better disentangle mechanisms underlying our results and follow-up on career progression for women entering the labor force.
Appendix
Appendix A Figures
Figure A.1: Distribution of WFH Exposure Using Job Postings Data
Notes: Author’s elaboration using WFH job postings data from Hansen et al. (2023). The figure shows the distribution of the difference in the share of hybrid and remote-work job postings between the post-pandemic period (2023–2024) and the pre-pandemic period (2019). In the paper, we define occupations with a WFH probability change at or above the median are classified as treated, and those below the median as control. We convert occupation codes from the original SOC 3-digit classification used in the job postings data to the occupation codes used in the ACS.
Figure A.2: Average of 5 Normalized Job Characteristics (O*Net)
(a) Pre
(b) Post
Notes: This figure plots the average of five normalized job characteristics from O*NET 2025 across Census 2010 occupations, using ACS male employment weights. The index—originally proposed by Goldin (2014)—includes Time pressure, Contact with others, Establishing and maintaining interpersonal relationships, Structured work, and Freedom to make decisions. Higher values indicate greater job inflexibility. Panel (a) uses pre-Covid weights (ACS 2019), while Panel (b) uses post-Covid weights (ACS 2023). The relationship between treatment intensity and this composite measure is relatively flat.
Figure A.3: Average of 7 Normalized Job Characteristics (O*Net)
(a) Pre
(b) Post
Notes: This figure extends the five-dimension index by adding two additional O*NET characteristics—Face-to-face discussions and Frequency of decision making—to construct a seven-dimension measure of job inflexibility, following Goldin (2014). Higher values indicate more structured, less substitutable jobs. Values are weighted by ACS male employment at the occupation level. Panel (a) uses 2019 weights (pre-pandemic), while Panel (b) uses 2023 weights. As with the five-variable index, treatment status is not systematically related to this flexibility measure.
Figure A.4: Evolution of WFH by industry (2 digits).
Notes: Author’s elaboration using WFH job postings data from Hansen et al. (2023). Classification based on NAICS 2022 2-digits.
Figure A.5: Adjusted Labor Market Trends by Treatment Group
(a) Annual Employment (b) Log Earnings
Refer to caption Refer to caption
(c) Log Wages (d) Usual Weekly Work Hours
Refer to caption Refer to caption
(c) Weeks Worked (d) Part-time Status
Refer to caption Refer to caption
Notes: Each panel plots adjusted yearly trends in labor market outcomes for treated (blue line) and control (red line) groups. These estimates are the predicted values from separate regressions by group:
$y_{i(H),s,t}=f_{s}+f_{o(H)}+\sum_{\begin{subarray}{c}y=2017\\ y\neq 2019\end{subarray}}^{2023}\beta^{y}\cdot 1(\text{year}(t)=y)+X_{i(H),t}\delta+\epsilon_{i(H),s,t},$
where outcomes are residualized on state and husband-occupation fixed effects and controls for wife’s age, number of children, and both spouses’ education. The omitted year is 2019.
Figure A.6: Effect of Husbands’ WFH Shock on Wives’ Labor Market Outcomes, by Age of Youngest Child
(a) Annual Employment (b) Weekly Work Hours
Refer to caption Refer to caption
(c) Weeks Worked (d) Part-time Status
Refer to caption Refer to caption
Notes: In each panel we estimate the baseline specification on 8 different subsamples of the ACS by age of the household’s youngest child.
Figure A.7: Assortative Matching by Spousal WFH Exposure, Pre- and Post-COVID
(a) Pre-COVID (b) Post-COVID
Refer to caption Refer to caption
Notes: Each panel shows the distribution of wives across treatment quartiles conditional on their husband’s occupation being in a given WFH shock quartile, separately for the pre-COVID (panel a) and post-COVID (panel b) periods, using pooled ACS data from 2017 to 2023 with children of child-rearing age. Quartiles are based on the occupation-level change in WFH probability described in Subsection 3.3. The figure is designed to assess whether assortative matching by WFH exposure changed differentially following the pandemic.
Figure A.8: Changes in Fertility Outcomes by Husbands’ WFH Shock
(a) Probability of Having a Child (b) Number of Children
Refer to caption Refer to caption
Notes: Each panel plots the change in fertility-related outcomes following the onset of COVID-19 across quartiles of the husband’s WFH exposure, defined by the occupation-level change in WFH probability as described in Subsection 3.3. Panel (a) displays the probability of having a child; Panel (b) reports the average number of children in the household. The figure uses pooled ACS data of dual households from 2017 to 2023 and is intended to assess whether fertility patterns shifted differentially across exposure groups in the post-pandemic period.
Figure A.9: Kernel Density Plot Showing Common Support for Reweighting
Notes: This figure uses the 2017–2023 ACS sample of dual-parent households with children of child-rearing age. It plots the estimated probability that an observation is from the pre-COVID-19 period, obtained from a logit model on a set of covariates including spouse occupation, wife’s age and education, and state. The predicted values are derived from separate regressions by group:
$\Pr(\text{Pre-COVID})=f_{s}+f_{o(H)}+f_{e}+f_{a}+\varepsilon,$
where $f_{s}$, $f_{o(H)}$, $f_{e}$, and $f_{a}$ represent fixed effects for state, husband’s occupation, wife’s education, and age bin, respectively. The densities indicate good overlap in covariate distributions before and after COVID, supporting the use of reweighting strategies such as DiNardo-Fortin-Lemieux (DiNardo et al. 1995) and inverse-propensity score weighting.
Figure A.10: Dynamic Effect of Husbands’ WFH Shock on Share of the Day Spent in Other Labor and Household Production, by Year
(a) Working (ATUS)
Refer to caption Refer to caption
Women Men
(b) Mainwork From Home (ATUS)
Refer to caption Refer to caption
Women Men
(c) Mainwork Outside Home (ATUS)
Refer to caption Refer to caption
Women Men
(a) Housework (ATUS)
Refer to caption Refer to caption
Women Men
Notes: Each panel from the first column plots coefficients from a dynamic difference-in-differences specification comparing women whose husbands were in occupations with large WFH shocks to those in occupations with smaller changes. Those in the second column compare husbands directly. The omitted year is 2019. All regressions include state-by-time and husband-occupation fixed effects, and control for wife’s age, number of children, and education levels of both spouses. Confidence intervals are shown at the 95% level and based on standard errors clustered at the husband’s occupation level.
Figure A.11: Effect of Husband’s WFH Shock on Daycare Hiring (SIPP)
(a) Paid for Daycare
(b) Log Daycare Expenditure
Notes: Each panel plots coefficients from Equation 5 using the 2018-2024 SIPP dataset. The omitted year is 2019. All regressions include state-by-time and husband-occupation fixed effects, and control for wife’s age, number of children, and education levels of both spouses. Confidence intervals are shown at the 95% level and are based on standard errors clustered at the husband-occupation level.
Appendix B Tables
Table B.1: Top 15 ACS Occupations by Treatment Status
Treated Occupations Control Occupations
Managers, nec (including Postmasters) (9.0%) First-Line Supervisors of Sales Workers (6.4%)
Software Developers, Applications and Systems Software (5.6%) Driver/Sales Workers and Truck Drivers (6.4%)
Computer Scientists and Systems Analysts/Network systems Analysts/Web Developers (3.9%) Construction Laborers (3.4%)
Police Officers and Detectives (3.4%) Elementary and Middle School Teachers (3.0%)
Chief executives and legislators/public administration (2.8%) Physicians and Surgeons (2.8%)
Lawyers, and judges, magistrates, and other judicial workers (2.5%) Electricians (2.7%)
General and Operations Managers (2.4%) Laborers and Freight, Stock, and Material Movers, Hand (2.6%)
Sales Representatives, Wholesale and Manufacturing (2.4%) Carpenters (2.4%)
Financial Managers (2.3%) Retail Salespersons (2.4%)
Managers in Marketing, Advertising, and Public Relations (2.2%) Secondary School Teachers (2.3%)
Constructions Managers (2.2%) Postsecondary Teachers (2.3%)
Accountants and Auditors (2.1%) Chefs and Cooks (2.3%)
Engineers, nec (1.8%) Other production workers including semiconductor processors and cooling and freezing equipment operators (2.1%)
First-Line Supervisors of Production and Operating Workers (1.7%) First-Line Supervisors of Construction Trades and Extraction Workers (2.1%)
Computer and Information Systems Managers (1.7%) Janitors and Building Cleaners (2.0%)
  • Notes: This table lists the 15 most common occupations in the treated and control groups, respectively, based on the classification variable occ2010 from the ACS. The treatment dummy is assigned at the occupation level, as described in Subsection 3.3. An occupation is classified as treated if the change in the share of remote/hybrid job postings before and after COVID-19, as measured using data from Hansen et al. (2023), is above the median across all occupations. Frequencies are expressed as percentages of the total number of observations within each group. Occupation labels correspond to the value labels associated with occ2010.
Table B.2: Difference between treated and control occupations using Occupational Employment Statistics (OES) 2019
Treated Control Difference S.E. N
Panel A: Total Employment and Annual Wage
Total Employment 349,041.91 585,176.13 -236,134.23 166,736.68 344
Mean annual wage 83,621.97 51,115.15 32,506.82*** 8,439.12 343
Annual 10th percentile wage 42,130.15 29,890.22 12,239.93*** 3,447.29 343
Annual 25th percentile wage 55,385.42 37,056.22 18,329.20*** 5,247.30 343
Annual median wage (50th percentile) 75,556.84 47,040.04 28,516.80*** 7,940.83 340
Annual 75th percentile wage 94,046.61 61,363.60 32,683.01*** 8,254.64 338
Annual 90th percentile wage 108,187.12 78,071.17 30,115.95*** 8,415.95 329
Panel B: Hourly Wage
Mean hourly wage 40.39 23.87 16.53*** 4.19 336
Hourly 10th percentile wage 20.60 14.04 6.57*** 1.69 336
Hourly 25th percentile wage 27.14 17.36 9.78*** 2.57 336
Hourly median wage (50th percentile) 36.88 21.99 14.89*** 3.92 333
Hourly 75th percentile wage 45.58 28.59 16.99*** 4.12 332
Hourly 90th percentile wage 52.25 36.45 15.79*** 4.21 324
  • Notes: Balance table comparing treated and control occupations. All statistics are weighted by ACS employment counts at the Census 2010 occupation level, using husbands’ occupations from survey year 2019. OES variables are reported at SOC 2018 (six-digit) in the source and were aggregated to Census 2010 occupations using employment-share weights; the OES reference period is May 2019. * $p<0.10$, ** $p<0.05$, *** $p<0.01$.
Table B.3: Difference between treated and control occupations using ACS 2019
Treated Control Difference S.E. N
Panel A: Occupational Characteristics (ACS)
Share of Women 0.41 0.40 0.01*** 0.00 182,286
Age (mean) 37.06 35.81 1.25*** 0.04 182,286
Share of Below Any College 0.10 0.37 -0.27*** 0.00 182,286
Share of Below Any High-School 0.02 0.11 -0.10*** 0.00 182,286
Share of College Graduates 0.68 0.32 0.35*** 0.00 182,286
Potential Years of Education 15.13 13.80 1.33*** 0.01 182,286
Potential Experience 15.93 16.01 -0.09** 0.04 182,286
Earnings 57,146.44 35,772.07 21,374.38*** 279.23 182,286
Wage 28.59 19.48 9.12*** 0.26 182,286
Usual Weekly Work Hours 41.49 40.91 0.58*** 0.07 182,286
Weeks Worked 49.33 48.42 0.91*** 0.06 182,286
Share Working Part-Time 0.13 0.15 -0.02*** 0.00 182,286
  • Notes: Balance table comparing treated and control occupations at the Census 2010 occupation level, using the ACS analysis sample for survey year 2019. All statistics are weighted by survey weights. The reference year is 2019 for all variables except wages, which refer to earnings in the previous year. * $p<0.10$, ** $p<0.05$, *** $p<0.01$.
Table B.4: Difference Between Treated and Control Occupations using O*NET 2025
Treated Control Difference S.E. N
Panel A: Work Context
Time Pressure 0.02 0.27 -0.25 0.19 342
Frequency of Conflict Situations 0.59 -0.02 0.61*** 0.21 342
Contact with Others 0.44 -0.00 0.44** 0.22 342
Coordinate or Lead Others 0.67 -0.05 0.71** 0.30 342
Degree of Automation 0.18 -0.36 0.55*** 0.15 342
Structured Work -0.55 0.19 -0.74*** 0.18 342
Duration of Typical Work Week 0.51 0.24 0.27 0.22 342
Face-to-Face Discussions 0.32 0.10 0.22 0.19 342
Freedom to Make Decisions 0.44 -0.22 0.66*** 0.18 342
Frequency of Decision Making 0.28 0.25 0.03 0.19 342
Impact of Decisions on Coworkers or Company Results 0.40 0.13 0.27 0.20 342
Level of Competition 0.40 -0.10 0.50*** 0.18 342
Physical Proximity -0.25 0.20 -0.45** 0.23 342
Regular Work Schedule -0.32 0.12 -0.44** 0.20 342
Work with a Workgroup or Team 0.48 -0.30 0.78*** 0.26 342
Panel B: Work Activities
Communicating with Supervisors, Peers, or Subordinates 0.74 -0.28 1.02*** 0.19 342
Coordinating the Work and Activities of Others 0.70 -0.22 0.92*** 0.23 336
Developing and Building Teams 0.75 -0.16 0.91*** 0.23 338
Establishing and Maintaining Interpersonal Relationships 0.74 -0.23 0.98*** 0.18 342
Panel C: O*Net Indexes (Goldin, 2014)
Avg. of 5 normalized characteristics 0.22 -0.00 0.22*** 0.08 342
Avg. of 7 normalized characteristics 0.24 0.05 0.19* 0.10 342
  • Notes: Balance table comparing treated and control occupations at the Census 2010 occupation level, using the ACS analysis sample restricted to survey year 2019, matched to the most recent O*NET occupational characteristics (2025). All statistics are weighted by ACS 2019 male employment counts at the Census 2010 occupation level. All O*NET variables are standardized to $z$-scores (demeaned and divided by the cross-occupation standard deviation). The O*NET indices of Goldin (2014) are constructed as the average of the normalized dimensions Time pressure, Contact with others, Establishing and maintaining interpersonal relationships, Structured work, and Freedom to make decisions (five-characteristic index); the seven-characteristic index additionally includes Face-to-face discussions and Frequency of decision making. Higher values of the indices indicate lower worker substitutability and greater job inflexibility, following Goldin (2014).
    Measurement note: The current “Structured work” item differs from that used in the original period due to O*NET wording changes: the present question asks “How much freedom does the worker have in determining the tasks, priorities, or goals of the job?”. We use this item and take its complement so that higher values correspond to more structure, aligning with the original construct.
    * $p<0.10$, ** $p<0.05$, *** $p<0.01$.
Table B.5: Descriptive Statistics of Main Labor Market Outcomes, by Treatment Group and Pre/Post-COVID Period
Pre-COVID Post-COVID
Mean SD N Mean SD N
Panel A: Control Sample
Spouse Employed (Last Week) 0.97 0.17 152,710 0.96 0.19 179,033
Spouse LFP (Last Week) 0.99 0.11 152,710 0.98 0.13 179,033
Spouse Employed (Last Year) 1.00 0.00 152,710 1.00 0.00 179,033
Spouse Log Earnings 10.26 0.84 152,710 10.25 0.88 179,033
Spouse Log Wage (Last Year) 2.64 0.71 152,710 2.69 0.76 179,033
Spouse Usual Weekly Hours Worked (Last Year) 43.98 10.76 152,710 43.43 11.02 179,033
Spouse Weeks Worked (Last Year) 49.14 7.53 152,710 48.90 9.42 179,033
Spouse Part-time Status (Last Year) 0.07 0.25 152,710 0.07 0.26 179,033
Wife Employed (Last Week) 0.65 0.48 152,710 0.66 0.47 179,033
Wife LFP (Last Week) 0.67 0.47 152,710 0.69 0.46 179,033
Wife Employed (Last Year) 0.67 0.47 152,710 0.68 0.46 179,033
Wife Log Earnings 9.77 1.13 152,710 9.81 1.16 179,033
Wife Log Wage (Last Year) 2.51 0.76 152,710 2.57 0.80 179,033
Wife Usual Weekly Hours Worked (Last Year) 36.41 11.00 152,710 36.43 11.31 179,033
Wife Weeks Worked (Last Year) 45.93 12.05 152,710 46.02 13.08 179,033
Wife Part-time Status (Last Year) 0.27 0.44 152,710 0.26 0.44 179,033
Panel B: Treated Sample
Spouse Employed (Last Week) 0.99 0.12 171,168 0.98 0.14 226,211
Spouse LFP (Last Week) 0.99 0.08 171,168 0.99 0.09 226,211
Spouse Employed (Last Year) 1.00 0.00 171,168 1.00 0.00 226,211
Spouse Log Earnings 10.84 0.78 171,168 10.83 0.82 226,211
Spouse Log Wage (Last Year) 3.15 0.71 171,168 3.17 0.75 226,211
Spouse Usual Weekly Hours Worked (Last Year) 45.10 9.49 171,168 44.24 9.63 226,211
Spouse Weeks Worked (Last Year) 50.33 5.21 171,168 50.51 6.76 226,211
Spouse Part-time Status (Last Year) 0.03 0.17 171,168 0.03 0.18 226,211
Wife Employed (Last Week) 0.68 0.46 171,168 0.70 0.46 226,211
Wife LFP (Last Week) 0.70 0.46 171,168 0.72 0.45 226,211
Wife Employed (Last Year) 0.70 0.46 171,168 0.72 0.45 226,211
Wife Log Earnings 10.11 1.18 171,168 10.16 1.17 226,211
Wife Log Wage (Last Year) 2.83 0.76 171,168 2.87 0.80 226,211
Wife Usual Weekly Hours Worked (Last Year) 36.64 11.87 171,168 37.07 11.50 226,211
Wife Weeks Worked (Last Year) 46.35 11.45 171,168 46.78 12.19 226,211
Wife Part-time Status (Last Year) 0.26 0.44 171,168 0.24 0.42 226,211
  • Notes: This table reports descriptive statistics for key labor market outcomes across treated and control households, before and after the COVID-19 pandemic. The sample corresponds to the pooled 2017–2023 CPS basic monthly files, restricted to dual-parent households with children of child-rearing age. All estimates are computed using the CPS sampling weights.
Table B.6: Descriptive Statistics of Main Control Variables, by Treatment Group and Pre/Post-COVID Period
Pre-COVID Post-COVID
Mean SD N Mean SD N
Panel A: Control Sample
Spouse - Below HS Level 0.14 0.35 152,710 0.13 0.34 179,033
Spouse - HS Level 0.30 0.46 152,710 0.30 0.46 179,033
Spouse - Above HS Level 0.56 0.50 152,710 0.57 0.50 179,033
Spouse Age 36.68 7.03 152,710 37.02 7.00 179,033
Wife - Below HS Level 0.11 0.32 152,710 0.09 0.29 179,033
Wife - HS Level 0.21 0.41 152,710 0.21 0.41 179,033
Wife - Above HS Level 0.67 0.47 152,710 0.70 0.46 179,033
Wife Age 34.26 6.06 152,710 34.61 6.00 179,033
Number of own children in the household 2.16 0.92 152,710 2.15 0.92 179,033
Age of Youngest Child 3.42 2.57 152,710 3.44 2.58 179,033
Age of Oldest Child 8.02 5.29 152,710 8.00 5.36 179,033
Youngest Child 0-4 0.65 0.48 152,710 0.64 0.48 179,033
Youngest Child 5-8 0.35 0.48 152,710 0.36 0.48 179,033
Panel B: Treated Sample
Spouse - Below HS Level 0.02 0.13 171,168 0.02 0.13 226,211
Spouse - HS Level 0.09 0.29 171,168 0.09 0.29 226,211
Spouse - Above HS Level 0.89 0.31 171,168 0.89 0.31 226,211
Spouse Age 37.83 6.31 171,168 38.12 6.19 226,211
Wife - Below HS Level 0.02 0.15 171,168 0.02 0.15 226,211
Wife - HS Level 0.09 0.28 171,168 0.09 0.28 226,211
Wife - Above HS Level 0.89 0.31 171,168 0.89 0.31 226,211
Wife Age 35.72 5.51 171,168 36.07 5.41 226,211
Number of own children in the household 2.02 0.86 171,168 2.02 0.86 226,211
Age of Youngest Child 3.35 2.58 171,168 3.40 2.59 226,211
Age of Oldest Child 6.92 4.75 171,168 6.97 4.79 226,211
Youngest Child 0-4 0.66 0.47 171,168 0.65 0.48 226,211
Youngest Child 5-8 0.34 0.47 171,168 0.35 0.48 226,211
  • Notes: This table presents summary statistics for key household- and individual-level control variables across treated and control groups, before and after the COVID-19 pandemic. The statistics are based on the pooled 2017–2023 ACS analysis sample, restricted to dual-parent households with children of child-rearing age. All estimates are computed using the CPS sampling weights.
Table B.7: Effect of Husbands’ WFH Shock on Their Own Labor Market Outcomes
Log Earnings Log Wages Weekly Work Hours Weeks Worked Part-time Status
$Exposed_{o(h(i))}\times Post_{t}$ -0.005 0.010 -0.580*** -0.096 -0.004*
(0.008) (0.008) (0.132) (0.074) (0.002)
Mean pre-treatment 10.588 2.930 44.740 49.750 0.046
Obs 729,120 729,120 729,120 729,120 729,120
$R^{2}$ 0.34 0.33 0.09 0.03 0.06
  • Notes: We estimate Equation (4) on the ACS 2017-2023 sample on male spouse intensive margin labor supply outcomes.
Table B.8: Heterogeneous Effects on Women’s Labor Market Outcomes by Quartile of Husbands’ WFH Shock
Employment Log Earnings Log Wages Usual Weekly Hours Weeks Worked Part-time
Worked Status
$Exposed_{o(h(i))}(Quartile2)\times Post_{t}$ 0.007 0.023 0.002 0.169 0.143 -0.004
(0.007) (0.018) (0.013) (0.172) (0.187) (0.008)
$Exposed_{o(h(i))}(Quartile3)\times Post_{t}$ 0.019*** 0.051*** 0.003 0.469*** 0.620*** -0.026***
(0.006) (0.016) (0.012) (0.144) (0.140) (0.006)
$Exposed_{o(h(i))}(Quartile4)\times Post_{t}$ 0.034*** 0.067*** 0.006 0.604*** 0.744*** -0.028***
(0.008) (0.015) (0.012) (0.150) (0.143) (0.007)
Mean pre-treatment 0.702 9.969 2.716 36.389 46.030 0.271
Obs 729,120 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.08 0.15 0.20 0.03 0.03 0.04
  • Notes: This table reports estimates from Equation (4), where treatment intensity is captured by quartiles of the husband’s occupation-level change in WFH probability, as defined in Subsection 3.3. The sample is the 2017–2023 ACS sample of dual-parent households with children of child-rearing age. The baseline (omitted) category is the bottom quartile of WFH exposure. For simplicity we only present coefficients for the interactions of the treatment and post-Covid time dummies. All regressions control for the wife’s age, education, and number of children, as well as the husband’s education. We include state-by-time and husband-occupation fixed effects. Standard errors are clustered at the occupation level.
Table B.9: Effect of Husbands’ WFH Shock on Wives’ Labor Market Outcomes - Unconditional on Employment
Employment Earnings Wage Usual Weekly Hours Weeks Worked Part-time
Worked Status
$Exposed_{o(h(i))}\times Post_{t}$ 0.024*** 0.058*** 0.051** 1.193*** 1.586*** -0.036***
(0.005) (0.015) (0.025) (0.226) (0.259) (0.006)
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ -0.009* -0.008 -0.013 -0.258 -0.599** 0.007
(0.005) (0.014) (0.034) (0.198) (0.240) (0.005)
Mean pre-treatment 0.698 24278.840 14.465 25.417 32.171 0.491
$\beta_{1}+\beta_{2}=0$ (p-val) 0.000 0.000 0.165 0.000 0.000 0.000
Obs 729,120 729,120 729,120 729,120 729,120 729,120
$R^{2}$ 0.08 0.19 0.12 0.09 0.09 0.08
  • Notes: This table reports estimates from Equation (4) using the 2017–2023 ACS sample of dual-parent households with children of child-rearing age. In this regression, we keep the whole sample across all estimations. For those unemployed women, we impute zeroes on all outcome variables, with the exception of part-time, that we impute a 1. Earnings and wage regressions are estimated using Poisson quasi-maximum likelihood. All regressions control for the wife’s age, education, and number of children, as well as the husband’s education. We include state-by-time and husband-occupation fixed effects. Standard errors are clustered at the occupation level.
Table B.10: Effect of Husbands’ Industry-Level WFH Shock on Wives’ Labor Market Outcomes
Employment Log Earnings Log Wages Usual Weekly Hours Weeks Worked Part-time
Worked Status
$Exposed_{n(h(i))}\times Post_{t}$ 0.013** 0.010 0.007 0.044 -0.076 0.002
(0.006) (0.014) (0.008) (0.128) (0.146) (0.004)
$Exposed_{n(h(i))}\times Post_{t}\times During_{t}$ -0.005 0.020 -0.005 0.221* 0.361** -0.012***
(0.004) (0.013) (0.010) (0.123) (0.146) (0.004)
Mean pre-treatment 0.692 9.950 2.686 36.475 46.054 0.269
$\beta_{1}+\beta_{2}=0$ (p-val) 0.036 0.000 0.778 0.006 0.014 0.003
Obs 758,105 530,048 530,048 530,048 530,048 530,048
$R^{2}$ 0.08 0.14 0.19 0.03 0.03 0.03
  • Notes: This table reports estimates from Equation (4), using the 2017–2023 ACS sample of dual-parent households with children of child-rearing age. Unlike in the main specification, treatment is defined using the husband’s industry (classified at the 2012 Census industry level), based on the change in remote job postings before and after the COVID-19 pandemic. Treated households are those in which the husband’s industry experienced a WFH probability increase above the median. We replace husband-occupation fixed effects with husband-industry fixed effects. All regressions include controls for the wife’s age, education, and number of children, as well as the husband’s education. State-by-time fixed effects are included throughout. Standard errors are clustered at the industry level.
Table B.11: Heterogeneous Effects of Husbands’ WFH Shock on Wives’ Labor Market Outcomes by Wives’ Education Level
Women Education Level Men Education level
Below HS Some College Below HS Some College
HS college Grad HS college Grad
A. Outcome: Annual Employment
$Exposed_{o(h(i))}\times Post_{t}$ 0.045** 0.013 -0.002 0.031*** 0.022 -0.006 0.004 0.026***
(0.021) (0.013) (0.008) (0.007) (0.028) (0.013) (0.008) (0.008)
Mean pre-treatment 0.392 0.564 0.678 0.773 0.531 0.678 0.726 0.714
Obs 36,110 97,529 188,412 407,046 45,195 131,857 193,871 358,156
$R^{2}$ 0.07 0.04 0.05 0.05 0.07 0.05 0.05 0.06
B. Outcome: Log Earnings
$Exposed_{o(h(i))}\times Post_{t}$ 0.026 -0.003 0.063** 0.065*** -0.092 0.068** 0.009 0.058***
(0.074) (0.041) (0.027) (0.016) (0.078) (0.031) (0.025) (0.020)
Mean pre-treatment 9.058 9.312 9.566 10.292 9.371 9.621 9.826 10.217
Obs 14,330 54,692 126,993 318,061 24,298 89,103 140,626 260,076
$R^{2}$ 0.11 0.06 0.06 0.07 0.09 0.09 0.09 0.09
C. Outcome: Log Wages
$Exposed_{o(h(i))}\times Post_{t}$ -0.091 -0.040* 0.004 0.026** -0.096* 0.009 -0.021 0.021
(0.056) (0.023) (0.017) (0.011) (0.055) (0.022) (0.017) (0.013)
Mean pre-treatment 2.036 2.172 2.382 2.963 2.202 2.369 2.553 2.950
Obs 14,330 54,692 126,993 318,061 24,298 89,103 140,626 260,076
$R^{2}$ 0.13 0.07 0.07 0.10 0.10 0.10 0.10 0.11
D. Outcome: Weekly Working Hours
$Exposed_{o(h(i))}\times Post_{t}$ 1.510* 0.580 0.444 0.540*** 0.003 0.657* 0.018 0.419**
(0.858) (0.366) (0.298) (0.143) (0.892) (0.336) (0.219) (0.203)
Mean pre-treatment 34.393 34.550 34.853 37.509 35.611 36.102 36.314 36.622
Obs 14,330 54,692 126,993 318,061 24,298 89,103 140,626 260,076
$R^{2}$ 0.08 0.04 0.04 0.03 0.06 0.03 0.03 0.04
E. Outcome: Weeks Worked
$Exposed_{o(h(i))}\times Post_{t}$ 1.292 0.704 1.041*** 0.411*** 0.219 0.857*** 0.714*** 0.546***
(1.071) (0.451) (0.286) (0.145) (0.964) (0.295) (0.244) (0.169)
Mean pre-treatment 41.754 44.376 45.287 46.931 44.101 45.759 46.224 46.279
Obs 14,330 54,692 126,993 318,061 24,298 89,103 140,626 260,076
$R^{2}$ 0.11 0.04 0.03 0.02 0.06 0.04 0.03 0.03
F. Outcome: Part-time Status
$Exposed_{o(h(i))}\times Post_{t}$ -0.083** -0.029* -0.028** -0.021*** 0.005 -0.032*** -0.011 -0.018**
(0.034) (0.015) (0.012) (0.006) (0.029) (0.011) (0.010) (0.008)
Mean pre-treatment 0.367 0.346 0.324 0.231 0.309 0.276 0.267 0.269
Obs 14,330 54,692 126,993 318,061 24,298 89,103 140,626 260,076
$R^{2}$ 0.09 0.04 0.04 0.04 0.06 0.04 0.03 0.05
  • Notes: We estimate the same baseline specification as in Table 2 in the 2017-2023 ACS sample for different samples. For simplicity, we only present the coefficient for the post-Covid–by-treatment interaction. In this Table, we split the sample by the education level of the both spouses in the household.
Table B.12: Heterogeneous Effects of the Husbands’ WFH Shock on Wives Labor Market Outcomes, by Women Age
Women Age Bins
20-24 25-29 30-34 35-39 40-44 45+
A. Outcome: Annual Employment
$Exposed_{o(h(i))}\times Post_{t}$ 0.072** 0.034*** 0.018* 0.023*** 0.023*** 0.018
(0.030) (0.011) (0.009) (0.007) (0.009) (0.019)
Mean pre-treatment 0.608 0.657 0.702 0.715 0.716 0.691
Obs 20,318 96,884 207,244 227,852 133,970 42,804
$R^{2}$ 0.12 0.11 0.09 0.08 0.08 0.10
B. Outcome: Log Earnings
$Exposed_{o(h(i))}\times Post_{t}$ -0.176* 0.034 0.067*** 0.062*** 0.071*** 0.029
(0.106) (0.030) (0.020) (0.024) (0.027) (0.041)
Mean pre-treatment 8.929 9.512 9.909 10.112 10.183 10.171
Obs 12,194 63,915 146,740 164,803 96,577 29,849
$R^{2}$ 0.12 0.10 0.12 0.11 0.12 0.13
C. Outcome: Log Wages
$Exposed_{o(h(i))}\times Post_{t}$ -0.026 -0.024 0.015 0.009 0.021 -0.009
(0.062) (0.025) (0.013) (0.016) (0.018) (0.031)
Mean pre-treatment 1.974 2.345 2.644 2.813 2.889 2.877
Obs 12,194 63,915 146,740 164,803 96,577 29,849
$R^{2}$ 0.15 0.12 0.17 0.17 0.17 0.20
D. Outcome: Weekly Work Hours
$Exposed_{o(h(i))}\times Post_{t}$ -1.119 1.079*** 0.383* 0.310 0.413 0.335
(0.849) (0.298) (0.201) (0.208) (0.305) (0.448)
Mean pre-treatment 33.971 35.258 36.349 36.804 36.777 36.961
Obs 12,194 63,915 146,740 164,803 96,577 29,849
$R^{2}$ 0.13 0.04 0.04 0.04 0.04 0.06
E. Outcome: Weeks Worked
$Exposed_{o(h(i))}\times Post_{t}$ -1.138 0.520 0.938*** 0.783*** 0.615** 0.742
(1.444) (0.370) (0.227) (0.213) (0.251) (0.453)
Mean pre-treatment 40.060 44.334 46.120 46.750 46.793 46.659
Obs 12,194 63,915 146,740 164,803 96,577 29,849
$R^{2}$ 0.11 0.04 0.03 0.02 0.03 0.05
F. Outcome: Part-time Status
$Exposed_{o(h(i))}\times Post_{t}$ 0.045 -0.033*** -0.026*** -0.025*** -0.027** -0.009
(0.034) (0.012) (0.008) (0.008) (0.011) (0.018)
Mean pre-treatment 0.381 0.302 0.264 0.258 0.270 0.272
Obs 12,194 63,915 146,740 164,803 96,577 29,849
$R^{2}$ 0.12 0.05 0.04 0.04 0.04 0.06
  • Notes: We estimate the same baseline specification as in Table 2 in the 2017-2023 ACS sample for different samples. For simplicity, we only present the coefficient for the post-Covid–by-treatment interaction. In this Table, we split the sample by the age of the female spouse in the household.
Table B.13: Effect of Spousal WFH Shocks on Women’s Labor Market Outcomes in Samples Without Children and with Older Children
Employment Log Earnings Log Wages Usual Weekly Hours Weeks Worked Part-time
Worked Status
A. No Children
$Exposed_{o(h(i))}\times Post_{t}$ 0.012** 0.020 0.003 0.147 0.170 -0.011**
(0.005) (0.014) (0.009) (0.143) (0.132) (0.005)
Mean pre-treatment 0.848 10.060 2.605 39.675 47.897 0.162
$\beta_{1}+\beta_{2}=0$ (p-val) 0.075 0.745 0.067 0.522 0.014 0.774
Obs 413,037 351,526 351,526 351,526 351,526 351,526
$R^{2}$ 0.06 0.16 0.18 0.03 0.03 0.03
B. Youngest Child is Teenager (12-17 Years Old)
$Exposed_{o(h(i))}\times Post_{t}$ 0.023*** 0.023 0.001 0.539*** -0.004 -0.013*
(0.006) (0.015) (0.012) (0.153) (0.155) (0.007)
Mean pre-treatment 0.778 10.000 2.622 37.657 47.677 0.238
$\beta_{1}+\beta_{2}=0$ (p-val) 0.038 0.130 0.253 0.241 0.020 0.031
Obs 267,310 207,836 207,836 207,836 207,836 207,836
$R^{2}$ 0.06 0.10 0.15 0.03 0.02 0.03
C. Youngest Child is Adult (18+ Years Old)
$Exposed_{o(h(i))}\times Post_{t}$ 0.007* 0.015 0.006 0.067 0.035 -0.009**
(0.004) (0.012) (0.008) (0.119) (0.116) (0.004)
Mean pre-treatment 0.821 10.012 2.573 39.318 47.777 0.172
$\beta_{1}+\beta_{2}=0$ (p-val) 0.113 0.388 0.082 0.719 0.016 0.452
Obs 545,021 449,335 449,335 449,335 449,335 449,335
$R^{2}$ 0.07 0.15 0.17 0.03 0.03 0.03
  • Notes: This table reports estimates from the cross-sectional specification described in Equation (4), using the 2017–2023 ACS sample of dual-parent households for three different samples. In Panel A we use a sample of households without children. In Panels B and C, we use a sample of household where the youngest child has 12 to 17 years old, and over 17 years old, respectively. All regressions include standard demographic controls (wife’s age, number of children, spouses’ education), as well as fixed effects for state-by-time and the husband’s occupation. Standard errors are clustered at the husband’s occupation level.
Table B.14: Effect of Wives’ WFH Shocks on Their Husbands’ Labor Market Outcomes
Annual Employment Log Earnings Log Wages Weekly Work Hours Weeks Worked Part-time Status
$Exposed_{o(h(i))}\times Post_{t}$ 0.005** 0.019** 0.019** -0.191 0.012 -0.006**
(0.003) (0.009) (0.008) (0.131) (0.094) (0.002)
Mean pre-treatment 0.920 10.574 2.934 44.242 49.675 0.049
Obs 396,791 364,830 364,830 364,830 364,830 364,830
$R^{2}$ 0.02 0.19 0.21 0.02 0.02 0.01
  • Notes: This table reports estimates from the cross-sectional specification described in Equation (4), using the 2017–2023 ACS sample of dual-parent households with children of child-rearing age. All regressions include standard demographic controls (wife’s age, number of children, spouses’ education), as well as fixed effects for state-by-time and the wife’s occupation (or industry, as appropriate). Standard errors are clustered at the wife’s occupation or industry level.
Table B.15: Effects of Husbands’ WFH Shock on Wives’ Intensive Margin of Labor Supply after Adding Wives’ Occupation Fixed Effects
Log Earnings Log Wages Usual Weekly Hours Weeks Worked Part-time status
$Exposed_{o(h(i))}\times Post_{t}$ 0.052*** 0.004 0.477*** 0.666*** -0.027***
(0.011) (0.008) (0.109) (0.110) (0.005)
Mean pre-treatment 9.981 2.732 36.489 46.013 0.266
Obs 514,144 514,144 514,144 514,144 514,144
$R^{2}$ 0.33 0.34 0.15 0.08 0.16
  • Notes: This table presents estimates from Equation (4) on the sample of women working. Treatment is defined based on the change in the probability of WFH before and after the COVID-19 pandemic, using job postings data at the husband’s occupation ($o(h(i))$, Census 2010) levels. The sample is the 2017–2024 CPS sample of dual-parent households with children of child-rearing age. All regressions include fixed effects for state-by-time, husband’s occupation, and women’s occupation fixed effects. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.16: Effect of Wives’ WFH Occupation Shock on Their Own Labor Market Outcomes
Log Earnings Log Wages Usual Weekly Hours Weeks Worked Part-time
Worked Status
$Exposed_{o(h(i))}\times Post_{t}$ 0.046** 0.019 0.185 0.381*** -0.012**
(0.018) (0.011) (0.150) (0.147) (0.005)
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ -0.036* -0.032*** -0.093 -0.109 0.007
(0.019) (0.010) (0.145) (0.212) (0.005)
Mean pre-treatment 9.959 2.692 36.522 46.112 0.269
$\beta_{1}+\beta_{2}=0$ (p-val) 0.430 0.159 0.475 0.078 0.345
Obs 560,544 560,544 560,544 560,544 560,544
$R^{2}$ 0.33 0.33 0.14 0.08 0.15
  • Notes: This table adapts our baseline cross-sectional specification to estimate the effect of women’s WFH shocks on their own labor market outcomes, using the 2017–2023 CPS sample of dual-parent households with children of child-rearing age.
Table B.17: Effects of Husbands’ Occupation and Women College Major WFH Shocks on Women Labor Supply
Employment Log Earnings Log Wage Usual Weekly Hours Weeks Worked Part-time
Worked Status
$Exposed_{o(h(i))}\times Post_{t}$ 0.011** 0.035** -0.011 0.432*** 0.652*** -0.026***
(0.005) (0.014) (0.009) (0.150) (0.143) (0.006)
$Exposed_{e(i)}\times Post_{t}$ 0.011 0.002 -0.006 -0.100 0.283 -0.010
(0.007) (0.023) (0.016) (0.211) (0.254) (0.007)
$Exposed_{o(h(i))}\times Exposed_{e(i)}\times Post_{t}$ 0.028*** 0.047* 0.043** 0.198 -0.122 0.005
(0.008) (0.028) (0.019) (0.275) (0.296) (0.010)
Mean pre-treatment 0.698 9.955 2.695 36.386 46.056 0.272
Obs 727,164 512,022 512,022 512,022 512,022 512,022
$R^{2}$ 0.11 0.21 0.27 0.07 0.06 0.07
  • Notes: This table presents estimates from Equation (7) in each column. Treatment is defined based on the change in the probability of WFH before and after the COVID-19 pandemic, using job postings data at the husband’s occupation ($o(h(i))$, Census 2010) levels and the actual WFH uptake data by women’s major college degree ($e(i)$). The sample is the 2017–2023 ACS sample of dual-parent households with children of child-rearing age. All regressions include fixed effects for state-by-time, husband’s occupation, and women’s major college degree. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.18: Effect of Husbands’ WFH Shock on Marriage, Divorce, and Fertility Outcomes
Married Divorced Had a child
in the last 12mm in the last 12mm in the last 12mm
$Exposed_{o(h(i))}\times Post_{t}$ 0.001 0.000 -0.007**
(0.002) (0.000) (0.003)
Mean pre-treatment 0.031 0.001 0.173
Obs 729,120 728,956 729,120
$R^{2}$ 0.04 0.00 0.32
  • Notes: This table reports estimates using data from the pooled ACS sample from 2017 to 2023. The outcomes—marriage, divorce, and fertility—are defined over the 12 months preceding the interview date. Occupation is measured only at the time of the interview, so we implicitly assume that individuals remained in the same occupation throughout the prior 12-month period.
Table B.19: Effect of Husbands’ WFH Shock on Fertility Outcomes
Pr(At least one child) Number of Children
$Exposed_{o(h(i))}\times Post_{t}$ -0.011*** -0.019**
(0.003) (0.008)
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ 0.006** 0.022***
(0.002) (0.007)
Mean pre-treatment 0.740 1.510
$\beta_{1}+\beta_{2}=0$ (p-val) 0.021 0.664
Obs 1,715,167 1,715,167
$R^{2}$ 0.08 0.11
  • Notes: This table reports estimates from Equation (4) using the pooled 2017–2023 ACS data. The outcomes include the probability of having had a child in the past year and the number of children in the household. Treatment is defined based on the husband’s occupation-level WFH exposure, as detailed in Subsection 3.3. All regressions include controls for the wife’s age, number of children, and education, as well as the husband’s education. State-by-time and husband-occupation fixed effects are included. Standard errors are clustered at the occupation level.
Table B.20: Effect of Husbands’ WFH Shock on Wives’ Labor Market Outcomes Using Reweighting Techniques
Employment Log Earnings Log Wages Usual Weekly Hours Weeks Worked Part-time
Worked Status
A. Using DFL Weights
$Exposed_{o(h(i))}\times Post_{t}$ 0.024*** 0.052*** 0.004 0.480*** 0.660*** -0.026***
(0.005) (0.012) (0.009) (0.113) (0.115) (0.005)
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ -0.009* -0.008 -0.005 0.085 -0.236* 0.001
(0.005) (0.012) (0.009) (0.113) (0.121) (0.005)
Mean pre-treatment 0.698 9.956 2.695 36.390 46.060 0.272
$\beta_{1}+\beta_{2}=0$ (p-val) 0.000 0.000 0.830 0.000 0.000 0.000
Obs 729,120 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.09 0.15 0.20 0.03 0.03 0.04
B. Using IPW
$Exposed_{o(h(i))}\times Post_{t}$ 0.023*** 0.053*** 0.005 0.464*** 0.671*** -0.025***
(0.005) (0.012) (0.009) (0.115) (0.114) (0.005)
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ -0.009* -0.009 -0.006 0.080 -0.235** 0.001
(0.005) (0.012) (0.009) (0.114) (0.119) (0.005)
Mean pre-treatment 0.698 9.956 2.695 36.390 46.060 0.272
$\beta_{1}+\beta_{2}=0$ (p-val) 0.000 0.000 0.884 0.000 0.000 0.000
Obs 729,120 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.08 0.14 0.20 0.03 0.03 0.04
  • Notes: This table reports estimates from Equation (4), using the 2017–2023 ACS sample of dual-parent households with children of child-rearing age. Panel A applies reweighting using DiNardo-Fortin-Lemieux (DFL) weights (DiNardo et al., 1995), while Panel B uses inverse-propensity score weights (IPW). Both methods aim to correct for potential changes in group composition before and after the COVID-19 period. Details on the construction of these weights are provided in Section 6.
Table B.21: Robustness of ACS Results to Alternative Definitions of WFH Shock
Employment Log Earnings Log Wages Usual Weekly Hours Weeks Worked Part-time
Worked Status
A. Shock: Difference in average 2023-2024 vs 2019 WFH vacancy rates at Census 2010 4d level (Original)
$Exposed_{o(h(i))}\times Post_{t}$ 0.024*** 0.052*** 0.004 0.472*** 0.666*** -0.025***
(0.005) (0.012) (0.009) (0.114) (0.114) (0.005)
Mean pre-treatment 0.702 9.969 2.716 36.389 46.030 0.271
Obs 729,120 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.08 0.15 0.20 0.03 0.03 0.04
B. Shock: Difference in 3 month moving average 2023-2024 vs 2019 WFH vacancy rates at Census 2010 4d level
$Exposed_{o(h(i))}\times Post_{t}$ 0.024*** 0.052*** 0.004 0.472*** 0.666*** -0.025***
(0.005) (0.012) (0.009) (0.114) (0.114) (0.005)
Mean pre-treatment 0.702 9.969 2.716 36.389 46.030 0.271
Obs 729,120 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.08 0.15 0.20 0.03 0.03 0.04
C. Shock: Difference in July 2023 vs July 2019 WFH vacancy rates at Census 2010 4d level
$Exposed_{o(h(i))}\times Post_{t}$ 0.025*** 0.050*** 0.006 0.444*** 0.556*** -0.024***
(0.005) (0.012) (0.009) (0.114) (0.116) (0.005)
Mean pre-treatment 0.702 9.969 2.716 36.389 46.030 0.271
Obs 729,120 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.08 0.15 0.20 0.03 0.03 0.04
D. Shock: WFH teleworkability a-la Dingel and Neumann at Census 2010 4d level
$Exposed_{o(h(i))}\times Post_{t}$ 0.029*** 0.040*** 0.001 0.420*** 0.416*** -0.023***
(0.005) (0.012) (0.008) (0.118) (0.118) (0.005)
Mean pre-treatment 0.704 9.973 2.718 36.398 46.066 0.270
Obs 721,305 510,086 510,086 510,086 510,086 510,086
$R^{2}$ 0.08 0.15 0.20 0.03 0.03 0.04
E. Shock: Difference in average 2023-2024 vs 2019 WFH vacancy rates at SOC 3d level
$Exposed_{o(h(i))}\times Post_{t}$ 0.023*** 0.055*** 0.008 0.473*** 0.644*** -0.027***
(0.005) (0.013) (0.008) (0.116) (0.126) (0.006)
Mean pre-treatment 0.703 9.970 2.717 36.391 46.042 0.271
Obs 726,511 512,849 512,849 512,849 512,849 512,849
$R^{2}$ 0.08 0.14 0.19 0.03 0.03 0.03
F. Shock: Difference in 3 month moving average 2023-2024 vs 2019 vacancy rates at SOC 3d level
$Exposed_{o(h(i))}\times Post_{t}$ 0.023*** 0.057*** 0.008 0.492*** 0.670*** -0.027***
(0.005) (0.013) (0.008) (0.117) (0.124) (0.006)
Mean pre-treatment 0.703 9.970 2.717 36.391 46.042 0.271
Obs 726,511 512,849 512,849 512,849 512,849 512,849
$R^{2}$ 0.08 0.14 0.19 0.03 0.03 0.03
G. Shock: Difference in July 2023 vs July 2019 WFH vacancy rates at SOC 3d level
$Exposed_{o(h(i))}\times Post_{t}$ 0.024*** 0.054*** 0.008 0.478*** 0.563*** -0.026***
(0.005) (0.013) (0.008) (0.117) (0.129) (0.006)
Mean pre-treatment 0.703 9.970 2.717 36.391 46.042 0.271
Obs 726,511 512,849 512,849 512,849 512,849 512,849
$R^{2}$ 0.08 0.14 0.19 0.03 0.03 0.03
H. Shock: WFH teleworkability a-la Dingel and Neumann at SOC 3d level
$Exposed_{o(h(i))}\times Post_{t}$ 0.025*** 0.050*** 0.009 0.441*** 0.509*** -0.025***
(0.005) (0.013) (0.008) (0.116) (0.137) (0.006)
Mean pre-treatment 0.704 9.973 2.718 36.398 46.066 0.270
Obs 721,307 510,088 510,088 510,088 510,088 510,088
$R^{2}$ 0.08 0.14 0.19 0.03 0.03 0.03
  • Notes:: In panel A, an occupation is treated if the difference in average 2023 and 2024 vs 2019 remote job vacancy shares is above the median of all other Census 2010 4d level occupations. Panel B uses the 3 month moving average version of the WFH monthly variables. In panel C, an occupation is shocked if the difference in July 2023 vs July 2019 WFH or remote job vacancy shares is above the median of all other Census 2010 4d level occupations. In panel D, an occupation is shocked if its Dingel and Neumann teleworkability index is above the median of all other Census 2010 4d level occupations. Panels E, F, G, H are analogous to A, B, C, D respectively, with the change that occupations are compared at the SOC 2018 3d level. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.22: Robustness to a Continuous ACS Measure of WFH
Employment Log Earnings Log Wages Usual Weekly Hours Weeks Worked Part-time
Worked Status
$Exposed_{o(h(i))}\times Post_{t}$ 0.014*** 0.023*** 0.007* 0.230*** 0.174*** -0.010***
(0.002) (0.005) (0.004) (0.045) (0.046) (0.002)
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ -0.007*** -0.007 -0.005 -0.066 -0.038 0.003
(0.002) (0.005) (0.004) (0.047) (0.050) (0.002)
Mean pre-treatment 0.698 9.956 2.695 36.390 46.060 0.272
$\beta_{1}+\beta_{2}=0$ (p-val) 0.000 0.000 0.291 0.000 0.001 0.000
Obs 729,119 514,150 514,150 514,150 514,150 514,150
$R^{2}$ 0.08 0.15 0.20 0.03 0.03 0.04
  • Notes: This table replicates the baseline ACS estimates using the continuous occupation–year WFH rate constructed from the ACS as the treatment variable, standardized by its mean and standard deviation. A one–standard-deviation increase in this WFH rate corresponds to a 12 percentage point increase in WFH. For comparison, a one–standard-deviation increase in the job-postings-based WFH measure corresponds to an 8 percentage point increase. All specifications include the same controls and fixed effects as in the baseline analysis. Standard errors are clustered at the occupation level.
Table B.23: Robustness of ACS Results to Alternative Fixed Effects Specifications
Employed Log Earnings Log Wages Usual work hours Weeks Worked Part-time status
A. FE: Year + Spouse Occupation (Group of 16) + State
$Exposed_{o(h(i))}\times Post_{t}$ 0.011* 0.081*** 0.044*** 0.361*** 0.536*** -0.020***
(0.006) (0.013) (0.011) (0.125) (0.118) (0.005)
Mean pre-treatment 0.698 9.956 2.695 36.390 46.060 0.272
Obs 516,559 364,843 364,843 364,843 364,843 364,843
$R^{2}$ 0.07 0.14 0.19 0.03 0.03 0.03
B. FE: Year-state + Spouse Occupation (Group of 16)
$Exposed_{o(h(i))}\times Post_{t}$ 0.011* 0.080*** 0.043*** 0.336*** 0.537*** -0.019***
(0.006) (0.013) (0.011) (0.126) (0.117) (0.005)
Mean pre-treatment 0.698 9.956 2.695 36.390 46.060 0.272
Obs 516,559 364,843 364,843 364,843 364,843 364,843
$R^{2}$ 0.07 0.14 0.19 0.03 0.03 0.03
C. FE: Year-state + Spouse Occupation (Group of 16)-by-state
$Exposed_{o(h(i))}\times Post_{t}$ 0.011* 0.080*** 0.044*** 0.340*** 0.540*** -0.019***
(0.006) (0.013) (0.011) (0.128) (0.115) (0.005)
Mean pre-treatment 0.698 9.956 2.695 36.390 46.060 0.272
Obs 516,559 364,843 364,843 364,843 364,843 364,843
$R^{2}$ 0.08 0.14 0.19 0.03 0.03 0.04
D. FE: Year + Spouse Occupation (4d) + State
$Exposed_{o(h(i))}\times Post_{t}$ 0.025*** 0.055*** 0.006 0.495*** 0.667*** -0.026***
(0.005) (0.012) (0.009) (0.113) (0.115) (0.005)
Mean pre-treatment 0.698 9.956 2.695 36.390 46.060 0.272
Obs 516,557 364,841 364,841 364,841 364,841 364,841
$R^{2}$ 0.08 0.15 0.20 0.03 0.03 0.04
E. FE: Year-state + Spouse Occupation (4d) (Baseline)
$Exposed_{o(h(i))}\times Post_{t}$ 0.024*** 0.053*** 0.004 0.469*** 0.669*** -0.025***
(0.005) (0.012) (0.009) (0.114) (0.114) (0.005)
Mean pre-treatment 0.698 9.956 2.695 36.390 46.060 0.272
Obs 516,557 364,841 364,841 364,841 364,841 364,841
$R^{2}$ 0.08 0.15 0.20 0.03 0.03 0.04
F. FE: Year-state + Spouse Occupation (4d)-by-state
$Exposed_{o(h(i))}\times Post_{t}$ 0.021*** 0.056*** 0.007 0.506*** 0.653*** -0.027***
(0.005) (0.013) (0.009) (0.121) (0.119) (0.005)
Mean pre-treatment 0.698 9.957 2.696 36.390 46.060 0.272
Obs 513,864 361,815 361,815 361,815 361,815 361,815
$R^{2}$ 0.12 0.19 0.25 0.08 0.09 0.09
G. FE: Year-MSA + Spouse Occupation (4d)
$Exposed_{o(h(i))}\times Post_{t}$ 0.023*** 0.052*** 0.001 0.492*** 0.677*** -0.026***
(0.006) (0.015) (0.010) (0.146) (0.143) (0.006)
Mean pre-treatment 0.692 10.041 2.774 36.490 46.127 0.270
Obs 395,604 277,637 277,637 277,637 277,637 277,637
$R^{2}$ 0.09 0.15 0.21 0.04 0.04 0.04
  • Notes: This table presents estimates from alternative versions of Equation (4) using the 2017–2023 ACS sample of dual-parent households with children of child-rearing age, varying the fixed effects (FE) structure. Panel E reports the baseline specification, which includes state-by-time and husband-occupation fixed effects. Panel A includes separate fixed effects for month-year, state, and broad occupation groups (collapsed into 16 categories such as Business/Finance, Education Professionals, Healthcare, Sales, Construction, etc.). Panel B adds state-by-time fixed effects, and Panel C includes state-by-time and broad occupation group fixed effects. Panels D to F replicate Panels A to C, respectively, but use detailed 4-digit Census occupation fixed effects instead of the broader groupings. Panel G includes MSA fixed effects, as an alternative to state fixed effects. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.24: Robustness of ACS Results to Adding Broad Occupations and Industries Time Trends
Employed Log Earnings Log Wages Usual work hours Weeks Worked Part-time status
A. FE: Year-state + Spouse Occupation (4d) (baseline)
$Exposed_{o(h(i))}\times Post_{t}$ 0.024*** 0.053*** 0.004 0.469*** 0.669*** -0.025***
(0.005) (0.012) (0.009) (0.114) (0.114) (0.005)
Mean pre-treatment 0.698 9.956 2.695 36.390 46.060 0.272
Obs 516,557 364,841 364,841 364,841 364,841 364,841
$R^{2}$ 0.08 0.15 0.20 0.03 0.03 0.04
B. Year-State + Spouse Occupation (4d) + year-by Spouse Occ (group of 16)
$Exposed_{o(h(i))}\times Post_{t}$ 0.012 0.021 -0.009 0.162 0.550*** -0.026***
(0.008) (0.019) (0.014) (0.174) (0.201) (0.008)
Mean pre-treatment 0.698 9.956 2.695 36.390 46.060 0.272
Obs 516,557 364,841 364,841 364,841 364,841 364,841
$R^{2}$ 0.09 0.15 0.20 0.03 0.03 0.04
C. Year-State + Spouse Occupation (4d) + year-by Spouse Even Broader Occ (group of 4)
$Exposed_{o(h(i))}\times Post_{t}$ 0.013** 0.042** 0.008 0.283* 0.604*** -0.018***
(0.006) (0.016) (0.012) (0.148) (0.132) (0.006)
Mean pre-treatment 0.698 9.956 2.695 36.390 46.060 0.272
Obs 516,557 364,841 364,841 364,841 364,841 364,841
$R^{2}$ 0.08 0.15 0.20 0.03 0.03 0.04
D. Year-State + Spouse Occupation (4d) + year-by Spouse Major Industry
$Exposed_{o(h(i))}\times Post_{t}$ 0.016*** 0.049*** -0.003 0.492*** 0.696*** -0.027***
(0.005) (0.013) (0.009) (0.126) (0.128) (0.005)
Mean pre-treatment 0.703 9.964 2.699 36.416 46.138 0.271
Obs 497,557 353,234 353,234 353,234 353,234 353,234
$R^{2}$ 0.09 0.15 0.21 0.03 0.03 0.04
E. Year-State + Spouse Occupation (4d) + year-by Spouse Broad Ind (group of 4)
$Exposed_{o(h(i))}\times Post_{t}$ 0.021*** 0.048*** 0.003 0.443*** 0.618*** -0.023***
(0.005) (0.013) (0.009) (0.123) (0.117) (0.005)
Mean pre-treatment 0.703 9.964 2.699 36.416 46.138 0.271
Obs 497,557 353,234 353,234 353,234 353,234 353,234
$R^{2}$ 0.09 0.15 0.21 0.03 0.03 0.04
  • Notes: This table presents estimates from alternative versions of Equation (4) using the 2017–2023 ACS sample of dual-parent households with children of child-rearing age, adding year-by-broad occupation or industry fixed effects. Panel A is our benchmark estimation. Panel B adds time trends of a broader occupation classification with 16 categories. Panel C does the same, but with an even broader classification, with 3 categories (blue-collar workers, professionals and those in service/sales occupations). Panel D and E are analogous, but using the industry classification. Panel E divides whether workers are in the public sector, primary sector, private services sector and manufacture. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.25: Robustness of ACS Results to Adding Educational Time Trends by Occupation
Employed Log Earnings Log Wages Usual work hours Weeks Worked Part-time status
A. FE: State + Spouse Occ (4d) + Years of Education (Quintiles) -year
$Exposed_{o(h(i))}\times Post_{t}$ 0.013* 0.059*** 0.016 0.386** 0.733*** -0.021***
(0.007) (0.018) (0.014) (0.167) (0.129) (0.007)
Mean pre-treatment 0.698 9.956 2.695 36.389 46.060 0.272
Obs 516,362 364,707 364,707 364,707 364,707 364,707
$R^{2}$ 0.08 0.15 0.20 0.03 0.03 0.04
B. FE: Year-State FE + Spouse Occ (4d) + Years of Education (Quintiles) -year
$Exposed_{o(h(i))}\times Post_{t}$ 0.012* 0.057*** 0.014 0.371** 0.725*** -0.020***
(0.006) (0.018) (0.014) (0.167) (0.129) (0.007)
Mean pre-treatment 0.698 9.956 2.695 36.389 46.060 0.272
Obs 516,362 364,707 364,707 364,707 364,707 364,707
$R^{2}$ 0.08 0.15 0.20 0.03 0.03 0.04
  • Notes: This table presents estimates from alternative versions of Equation (4) using the 2017–2023 ACS sample of dual-parent households with children of child-rearing age, adding year-by-quintiles in average years of education of occupations fixed effects. Panel A controls by state, occupation and the educational trends fixed effects. Panel B alternatively includes state-by-year fixed effects. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.26: Robustness of ACS Results to Sequential Removal of Demographic Controls
Control Set 1 Control Set 2 Control Set 3 Control Set 4 Control Set 5
Outcome: Employment
$Exposed_{o(h(i))}\times Post_{t}$ 0.019*** 0.019*** 0.020*** 0.024*** 0.024***
(0.005) (0.005) (0.005) (0.005) (0.005)
Mean pre-treatment 0.698 0.698 0.698 0.698 0.698
Obs 729,120 729,120 729,120 729,120 729,120
$R^{2}$ 0.03 0.05 0.05 0.08 0.08
Outcome: Log Earnings
$Exposed_{o(h(i))}\times Post_{t}$ 0.040*** 0.039*** 0.046*** 0.053*** 0.052***
(0.013) (0.013) (0.012) (0.012) (0.012)
Mean pre-treatment 9.956 9.956 9.956 9.956 9.956
Obs 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.08 0.09 0.12 0.15 0.15
Outcome: Log wages
$Exposed_{o(h(i))}\times Post_{t}$ -0.004 -0.004 0.000 0.004 0.004
(0.009) (0.009) (0.009) (0.009) (0.009)
Mean pre-treatment 2.695 2.695 2.695 2.695 2.695
Obs 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.14 0.14 0.17 0.20 0.20
Outcome: Usual work hours
$Exposed_{o(h(i))}\times Post_{t}$ 0.425*** 0.426*** 0.454*** 0.472*** 0.472***
(0.114) (0.115) (0.115) (0.114) (0.114)
Mean pre-treatment 36.390 36.390 36.390 36.390 36.390
Obs 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.02 0.02 0.03 0.03 0.03
Outcome: Weeks Worked
$Exposed_{o(h(i))}\times Post_{t}$ 0.597*** 0.587*** 0.631*** 0.666*** 0.666***
(0.116) (0.116) (0.113) (0.114) (0.114)
Mean pre-treatment 46.060 46.060 46.060 46.060 46.060
Obs 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.01 0.02 0.03 0.03 0.03
Outcome: Part-time status
$Exposed_{o(h(i))}\times Post_{t}$ -0.023*** -0.023*** -0.025*** -0.025*** -0.025***
(0.005) (0.005) (0.005) (0.005) (0.005)
Mean pre-treatment 0.272 0.272 0.272 0.272 0.272
Obs 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.02 0.03 0.03 0.04 0.04
Control: Number of children No Yes Yes Yes Yes
Control: Women’s age (5y sbin) No No Yes Yes Yes
Control: Women’s education level No No No Yes Yes
Control: Husband’s education level No No No No Yes
  • Notes: This table reports results from estimating Equation (4) on the 2017–2023 ACS sample of dual-parent households with children of child-rearing age. The baseline specification, shown in Column 5, includes controls for the wife’s age (in 5-year bins), number of children in the household, and both spouses’ education levels. Columns 1–4 sequentially remove these controls to assess the sensitivity of the results to the inclusion of covariates. All specifications include state-by-time and husband-occupation fixed effects.
Table B.27: Robustness of ACS Results to Adding Controls for Children’s Age
Control Set 1 Control Set 2 Control Set 3 Control Set 4 Control Set 5
Outcome: Employment
$Exposed_{o(h(i))}\times Post_{t}$ 0.024*** 0.023*** 0.024*** 0.024*** 0.024***
(0.005) (0.005) (0.005) (0.005) (0.005)
Mean pre-treatment 0.698 0.698 0.698 0.698 0.698
Obs 729,120 729,120 729,120 729,120 729,120
$R^{2}$ 0.08 0.09 0.09 0.09 0.09
Outcome: Log Earnings
$Exposed_{o(h(i))}\times Post_{t}$ 0.052*** 0.051*** 0.052*** 0.052*** 0.052***
(0.012) (0.012) (0.012) (0.012) (0.012)
Mean pre-treatment 9.956 9.956 9.956 9.956 9.956
Obs 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.15 0.15 0.15 0.15 0.15
Outcome: Log wages
$Exposed_{o(h(i))}\times Post_{t}$ 0.004 0.004 0.004 0.004 0.004
(0.009) (0.009) (0.009) (0.009) (0.009)
Mean pre-treatment 2.695 2.695 2.695 2.695 2.695
Obs 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.20 0.20 0.21 0.21 0.21
Outcome: Usual work hours
$Exposed_{o(h(i))}\times Post_{t}$ 0.472*** 0.465*** 0.482*** 0.472*** 0.473***
(0.114) (0.114) (0.114) (0.113) (0.113)
Mean pre-treatment 36.390 36.390 36.390 36.390 36.390
Obs 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.03 0.03 0.03 0.03 0.03
Outcome: Weeks Worked
$Exposed_{o(h(i))}\times Post_{t}$ 0.666*** 0.650*** 0.670*** 0.654*** 0.658***
(0.114) (0.115) (0.115) (0.116) (0.115)
Mean pre-treatment 46.060 46.060 46.060 46.060 46.060
Obs 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.03 0.03 0.03 0.03 0.03
Outcome: Part-time status
$Exposed_{o(h(i))}\times Post_{t}$ -0.025*** -0.025*** -0.026*** -0.026*** -0.026***
(0.005) (0.005) (0.005) (0.005) (0.005)
Mean pre-treatment 0.272 0.272 0.272 0.272 0.272
Obs 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.04 0.04 0.04 0.04 0.04
Control: Baseline controls Yes Yes Yes Yes Yes
Control: Age youngest child No Yes No Yes Yes
Control: Age oldest child (raw) No No Yes Yes No
Control: Age oldest child (4y bins) No No No No Yes
  • Notes: This table reports estimates from Equation (4) using the 2017–2023 ACS sample of dual-parent households with children of child-rearing age. Column 1 reproduces the baseline specification, which includes controls for the wife’s age (in 5-year bins), number of children in the household, and both spouses’ education levels. In Columns 2 through 5, we sequentially add controls for the age of the youngest and oldest child, both as continuous variables and in four-year bins. All regressions include state-by-time and husband-occupation fixed effects.
Table B.28: Robustness to Excluding Households with Self-Employed Husbands
Employment Log Earnings Log Wages Usual Weekly Hours Weeks Worked Part-time
Worked Status
$Exposed_{o(h(i))}\times Post_{t}$ 0.024*** 0.052*** 0.004 0.472*** 0.666*** -0.025***
(0.005) (0.012) (0.009) (0.114) (0.114) (0.005)
Mean pre-treatment 0.702 9.969 2.716 36.389 46.030 0.271
Obs 729,120 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.08 0.15 0.20 0.03 0.03 0.04
  • Notes: This table reports results from Equation (4) using the 2017–2023 ACS sample of dual-parent households with children of child-rearing age. We restrict the sample to households where the husband is classified as an employee (i.e., excluding households in which the husband reports being self-employed). The treatment is defined based on the husband’s occupation-level WFH shock, as in the baseline specification. All regressions include state-by-time and husband-occupation fixed effects, as well as controls for the wife’s age, number of children, and both spouses’ education levels.
Table B.29: Robustness of Wage Results to Trimming High-Income Households
Employment Log Earnings Log Wages Usual Weekly Hours Weeks Worked Part-time
Worked Status
A. No income trim (Baseline)
$Exposed_{o(h(i))}\times Post_{t}$ 0.024*** 0.052*** 0.004 0.472*** 0.666*** -0.025***
(0.005) (0.012) (0.009) (0.114) (0.114) (0.005)
Mean pre-treatment 0.702 9.969 2.716 36.389 46.030 0.271
Obs 729,120 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.08 0.15 0.20 0.03 0.03 0.04
B. Trim top 1% income earners
$Exposed_{o(h(i))}\times Post_{t}$ 0.023*** 0.051*** 0.003 0.481*** 0.670*** -0.026***
(0.005) (0.012) (0.009) (0.117) (0.115) (0.005)
Mean pre-treatment 0.701 9.949 2.700 36.316 45.995 0.272
Obs 721,361 507,784 507,784 507,784 507,784 507,784
$R^{2}$ 0.08 0.14 0.19 0.03 0.03 0.04
C. Trim top 5% income earners
$Exposed_{o(h(i))}\times Post_{t}$ 0.022*** 0.045*** -0.003 0.478*** 0.682*** -0.026***
(0.005) (0.012) (0.009) (0.117) (0.117) (0.005)
Mean pre-treatment 0.702 9.912 2.665 36.222 45.961 0.273
Obs 691,771 487,504 487,504 487,504 487,504 487,504
$R^{2}$ 0.09 0.13 0.18 0.03 0.03 0.04
  • Notes: This table reports robustness checks on women’s log wage outcomes using the 2017–2023 ACS sample of dual-parent households with children of child-rearing age. We examine whether our baseline wage results are sensitive to the exclusion of households at the top of the income distribution. Panel A reports the baseline specification without trimming, included for comparison. Subsequent panels show results after progressively trimming the top percentiles of the income distribution. The treatment is based on the husband’s occupation-level WFH shock, as described in Subsection 3.3. All regressions include state-by-time and husband-occupation fixed effects, as well as controls for the wife’s age, number of children, and both spouses’ education levels.
Table B.30: Test: Leave-One Occupation Out using Census 2d
Employment Log Earnings Log Wages Usual Weekly Hours Weeks Worked Part-time
Worked Status
Benchmark Estimate 0.024 0.053 0.004 0.469 0.669 -0.025
Average Bootstrap Estimate 0.024 0.053 0.004 0.469 0.669 -0.025
Average Bootstrap SE 0.005 0.012 0.009 0.115 0.115 0.005
P-Statistic 0.000 0.000 1.000 0.000 0.000 0.000
  • Notes: The table reports the main estimates of the analysis for the labor outcomes of women in the ACS sample between 2017 and 2023. Benchmark Estimate reports estimates from Equation (4). Average Boostrap Estimate is the average from the estimates obtained from dropping in every iteration a different 2-digit occupation at a time. The Average Standard Error is the mean from the standard errors in each iteration from leaving a different occupation out from estimations. The P-statistic captures the percentage of estimates from the iterations that were not significant at the 5% level. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors from each iteration are clustered at the husband’s occupation level.
Table B.31: Effect of Husbands’ WFH Shock on Wives Labor Market Outcomes Using a Placebo Shock Timing
Employment Log Earnings Log Wages Usual Weekly Hours Weeks Worked Part-time
Worked Status
$Exposed_{o(h(i))}\times Post_{t}$ 0.011*** -0.004 -0.024*** 0.031 0.226* 0.000
(0.004) (0.011) (0.008) (0.105) (0.122) (0.004)
Mean pre-treatment 0.689 9.933 2.673 36.390 45.664 0.276
Obs 408,112 286,403 286,403 286,403 286,403 286,403
$R^{2}$ 0.09 0.15 0.21 0.03 0.03 0.04
  • Notes: This table reports results from a placebo test in which the COVID-19 treatment period is artificially shifted to 2018, so that the post-Covid period comprises 2018 and 2019. The specification follows the baseline cross-sectional model, estimated using the 2017–2020 ACS subsample where the reference years are 2016-2019.
Table B.32: Effect of Husbands’ WFH Shock on Wives’ Labor Market Outcomes Using the CPS
Previous week Previous year (March CPS)
Employment Usual Weekly Part-time Employment Log Earnings Log Wages Usual Weekly Weeks Worked Part-time
Work hours status Work Hours Status
$Exposed_{o(h(i))}\times Post_{t}$ 0.027*** 0.300 -0.007 0.035** 0.082** 0.010 1.199*** 0.596 -0.030*
(0.009) (0.194) (0.008) (0.015) (0.036) (0.026) (0.362) (0.426) (0.016)
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ -0.006 0.057 -0.001 -0.010 0.007 0.000 -0.143 0.520 -0.009
(0.008) (0.228) (0.010) (0.015) (0.035) (0.025) (0.382) (0.424) (0.016)
Mean pre-treatment 0.672 36.631 0.313 0.665 9.992 2.696 36.441 46.702 0.247
Obs 395,431 254,834 251,572 58,927 39,785 39,785 39,785 39,785 39,785
$R^{2}$ 0.12 0.06 0.05 0.11 0.17 0.21 0.06 0.06 0.06
  • Notes: This table reports estimates from Equation (4) using the 2016–2024 CPS sample of dual-parent households with children of child-rearing age. The Monthly CPS sample tracks individuals monthly in a short panel rotational basis, which we do not exploit, and does not ask about income. The March CPS tracks income in addition to other variables and annual employment variables. All regressions control for the wife’s age, education, and number of children, as well as the husband’s education. Specifications include state-by-time and husband-occupation fixed effects. Standard errors are clustered at the occupation level.
Table B.33: Effect of Husbands’ WFH Shock on Wives’ Labor Market Outcomes Using the SIPP Panel Component
Employed Log Earnings Log Wages Avg Weekly Part-time
last month Work Hours Status
$Exposed_{o(h(i))}\times Post_{t}$ 0.050 0.248** 0.058 3.392* -0.062
(0.037) (0.125) (0.120) (1.935) (0.068)
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ -0.047 -0.252** -0.097 -2.487 0.005
(0.031) (0.125) (0.119) (1.862) (0.064)
Mean pre-treatment 0.694 7.736 2.857 35.429 0.298
Total Obs 12,763 8,309 8,133 8,713 8,713
Unique HHs 594 438 433 449 449
$R^{2}$ 0.84 0.80 0.76 0.74 0.77
  • Notes: We estimate an alternative to the baseline specification that exploits the SIPP panel component. We replace cross-section FE with household-level and time-level (i.e. month-year) fixed effects. In addition we also ensure that treatment status has not change from treated to untreated (or viceversa) in the sample (i.e. we enforce that spouses must have been in similarly exposed occupations across any waves they appear). We use a sample of households that joined the SIPP on 2020 and thus were covered during reference years 2019 to 2022. In practice, this is the only subsample of the SIPP that covers the pre- and post-Covid period. Earnings are monthly since the data is recorded at the monthly level. All regressions include controls for the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.34: Effect of Husbands’ WFH Shock on Time Spent in Labor and Household Supply - Share of the Day
Time Spent in Labor Production Time Spent for Personal Care and Leisure
Working Main Work Main Work Main Work Secondary Work Commuting Personal Care Leisure
From Home Outside Home
Wives:
$Exposed_{o(h(i))}\times Post_{t}$ 0.072*** 0.071*** 0.049*** 0.022 -0.000 0.001 -0.010 0.025**
(0.017) (0.017) (0.013) (0.016) (0.001) (0.002) (0.011) (0.011)
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ -0.080*** -0.083*** -0.039** -0.044** 0.002** -0.000 0.005 -0.030***
(0.020) (0.021) (0.016) (0.019) (0.001) (0.002) (0.011) (0.011)
Mean pre-treatment 0.172 0.170 0.025 0.145 0.001 0.012 0.374 0.112
$\beta_{1}+\beta_{2}=0$ (p-val) 0.675 0.547 0.398 0.228 0.097 0.906 0.591 0.656
Obs 11,235 11,235 11,235 11,235 11,235 11,235 11,235 11,235
$R^{2}$ 0.83 0.83 0.85 0.82 0.91 0.81 0.80 0.84
Husbands:
$Exposed_{o(h(i))}\times Post_{t}$ 0.013 0.013 0.051*** -0.038** 0.004 -0.012*** 0.006 -0.023*
(0.016) (0.016) (0.011) (0.018) (0.003) (0.004) (0.011) (0.014)
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ -0.016 -0.012 0.049*** -0.061** -0.003 -0.000 0.023** 0.010
(0.021) (0.022) (0.018) (0.025) (0.004) (0.004) (0.011) (0.014)
Mean pre-treatment 0.317 0.309 0.028 0.280 0.007 0.029 0.350 0.116
$\beta_{1}+\beta_{2}=0$ (p-val) 0.859 0.985 0.000 0.000 0.882 0.001 0.005 0.320
Obs 11,440 11,440 11,440 11,440 11,440 11,440 11,440 11,440
$R^{2}$ 0.84 0.84 0.85 0.85 0.79 0.83 0.82 0.82
Time Spent in Household Production
Housework Cooking Shopping Primary Secondary Total Total Prop. Secondary CC
Childcaring CC while working Secondary CC Childcaring in Total CC
Wives:
$Exposed_{o(h(i))}\times Post_{t}$ -0.007 -0.006 -0.006** -0.025** 0.012 0.011 -0.014 0.060**
(0.008) (0.005) (0.003) (0.011) (0.007) (0.020) (0.020) (0.030)
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ 0.009 0.008 0.004 0.050*** -0.022** -0.005 0.045* -0.064**
(0.011) (0.006) (0.003) (0.013) (0.011) (0.022) (0.025) (0.032)
Mean pre-treatment 0.093 0.045 0.016 0.106 0.016 0.228 0.334 0.645
$\beta_{1}+\beta_{2}=0$ (p-val) 0.831 0.625 0.519 0.027 0.363 0.786 0.190 0.882
Obs 11,235 11,235 11,235 11,235 11,235 11,235 11,235 11,046
$R^{2}$ 0.84 0.85 0.82 0.82 0.83 0.82 0.82 0.83
Husbands:
$Exposed_{o(h(i))}\times Post_{t}$ 0.001 0.002 -0.002 0.000 0.015** 0.011 0.012 -0.053
(0.008) (0.003) (0.002) (0.008) (0.007) (0.018) (0.019) (0.044)
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ 0.001 -0.006* 0.004 0.004 0.022 0.019 0.023 0.035
(0.011) (0.003) (0.003) (0.007) (0.013) (0.029) (0.029) (0.047)
Mean pre-treatment 0.041 0.016 0.007 0.051 0.008 0.127 0.177 0.649
$\beta_{1}+\beta_{2}=0$ (p-val) 0.811 0.125 0.318 0.531 0.004 0.223 0.142 0.713
Obs 11,440 11,440 11,440 11,440 11,440 11,440 11,440 10,585
$R^{2}$ 0.83 0.88 0.85 0.84 0.80 0.80 0.81 0.83
  • Notes: We use data from the the 2017-2024 ATUS panel datasets. We show results for the time spent in labor, household and personal activities available in ATUS. Time-use variables are recorded by one member in the household. Then, as Section 3 points out, we have two draws from ATUS: the men-respondent sample and the women-respondent sample. While the former has data on the time spent doing activities by husbands, the latter will have for wives. There is a first panel that has dependent variables that capture time spent in labor production, as a share of the whole day: working, working at their main work, working at their main work from home and outside home, working at their secondary work and commuting to work. Those that do not work should have a 0 imputed in all working variables. There is a panel with dependent variables related to personal care (i.e. sleeping or taking a shower) and leisure (ie. going to religious events or sport activities). Lastly there is a household panel with dependent variables that capture the time spent doing the following activities: housework, cooking, shopping, primary childcaring (report doing childcaring as main activity), childcaring while working, total childcaring as secondary activity (childcaring while doing primary activities), total childcaring (the sum of primary and secondary childcaring) and the share of secondary childcaring on total childcaring. We restrict our sample to weekdays and non-holiday days. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the occupation level.
Table B.35: Effects of Husbands’ Occupation and Women College Major WFH Shocks on Women WFH Uptake and Commuting
WFH Uptake Commuting
$Exposed_{o(h(i))}\times Post_{t}$ 0.019*** -0.373
(0.005) (0.265)
$Exposed_{e(i)}\times Post_{t}$ 0.044*** -2.287***
(0.008) (0.385)
$Exposed_{o(h(i))}\times Exposed_{e(i)}\times Post_{t}$ 0.038*** -0.704
(0.010) (0.464)
Mean pre-treatment 0.083 24.488
Obs 482,358 482,298
$R^{2}$ 0.11 0.04
  • Notes: This table presents estimates from Equation (7) in each column. Treatment is defined based on the change in the probability of WFH before and after the COVID-19 pandemic, using job postings data at the husband’s occupation ($o(h(i))$, Census 2010) levels and the actual WFH uptake data by women’s major college degree ($e(i)$). The sample is the 2017–2023 ACS sample of dual-parent households with children of child-rearing age. All regressions include fixed effects for state-by-time, husband’s occupation, and women’s major college degree. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.36: Effect of Husbands’ WFH Occupation Shocks on Childcare Time and Child-Related Activities
Childcare by Childcare by Childcare by Childcare by Childcare by Childcare Paid for Log Childcare
Wife while Spouse while Grandparent Relative Non-relative by Daycare Childcare Expenditure
Working Wife Works
A. Entire sample
$Exposed_{o(h(i))}\times Post_{t}$ 0.004 0.064* -0.001 0.023 -0.009 -0.053* -0.073** -0.165
(0.032) (0.036) (0.031) (0.032) (0.021) (0.028) (0.031) (0.117)
Mean pre-treatment 0.330 0.485 0.416 0.454 0.186 0.266 0.484 4.710
Obs 60,264 60,264 85,379 85,379 85,379 85,379 68,047 32,910
$R^{2}$ 0.15 0.16 0.15 0.14 0.13 0.18 0.19 0.33
  • Notes: We use data from the the 2018-2024 SIPP panel datasets. We estimate our baseline cross-section specification which has state-by-month-year and spouse occupation fixed effects. We show results for childcare variables available in the SIPP, which ask households with minor of age children if in a typical week, the following are involved at least one day in looking after children: The wife and her spouse (when the wife is working), a grandparent, another relative (includes grandparent), a daycare. We also include a dummy on whether the family pays for daycare, and self-reported log expenditure in childcare. All regressions include fixed effects for state-by-time, and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.37: Effect of Husbands’ WFH Shock on Child Well-Being and Activities
Days wife Days spouse Days wife Days spouse Child is in Child takes Child plays
has dinner has dinner takes child takes child club/activity lessons after sports in/out
with child with child on outings on outings after school school of school
A. Entire sample
$Exposed_{o(H)}\times Post_{t}$ -0.029 0.082 0.175 0.080 0.056 0.042 -0.085
(0.086) (0.180) (0.234) (0.245) (0.059) (0.062) (0.060)
Mean pre-treatment 6.532 5.817 3.533 2.677 0.353 0.486 0.535
Obs 44,620 44,620 31,873 31,873 27,705 27,705 27,705
$R^{2}$ 0.18 0.23 0.25 0.24 0.27 0.32 0.30
  • Notes: We use data from the the 2018-2024 SIPP panel datasets. We estimate our baseline cross-section specification which has state-by-month-year and spouse occupation fixed effects. We show results for childcare variables available in the SIPP, which ask households with minor of age children if in a typical week, the following are involved at least one day in looking after children: The wife and her spouse (when the wife is working), a grandparent, another relative (includes grandparent), a daycare. We also include a dummy on whether the family pays for daycare, and self-reported log expenditure in childcare.
Table B.38: Mean and SD from CPS Compared to Women-Respondent and Men-Respondent in the ATUS Sample
CPS ATUS
Sample Women-Resp Sample Men-Resp Sample
Mean SD Mean SD Mean SD
Panel A: Husband Characteristics
Husband In Shocked Occupation (Treatment) 0.51 0.50 0.56 0.50 0.61 0.49
Husband Employed (Last Year) 0.91 0.29 0.92 0.27 0.93 0.25
Husband Log Earnings 10.97 0.69 11.09 0.65 10.82 0.70
Husband Wage 34.98 47.64 35.23 19.94 34.63 69.97
Husband Usual Work Hours 43.51 9.35 43.88 9.40 43.62 9.05
Husband Weeks Worked 50.11 6.78 50.27 6.33 50.08 7.00
Husband Part-time (Last Year) 0.03 0.18 0.03 0.16 0.04 0.19
Husband - Below HS Level 0.08 0.27 0.05 0.22 0.04 0.21
Husband - HS Level 0.23 0.42 0.18 0.38 0.15 0.36
Husband - Above HS Level 0.69 0.46 0.77 0.42 0.81 0.40
Husband Age 37.67 6.82 37.92 6.33 37.93 6.43
Panel B: Wife Characteristics
Wife Employed (Last Year) 0.67 0.47 0.68 0.47 0.71 0.45
Log Earnings 10.98 0.72 10.82 0.81 11.15 0.58
Log Wages 3.57 0.74 3.38 0.66 3.75 0.76
Wife Usual Work Hours 37.23 10.44 37.50 10.94 36.90 10.69
Wife Weeks Worked 48.29 9.42 48.92 8.33 48.51 8.69
Wife Part-time (Last Year) 0.22 0.42 0.26 0.44 0.23 0.42
Wife - Below HS Level 0.06 0.24 0.05 0.21 0.03 0.18
Wife - HS Level 0.17 0.38 0.11 0.32 0.10 0.30
Wife - Above HS Level 0.76 0.42 0.84 0.37 0.86 0.34
Wife Age 35.23 5.93 35.52 5.35 35.84 5.50
Panel C: HH Characteristics
HH Wage 76074.47 77836.21 85452.19 83882.00 86731.10 77692.49
Number of own children in household 2.08 0.89 2.01 0.86 2.01 0.84
Age of Youngest Child 3.48 2.54 3.39 2.54 3.47 2.55
Age of Oldest Child 7.54 5.10 6.89 4.54 6.87 4.54
Youngest Child 0-4 0.64 0.48 0.66 0.47 0.65 0.48
Youngest Child 5-8 0.36 0.48 0.34 0.47 0.35 0.48
  • Notes: The table reports mean and standard deviations between the different samples: the benchmark CPS sample from the main findings and the ATUS samples. There are two ATUS samples: women respondents sample and men respondents sample. The former is when wives are those that report their time diary of their activities, while the latter is when husbands report them. Panel A captures variables related to husband characteristics, Panel B is related to wives characteristics and Panel C are related to household variables.
Table B.39: Effect of Husbands’ WFH Shock on Wives’ ATUS Labor Market Outcomes
Women-Reported Sample Men-Reported Sample
Employed Log Earnings Log Wages Usual Work Weeks Worked Part-time Employed Log Earnings Log Wages Usual Work Weeks Worked Part-time
Hours Status Hours Status
$Exposed_{o(h(i))}\times Post_{t}$ 0.047 8.980*** -1.152*** 4.404** 14.176*** -0.117* 0.073 2.759*** 2.908*** 1.962 -2.626 -0.046
(0.060) (0.000) (0.000) (1.883) (3.530) (0.064) (0.057) (0.000) (0.087) (1.539) (1.784) (0.058)
Mean pre-treatment 0.722 10.472 3.080 35.894 49.632 0.300 0.666 10.699 3.014 35.517 51.241 0.294
Obs 7,989 202 172 5,476 256 5,780 8,175 210 203 5,506 465 5,688
$R^{2}$ 0.84 1.00 0.96 0.89 0.84 0.90 0.85 1.00 0.99 0.91 0.71 0.92
  • Notes: We use data from the the 2017-2024 ATUS panel datasets. Time-use variables are recorded by one member in the household. Then, as Section 3 points out, we have two draws from ATUS: the men-respondent sample and the women-respondent sample. We use Labor Market Outcomes of wives from ATUS data, different from the ones from the CPS surveys. While in the women-respondent sample, wives report their own labor outcomes, in the men-respondent, their spouses respond for them. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.40: Effect of Husbands’ WFH Shock on Time Spent in Labor and Household Supply - Share of the Day (Placebo Timing)
Time Spent in Labor Production Time Spent for Personal Care and Leisure
Working Main Work Main Work Main Work Secondary Work Commuting Personal Care Leisure
From Home Outside Home
Wives:
$Exposed_{o(h)}\times Post_{t}$ -0.024 -0.018 -0.003 -0.015 -0.001 -0.006 -0.008 -0.013
(0.025) (0.025) (0.015) (0.030) (0.002) (0.005) (0.012) (0.018)
Mean pre-treatment 0.166 0.165 0.025 0.140 0.001 0.013 0.373 0.108
Obs 4,488 4,488 4,488 4,488 4,488 4,488 4,488 4,488
$R^{2}$ 0.86 0.86 0.88 0.85 0.86 0.84 0.87 0.88
Husbands:
$Exposed_{o(h)}\times Post_{t}$ -0.024 -0.025 0.004 -0.029 0.002 -0.010 0.014 0.031
(0.032) (0.030) (0.015) (0.032) (0.007) (0.008) (0.014) (0.026)
Mean pre-treatment 0.324 0.312 0.026 0.286 0.009 0.036 0.353 0.113
Obs 4,483 4,483 4,483 4,483 4,483 4,483 4,483 4,483
$R^{2}$ 0.88 0.88 0.85 0.87 0.86 0.83 0.87 0.86
Time Spent in Household Production
Housework Cooking Shopping Primary Secondary Total Total Prop. Secondary CC
Childcaring CC while working Secondary CC Childcaring in Total CC
Wives:
$Exposed_{o(h)}\times Post_{t}$ 0.023 0.011 -0.007 0.027 0.062 -0.001 0.026 -0.037
(0.016) (0.007) (0.004) (0.019) (0.097) (0.032) (0.034) (0.045)
Mean pre-treatment 0.097 0.047 0.017 0.110 0.104 0.237 0.347 0.641
Obs 4,488 4,488 4,488 4,488 2,605 4,488 4,488 4,407
$R^{2}$ 0.89 0.86 0.85 0.85 0.93 0.87 0.86 0.86
Husbands:
$Exposed_{o(h)}\times Post_{t}$ -0.012 -0.002 0.003 0.010 -0.031 0.023 0.033 0.026
(0.010) (0.005) (0.003) (0.012) (0.055) (0.039) (0.045) (0.059)
Mean pre-treatment 0.035 0.014 0.007 0.051 0.031 0.115 0.166 0.629
Obs 4,483 4,483 4,483 4,483 3,932 4,483 4,483 4,122
$R^{2}$ 0.86 0.84 0.88 0.84 0.84 0.85 0.85 0.86
  • Notes: We use data from the the 2017-2024 ATUS panel datasets. We show results for the time spent in labor, household and personal activities available in ATUS. Time-use variables are recorded by one member in the household. Then, as Section 3 points out, we have two draws from ATUS: the men-respondent sample and the women-respondent sample. While the former has data on the time spent doing activities by husbands, the latter will have for wives. There is a first panel that has data on the time spent in labor production, capturing the share of the whole day - over 24 hours - that respondents spent: working, working at their main work, working at their main work from home and outside home, working at their secondary work and commuting to work. There is a panel on personal care (i.e. sleeping or taking a shower) and leisure (ie. going to religious events or sport activities). Lastly there is household panel with the time spent doing the following activities: housework, cooking, shopping, primary childcaring (report doing childcaring as main activity), childcaring while working, total childcaring as secondary activity while doing other activities, total childcaring (the sum of primary and secondary childcaring) and the share of secondary childcaring on total childcaring. We restrict our sample to weekdays and non-holiday days. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level. We exploit a fake timing as in Table B.31 by moving the arrival of the pandemic shock to March 2018.
Appendix C Discussion of Identification Strategy
Our empirical design estimates the causal effect of an increase in the husband’s WFH probability on the wife’s labor market outcomes. A key data limitation, however, is that in many surveys – including the ACS – occupational information is observed only for individuals currently employed. This constraint implies two challenges. First, we cannot systematically observe a wife’s own occupation when she is not employed, preventing us from measuring her WFH exposure during non-employment spells. Second, because women’s own exposure is unobserved for part of the sample, our baseline estimates necessarily combine the spillover effect arising from husbands’ WFH exposure with the direct effect operating through women’s own WFH exposure, which in the model is generated by assortative matching, as explained in Section 4.
To illustrate the implications, consider a simplified version of our baseline specification:
$y_{it}=\alpha+\beta_{0}D^{o(h(i))}+\gamma D_{t}+\beta_{h}D^{o(h(i))}\times D_{t}+u_{it},$ (8)
where $D^{o(h(i))}=1$ if the husband of woman $i$ is in a treated occupation (i.e., one with a large WFH shock), and $D_{t}=1$ in post-COVID years.
Suppose that unobserved variation in the wife’s own WFH exposure enters the error term as:
$u_{it}=\beta_{1}D^{o(i)}+\beta_{w}D^{o(i)}\times D_{t}+\beta_{2}D^{o(i)}\times D^{o(h(i))}+\beta_{hw}D^{o(i)}\times D^{o(h(i))}\times D_{t}+\varepsilon_{it},$ (9)
where $D^{o(i)}=1$ if the wife herself works in a treated occupation.
Even if we assume the cross-occupation interaction terms $\beta_{2}$ and $\beta_{hw}$ are zero, the difference-in-differences estimator $\hat{\beta}_{h}$ from (8) identifies:
$\begin{split}\hat{\beta}_{h}=&\;\beta_{h}+\beta_{1}\cdot\Big[\underbrace{P(D^{o(i)}=1\mid D^{o(h)}=1,\text{Post})-P(D^{o(i)}=1\mid D^{o(h)}=1,\text{Pre})}_{\Delta P(D^{o(i)}=1\mid D^{o(h)}=1)}\\[2.5pt] &\hskip 31.0001pt-\underbrace{\big(P(D^{o(i)}=1\mid D^{o(h)}=0,\text{Post})-P(D^{o(i)}=1\mid D^{o(h)}=0,\text{Pre})\big)}_{\Delta P(D^{o(i)}=1\mid D^{o(h)}=0)}\Big]\\[5.0pt] &\;-\;\beta_{w}\cdot\Big(P(D^{o(i)}=1\mid D^{o(h)}=1,\text{Post})-P(D^{o(i)}=1\mid D^{o(h)}=0,\text{Post})\Big).\end{split}$ (10)
Even under the assumption that assortative matching is stable before and after COVID— i.e.,
$\Delta P(D^{o(i)}=1\mid D^{o(h)}=1)=\Delta P(D^{o(i)}=1\mid D^{o(h)}=0)=0,$
the estimator $\hat{\beta}_{h}$ will still load on the direct effect $\beta_{w}$ through differences in wives’ occupational exposure across treated and untreated husbands in the post period.
Section 6 presents several exercises designed to separate these two components and quantify the relative contributions of spillover and direct effects.